### v1

# A Simple, Optimal and Efficient Algorithm for Online Exp-Concave Optimization

Abstract
[…]
This paper proposes a simple variant of ONS, called LightONS, which reduces the total runtime to $O(d^{2}T+d^{\omega}\sqrt{T\log T})$ while preserving the optimal $O(d\log T)$ regret. For SXO, LightONS implies a stochastic optimization method with runtime $\widetilde{O}(d^{3}/\epsilon)$, thereby answering the open problem. Importantly, LightONS preserves the elegant structure of ONS by leveraging domain-conversion techniques from parameter-free online learning to introduce a hysteresis mechanism that delays expensive Mahalanobis projections until necessary. This design enables LightONS to serve as an efficient plug-in replacement of ONS in broader scenarios, even beyond regret minimization, including gradient norm adaptive regret, parametric stochastic bandits, and memory-efficient online learning.
1 Introduction
Online Convex Optimization (OCO) provides a versatile framework for online learning, with deep connections to stochastic optimization, game theory, and information theory (Cesa-Bianchi and Lugosi, 2006; Hazan, 2016). Online eXp-concave Optimization (OXO), is an important instance, where each online loss function is exponentially concave (exp-concave), meaning that $\exp(-\alpha f(\mathbf{x}))$ is concave for some $\alpha>0$ with $f(\mathbf{x})$ denoting the loss function. Exp-concavity naturally arises in many machine learning applications such as linear/logistic regression (Foster, 1991; Vovk, 1997; Foster et al., 2018), portfolio selection (Cover, 1991), linear-quadratic regulator control (Foster and Simchowitz, 2020), and so on. From a theoretical perspective, it introduces rich structures beyond convexity, allowing algorithms to exploit the local geometry of the loss landscape, often through local norms induced along the optimization trajectory. Such structures yield sharper statistical guarantees: Exp-concave losses admit algorithms with minimax-optimal regret $O(d\log T)$ (Ordentlich and Cover, 1998), an exponential improvement over the $\Omega(\sqrt{T})$ lower bound for convex losses (Abernethy et al., 2008).
Online Newton Step (ONS) (Hazan et al., 2007) is the de facto standard for OXO, which achieves an optimal regret bound of $O(d\log T)$ with only constant time and space per round. ONS exhibits remarkable simplicity, which has driven many advances in diverse optimization settings (Cutkosky and Orabona, 2018; Orabona et al., 2012; Luo et al., 2016) and in machine learning applications even beyond regret minimization, such as generalized linear bandits (Zhang et al., 2016; Jun et al., 2017; Zhang et al., 2025). The important insight of ONS is to leverage exp-concavity by maintaining a Hessian-related matrix that captures the local geometry of the optimization trajectory, which is crucial for a balance between statistical optimality and computational practicality. Indeed, all known other optimal OXO algorithms that avoid ONS-like Hessian maintenance, such as Exponential Weight Online Optimization (EWOO) (Hazan et al., 2007), incur prohibitive $O(T^{24})$ time complexity due to integrating over the domain (Bubeck et al., 2018).
However, the time-varying Hessian-related matrix in ONS necessitates a Mahalanobis projection at each round to ensure feasibility, which introduces the computational bottleneck. Specifically, the Mahalanobis projection solves the quadratic program $\Pi_{\mathcal{X}}^{M}[\mathbf{y}]=\operatorname*{arg\,min}_{\mathbf{x}\in\mathcal{X}}(\mathbf{x}-\mathbf{y})^{\mathsf{T}}M(\mathbf{x}-\mathbf{y})$ for some positive-definite and symmetric matrix $M$ and convex domain $\mathcal{X}\subseteq\mathbb{R}^{d}$. When the domain $\mathcal{X}$ is bounded, this entails $\Omega(d^{\omega})$ arithmetic operations due to matrix factorizations such as matrix square root (Golub and Van Loan, 2013).  [^1] The factorization underlying Mahalanobis projections is related to eigendecomposition, equivalent to finding roots of a $d$-degree polynomial, which is not exactly solvable by finitely many arithmetic operations when $d\geq 5$. Even for simple domains such as the unit ball and the probability simplex, Mahalanobis projection requires $\widetilde{O}(d^{\omega})$ time (see Section˜3.3) and $\widetilde{O}(d^{\omega+0.5})$ time (via interior-point methods (Nesterov and Nemirovskii, 1994)), respectively. Thus, the total runtime of ONS can reach $\widetilde{O}(d^{\omega}T)$, particularly when iterates frequently oscillate near the domain boundary. Although the matrix-multiplication exponent satisfies $\omega<2.3714$ theoretically (Alman et al., 2025), linear algebra libraries typically operate with $\omega=3$, resulting in a runtime of $\widetilde{O}(d^{3}T)$ in practice. In contrast, for convex and strongly convex online optimization, Online Gradient Descent (OGD) (Zinkevich, 2003) requires only Euclidean projections and achieves minimax-optimal regret with a runtime of $O(dT)$ for these simple domains.
[…]
Related works.
The quest for computationally efficient OXO algorithms bifurcates into two main research lines. The first, which we focus on in this paper, aims to minimize runtime while preserving optimal regret. The most relevant work is by Mhammedi and Gatmiry (2023), who proposed the OQNS (Online Quasi-Newton Steps) algorithm that attains the optimal regret $O(d\log T)$ with runtime $O(d^{2}T\log T+d^{\omega}\sqrt{T\log T})$. They employ a log-barrier to eliminate Mahalanobis projections, transferring the computational burden to Hessian-inverse evaluations under log-barrier regularization. However, their method breaks from canonical algorithmic frameworks such as online mirror descent (OMD) and Follow-the-Regularized-Leader (FTRL) (Hazan, 2016), making it difficult to accommodate various local norms and thereby limiting applicability beyond OXO. A detailed comparison of ONS and our method with OQNS (Mhammedi and Gatmiry, 2023) is deferred to Section˜5.
The second research line aims to reduce regret within a time or space budget. Prominent in this line are works using matrix sketching, such as (Luo et al., 2016). Their method achieves both runtime and working memory linear in $d$, albeit under additional assumptions on the loss functions and domains. Other works in this line include projection-free methods that trade statistical optimality for computational gains, leading to suboptimal regret bounds like $O(T^{2/3})$ (Garber and Kretzu, 2023; Wan et al., 2022). The trade-off between statistical performance and computational cost is a foundational theme in machine learning. Our work advances this direction from an online optimization perspective, achieving a balance among these competing goals while preserving the flexibility to exploit problems’ local geometry. Our focus on asymptotic online behaviors contrasts with progress on offline optimization problems (Servedio, 1999; Raz, 2016; Lyu et al., 2023; Sharan et al., 2019; Peng and Rubinstein, 2023).
Table 1: Algorithmic upper bounds on regret and total runtime of OXO algorithms with respect to $d$ and $T$ over a simple domain of the unit ball. “ONS-like” indicates that the method can be integrated into other settings where ONS is a key tool, including gradient-norm adaptivity, parametric bandits, and memory-efficient OXO.
Algorithm Regret Total Runtime ONS-like
OGD (Zinkevich, 2003) $\sqrt{T}$ dTdT
ONS (Hazan et al., 2007) $d\log{T}$ $d^{\omega}T\log T$
[…]
Contributions.
We propose Light Online Newton Step (LightONS), a simple ONS variant that substantially reduces the total runtime while achieving optimal regret. Our method preserves the elegant structure of ONS and thus inherits its applicability across various scenarios. Our contributions are three-fold:
  • An optimal and efficient algorithm for OXO. As summarized in Table˜1, LightONS attains the best-known total runtime $O(d^{2}T+d^{\omega}\sqrt{T\log T})$ to achieve the minimax-optimal regret $O(d\log T)$ for OXO. Crucially, LightONS matches ONS’s dependence on all problem parameters in OXO ($T$, $d$, $D$, $G$, $\alpha$), whereas the prior method (Mhammedi and Gatmiry, 2023) suffers from large multiplicative constants. Empirical validations in Section˜3.4 corroborate the theoretical superiority of our method.
  • An optimal and efficient algorithm for SXO. With the online-to-batch conversion, LightONS yields, up to poly-logarithmic factors, the optimal sample complexity $T=\widetilde{O}(d/\epsilon)$ to achieve excess risk $\epsilon$, thus reducing the total runtime to $\widetilde{O}(d^{3}/\epsilon)$. Our result answers a COLT’13 open problem posed by Koren (2013). In Section˜4.3, we provide evidence that the runtime $\widetilde{O}(d^{3}/\epsilon)$ is unlikely to be improved in practice.
  • Applicability across various scenarios. LightONS preserves the online mirror descent (OMD) framework of ONS and inherits ONS’s structural flexibility, especially in accommodating various local norms. In Section˜5, we demonstrate plugging LightONS into scenarios which extend beyond regret minimization settings, including gradient-norm adaptive regret, parametric stochastic bandits, and memory-efficient online learning. In contrast, the prior method (Mhammedi and Gatmiry, 2023), tailored for OXO, lacks the flexibility to fit for these scenarios.
Outline.
The remainder of this paper is organized as follows. Section˜2 presents preliminaries. Section˜3 presents our method and key analytical ingredients. Section˜4 discusses the implications of LightONS to SXO, which answers a COLT’13 open problem. Section˜5 demonstrates the broad applicability of LightONS inherited from ONS. Section˜6 concludes the paper and discusses future directions. All omitted details of algorithms and proofs are deferred to the appendices.
2 Preliminaries
In this section, we introduce notations used throughout this paper, formalize the problem setting of Online eXp-concave Optimization (OXO), and review important progress.
Notations.
Let $a]_{+}=\max\left\{0,a\right\}$ be the hinge function, let $N]=\left\{1,\dots,N\right\}$ be the index set, and let $\mathcal{B}(R)=\left\{\mathbf{x}\,|\,\left\|\mathbf{x}\right\|_{2}\leq R\right\}$ be the Euclidean ball of radius $R$ centered at the origin. Let $\nabla^{2}f(\mathbf{x})^{\mathsf{T}}$ and $\nabla^{2}f(\mathbf{x})^{-1}$ denote the transpose and inverse of the Hessian matrix, and let $\left\|\mathbf{x}\right\|_{M}=\sqrt{\mathbf{x}^{\mathsf{T}}M\mathbf{x}}$ be the Mahalanobis norm induced by a positive-definite and symmetric matrix $M$. Let $\Pi_{\mathcal{X}}^{M}[\mathbf{y}]=\operatorname*{arg\,min}_{\mathbf{x}\in\mathcal{X}}\left\|\mathbf{x}-\mathbf{y}\right\|_{M}^{2}$ be the Mahalanobis projection of $\mathbf{y}$ onto a compact convex set $\mathcal{X}$, which exists and is unique (Boyd and Vandenberghe, 2004), and let $\Pi_{\mathcal{X}}[\mathbf{y}]=\Pi^{I}_{\mathcal{X}}[\mathbf{y}]=\operatorname*{arg\,min}_{\mathbf{x}\in\mathcal{X}}\left\|\mathbf{x}-\mathbf{y}\right\|_{2}^{2}$ be the Euclidean projection. For shorthand, we write $\texttt{EP}_{\mathcal{X}}$ and $\texttt{MP}_{\mathcal{X}}$ for runtime of Euclidean and Mahalanobis projection onto $\mathcal{X}$ respectively.
2.1 Problem Setting
Online Convex Optimization (OCO) unfolds as a game between a learner and an environment over $T$ rounds. At each round $t\in[T]$, the learner selects a decision $\mathbf{x}_{t}$ from a compact convex domain $\mathcal{X}\subseteq\mathbb{R}^{d}$, and the environment simultaneously reveals a convex loss function $f_{t}:\mathcal{X}\to\mathbb{R}$; Then the learner incurs a loss $f_{t}(\mathbf{x}_{t})$ and observes a gradient $\nabla f_{t}(\mathbf{x}_{t})$ for updates. The performance of the learner is measured by its regret against some comparator $\mathbf{u}\in\mathcal{X}$, which is defined as:
$\textsc{Reg}_{T}(\mathbf{u})=\sum_{t=1}^{T}f_{t}(\mathbf{x}_{t})-\sum_{t=1}^{T}f_{t}(\mathbf{u}).$
The regularity of the domain and the loss functions is formally stated below (Hazan, 2016).
Assumption 1 (bounded domain).
The domain $\mathcal{X}\subseteq\mathbb{R}^{d}$ is compact and convex, and has a diameter of $D$, i.e., $\max_{(\mathbf{x},\mathbf{y})\in\mathcal{X}^{2}}\left\|\mathbf{x}-\mathbf{y}\right\|_{2}\leq D$. Without loss of generality, let the domain $\mathcal{X}$ be centered at the origin, i.e., $\max_{\mathbf{x}\in\mathcal{X}}\left\|\mathbf{x}\right\|_{2}\leq D/2$.
Assumption 2 (bounded gradient).
For any $t\in[T]$, the loss function $f_{t}:\mathcal{X}\to\mathbb{R}$ is differentiable and $G$-Lipschitz, i.e., $\max_{\mathbf{x}\in\mathcal{X}}\left\|\nabla f_{t}(\mathbf{x})\right\|_{2}\leq G$.
Exponentially concave (exp-concave) functions are defined below (Kivinen and Warmuth, 1999; Cesa-Bianchi and Lugosi, 2006).
Definition 1 (exp-concavity).
For a compact convex set $\mathcal{X}\subseteq\mathbb{R}^{d}$ and a positive real number $\alpha$, a function $f:\mathcal{X}\to\mathbb{R}$ is $\alpha$-exp-concave if and only if $g(\mathbf{x})=\exp\left(-\alpha f(\mathbf{x})\right)$ is concave.
Online eXp-concave Optimization (OXO) follows OCO with the exp-concavity of loss functions (Hazan et al., 2007; Mhammedi and Gatmiry, 2023), as stated below.
Assumption 3 (exp-concave loss).
For any $t\in[T]$, the loss function $f_{t}:\mathcal{X}\to\mathbb{R}$ is $\alpha$-exp-concave.
2.2 Important Progress
In this subsection, we review two important OXO algorithms, namely Online Newton Step (ONS) (Hazan et al., 2007) and Online Quasi-Newton Step (OQNS) (Mhammedi and Gatmiry, 2023).
2.2.1 Online Newton Step (ONS)
Online Newton Step (ONS) (Hazan et al., 2007) is summarized in Algorithm˜1. The following proposition states the theoretical guarantees of ONS.
Proposition 1 (Theorem 2 of Hazan et al. (2007)).
Under Assumptions˜1, 2 and 3, ONS (Algorithm˜1) satisfies that, for any $\mathbf{u}\in\mathcal{X}$,
$\textsc{Reg}_{T}(\mathbf{u})\leq\frac{d}{2\gamma_{0}}\log\left(1+\frac{G^{2}}{d\epsilon}T\right)+\frac{\gamma_{0}\epsilon D^{2}}{8},$ (1)
where $\gamma_{0}=\frac{1}{2}\min\left\{\frac{1}{DG},\alpha\right\}$. The total runtime of ONS is
$O\left(\left(\texttt{MP}_{\mathcal{X}}+d^{2}\right)T\right)=\widetilde{O}(d^{\omega}T).$ (2)
ONS follows the classical framework of online mirror descent (OMD) (Orabona, 2019), equipped with a time-varying Mahalanobis norm induced by the Hessian-related matrix $A_{t}$. The computational bottleneck of ONS lies in Line 6 in Algorithm˜1, the Mahalanobis projection $\mathbf{x}_{t+1}=\Pi^{A_{t}}_{\mathcal{X}}[\widehat{\mathbf{x}}_{t+1}]$. Once the decision $\widehat{\mathbf{x}}_{t+1}$ exits the domain $\mathcal{X}$, it must be projected back. Since each Mahalanobis projection requires $\Omega(d^{\omega})$ arithmetic operations and ONS projects in $O(T)$ rounds in the worst case, the crippling total runtime $\widetilde{O}(d^{\omega}T)$ emerges, which in practice evaluates to $\widetilde{O}(d^{3}T)$.
ONS and many other practical OXO algorithms are built upon a key property of exp-concavity that bridges its curvature and the Hessian-related matrix, as stated in the following lemma, which is essentially Lemma 3 of Hazan et al. (2007) with improved constants. The proof of this lemma is provided in Section˜A.2.
Lemma 1.
If a function $f:\mathcal{X}\to\mathbb{R}$ is $\alpha$-exp-concave and differentiable, then for any $(\mathbf{x},\mathbf{u})\in\mathcal{X}^{2}$, $D\geq\left\|\mathbf{x}-\mathbf{u}\right\|_{2}$, $G\geq\left\|\nabla f(\mathbf{x})\right\|_{2}$, $\gamma\leq\gamma_{0}=\frac{1}{2}\min\left\{\frac{1}{DG},\alpha\right\}$, it holds that
$f(\mathbf{x})-f(\mathbf{u})\leq\nabla f(\mathbf{x})^{\mathsf{T}}(\mathbf{x}-\mathbf{u})-\frac{\gamma}{2}\left(\nabla f(\mathbf{x})^{\mathsf{T}}(\mathbf{x}-\mathbf{u})\right)^{2},$ (3)
or equivalently,
$f(\mathbf{u})\geq f(\mathbf{x})+\nabla f(\mathbf{x})^{\mathsf{T}}(\mathbf{u}-\mathbf{x})+\frac{1}{2}\left\|\mathbf{u}-\mathbf{x}\right\|_{\gamma\nabla f(\mathbf{x})\nabla f(\mathbf{x})^{\mathsf{T}}}^{2}.$ (4)
We remark that, with this lemma, Assumption˜1 turns crucial for achieving the optimal regret $O(d\log T)$. Because this lemma always requires an explicit diameter $D$ to determine the curvature parameter $\gamma$, even if the problem is unconstrained. Otherwise, as $D$ increases to infinity, $\gamma$ shrinks to zero and the curvature of exp-concavity vanishes.
2.2.2 Online Quasi-Newton Step (OQNS)
Online Quasi-Newton Step (OQNS) (Mhammedi and Gatmiry, 2023) effectively reduces the total runtime to reach the asymptotically optimal regret $O(d\log T)$ for OXO, as stated in the following proposition.
Proposition 2 (Theorem 9 of Mhammedi and Gatmiry (2023)).
Under Assumptions˜1, 2 and 3, OQNS (Algorithm 3 of Mhammedi and Gatmiry (2023)) satisfies that, for any $\mathbf{u}\in\mathcal{X}$,
$\textsc{Reg}_{T}(\mathbf{u})\leq\frac{5d}{\overset{\scriptscriptstyle\circ}{\gamma}}\log\left(d+T\right)+\frac{11DGd}{2}\log T+3DGd,$ (5)
where $\overset{\scriptscriptstyle\circ}{\gamma}=\frac{1}{2}\min\left\{\frac{1}{2DG},\alpha\right\}$. The total runtime of OQNS is
$O\left(\left(\texttt{EP}_{\mathcal{X}}+d^{2}\log T\right)T+d^{\omega}\sqrt{T\log T}\right).$ (6)
OQNS eliminates Mahalanobis projections with a log-barrier, such as $\log(1-\left\|\mathbf{x}\right\|_{2}^{2})$ for $\mathcal{X}=\mathcal{B}(1)$, shifting the computational burden to Hessian-inverse evaluations. Key components of OQNS is illustrated below, with $\mathcal{X}=\mathcal{B}(1)$ and $G=1$ as in (Mhammedi and Gatmiry, 2023).
$\mathbf{x}_{t+1}$ $=\mathbf{x}_{t}-\textrm{Approx}\left(\nabla^{2}\Phi_{t}(\mathbf{x}_{t})^{-1}\nabla\Phi_{t}(\mathbf{x}_{t})\right),~~\text{where}$ (7a)
$\Phi_{t}(\mathbf{x})$ $\triangleq-\eta d\log(1-\left\|\mathbf{x}\right\|_{2}^{2})+\frac{d+\eta}{2}\left\|\mathbf{x}\right\|_{2}^{2}+\sum_{s=1}^{t}\left(\nabla f_{s}(\mathbf{x}_{s})^{\mathsf{T}}\mathbf{x}+\frac{\gamma}{2}\left(\nabla f_{s}(\mathbf{x}_{s})^{\mathsf{T}}(\mathbf{x}-\mathbf{x}_{s})\right)^{2}\right).$ (7b)
Evaluating the Hessian-inverse still takes $O(d^{\omega})$ time and OQNS mitigates this issue by approximating the Hessian-inverse gradient product with incremental updates. The approximation procedure, $\textrm{Approx}(\cdot)$, returns within $O(d^{2}\log T)$ time under proper conditions. OQNS controls the number of exact Hessian-inverse evaluations to $O(\sqrt{T\log T})$, leading to a total runtime of $O(d^{2}T\log T+d^{\omega}\sqrt{T\log T})$, while achieving the asymptotical optimal regret $O(d\log T)$.
However, their improvement comes at the cost of introducing large constant factors and abandoning the OMD framework of ONS. Specifically, when $\alpha\geq\frac{1}{DG}$, the leading $\log T$ term of ONS carries coefficient $DGd$ in Eq.˜1, whereas OQNS’s leading coefficient is $\frac{51}{2}DGd$ in Eq.˜5. Moreover, the log-barrier limits the applicability of OQNS, which will be dicussed in Section˜5.
Algorithm 1 ONS (Hazan et al., 2007)
0: preconditioner coefficient $\epsilon$.
1: Initialize $\gamma_{0}=\frac{1}{2}\min\left\{\frac{1}{DG},\alpha\right\}$, $A_{0}=\epsilon I$, $\mathbf{x}_{1}=\bm{0}$.
2:for $t=1,\dots,T$ do
3:  Observe $\nabla f_{t}(\mathbf{x}_{t})$.
4:  $A_{t}=A_{t-1}+\nabla f_{t}(\mathbf{x}_{t})\nabla f_{t}(\mathbf{x}_{t})^{\mathsf{T}}$.
5:  $\widehat{\mathbf{x}}_{t+1}=\mathbf{x}_{t}-\frac{1}{\gamma_{0}}A_{t}^{-1}\nabla f_{t}(\mathbf{x}_{t})$.
6:  $\mathbf{x}_{t+1}=\Pi^{A_{t}}_{\mathcal{X}}[\widehat{\mathbf{x}}_{t+1}]$.
7:end for
Algorithm 2 LightONS.Core
0: preconditioner coefficient $\epsilon$, hysteresis coefficient $k$.
1: Initialize $\gamma=\frac{1}{2}\min\left\{\frac{2}{(k+1)DG},\alpha\right\}$, $A_{0}=\epsilon I$, $\mathbf{x}_{1}=\bm{0}$.
2:for $t=1,\dots,T$ do
3:  Observe $\nabla f_{t}(\mathbf{x}_{t})$.
4:  $A_{t}=A_{t-1}+\nabla f_{t}(\mathbf{x}_{t})\nabla f_{t}(\mathbf{x}_{t})^{\mathsf{T}}$.
5:  $\widehat{\mathbf{x}}_{t+1}=\mathbf{x}_{t}-\frac{1}{\gamma}A_{t}^{-1}\nabla f_{t}(\mathbf{x}_{t})$.
6:  $\mathbf{x}_{t+1}=\begin{cases}\widehat{\mathbf{x}}_{t+1}&\text{if}~\left\|\widehat{\mathbf{x}}_{t+1}\right\|_{2}\leq kD/2\\ \Pi^{A_{t}}_{\mathcal{X}}[\widehat{\mathbf{x}}_{t+1}]&\text{otherwise}\end{cases}$.
7:end for
3 Our Algorithm: LightONS
Prior progress naturally raises a question: Can we retain the simplicity and elegance of ONS, while achieving a total runtime that is competitive with, or even superior to, the state-of-the-art? Our work is motivated by answering this question in the affirmative.
In this spirit, we present our algorithm LightONS. In Section˜3.1, to illustrate the key idea, we introduce the core algorithm that only differs from ONS by one line of code. However, the core algorithm is essentially improper learning. To address this issue, we introduce the improper-to-proper conversion in Section˜3.2, which yields the complete version of LightONS. Next, in Section˜3.3, we provide efficient implementations of LightONS. Finally, in Section˜3.4, we validate the empirical superiority of our method, which corroborates its theoretical guarantees.
3.1 Amortizing Projections with Hysteresis
The key idea of LightONS is to amortize the costly Mahalanobis projections with a hysteresis mechanism and enforce feasibility with computationally cheap Euclidean projections.
To illustrate this idea, we first introduce the core algorithm LightONS.Core in Algorithm˜2. As shown in Line 6, the proposed LightONS.Core performs a projection only when the decision moves outside an expanded domain $\widetilde{\mathcal{X}}\subseteq\mathbb{R}^{d}$ defined as
$\widetilde{\mathcal{X}}\triangleq\mathcal{B}(kD/2)=\{\mathbf{x}\in\mathbb{R}^{d}\,|\,\left\|\mathbf{x}\right\|_{2}\leq kD/2\},$ (8)
where $k>1$ is the hysteresis coefficient. Essentially, $\widetilde{\mathcal{X}}$ is the Euclidean ball of radius $kD/2$ centered at the origin, obtained by scaling by a factor of $k$ the minimal ball that contains $\mathcal{X}$. The key difference between LightONS.Core and the standard ONS algorithm (Hazan et al., 2007) is that ONS projects onto $\mathcal{X}$ immediately once the decision leaves $\mathcal{X}$, whereas LightONS.Core delays projection and continues updating outside $\mathcal{X}$, projecting back onto $\mathcal{X}$ only when the decision exits the expanded domain $\widetilde{\mathcal{X}}$.
Clearly, allowing a sufficiently large $k$ significantly improves projection efficiency, since projections are invoked less frequently. The following lemma precisely characterizes the relationship between the hysteresis coefficient $k$ and the number of Mahalanobis projections performed by the core algorithm. The proof of this lemma is provided in Section˜B.1.
Lemma 2.
Under Assumption˜1, and that for any $t\in[T]$ the loss function $f_{t}:\widetilde{\mathcal{X}}\to\mathbb{R}$ is $\alpha$-exp-concave, differentiable and $G$-Lipschitz on $\widetilde{\mathcal{X}}$, let $N$ denote the number of Mahalanobis projections in LightONS.Core (Algorithm˜2) over $T$ rounds, then
$N\leq\left\lfloor\frac{2}{(k-1)D\gamma}\sqrt{\frac{d}{\epsilon}T}\right\rfloor.$ (9)
On the other hand, because the proposed hysteresis mechanism alters the projection rule, it is essential to verify that the regret guarantee is preserved. In particular, increasing $k$ degrades the curvature parameter $\gamma$, leading to weaker regret bounds. In the extreme case $k$ approaches infinity, $\gamma$ collapses to zero, and exp-concavity degenerates to mere convexity. The following theorem establishes the relationship between the hysteresis coefficient $k$ and the regret bound of LightONS.Core. The proof of this theorem is provided in Section˜B.2.
Theorem 1.
Under Assumption˜1, and assuming that for any $t\in[T]$ the loss function $f_{t}:\widetilde{\mathcal{X}}\to\mathbb{R}$ is $\alpha$-exp-concave, differentiable and $G$-Lipschitz on $\widetilde{\mathcal{X}}$, LightONS.Core (Algorithm˜2) satisfies that, for any $\mathbf{u}\in\mathcal{X}$,
$\textsc{Reg}_{T}(\mathbf{u})\leq\frac{d}{2\gamma}\log\left(1+\frac{G^{2}}{d\epsilon}T\right)+\frac{\gamma\epsilon D^{2}}{8},$ (10)
where $\gamma=\frac{1}{2}\min\left\{\frac{2}{(k+1)DG},\alpha\right\}$. The total runtime of LightONS.Core over $T$ rounds is
$O\left(d^{2}T+(k-1)^{-1}\sqrt{dT/\epsilon}\cdot\texttt{MP}_{\mathcal{X}}\right).$ (11)
As $k$ increases, Eq.˜10 shows that the regret bound scales as $O(\frac{1}{\gamma})=O(k+1)$, whereas Eq.˜11 indicates that the total runtime spent on Mahalanobis projections decreases as $O(N)=O(\frac{1}{k-1})$. Choosing $k=2$ therefore yields a favorable regret-efficiency trade-off: the regret bound grows by at most a factor of $\frac{3}{2}$, while the number of projections is substantially reduced. This suggests that LightONS.Core already achieves a significant improvement over standard ONS in terms of computational efficiency without sacrificing regret guarantees. However, as discussed below, LightONS.Core actually suffers from an improper learning issue.
Improper Learning Issue.
However, our LightONS.Core algorithm falls in the scope of improper learning (Shalev-Shwartz and Ben-David, 2014), as the algorithm’s decisions $\mathbf{x}_{t}\in\widetilde{\mathcal{X}}\supset\mathcal{X}$ may reside beyond the domain while the comparator $\mathbf{u}\in\mathcal{X}$ is strictly constrained to the domain. The core algorithm requires additional assumptions that the Lipschitzness and exp-concavity of the loss functions extend to the enlarged domain $\widetilde{\mathcal{X}}$. Vitally, such additional assumptions suppress the theoretical performance limits of its proper counterparts. A notable illustration is online logistic regression, where an improper learner achieves regret bounds $O(d\log(GT))$ (Foster et al., 2018), while proper learners are limited to a regret lower bound $\Omega(d\mathrm{e}^{G}\log T)$ (Hazan et al., 2014).
In the next subsection, we build upon LightONS.Core to develop a learning algorithm for OXO that achieves efficient updates and optimal regret, while guaranteeing that all decisions remain within the domain $\mathcal{X}$.
3.2 The Improper-to-Proper Conversion
Based on techniques from parameter-free online learning (Cutkosky and Orabona, 2018; Cutkosky, 2020), we incorporate a domain conversion into LightONS.Core, which yields a fully proper algorithm LightONS. This incorporation is conceptually simple yet technically delicate, as it recovers the optimal regret of ONS while preserving the efficiency gains of LightONS.Core.
We present the complete version of LightONS in Algorithm˜3. The domain conversion works by constructing surrogate loss functions. The core algorithm inside LightONS interacts with the surrogate loss $g_{t}$ (instead of the true loss $f_{t}$) and generates the surrogate decision $\mathbf{y}_{t}\in\widetilde{\mathcal{X}}$ (instead of the true decision $\mathbf{x}_{t}\in\mathcal{X}$). The true decision $\mathbf{x}_{t}$ is then obtained as the Euclidean projection of $\mathbf{y}_{t}$ onto the domain $\mathcal{X}$. For LightONS, any surrogate loss satisfying the following condition accomplishes this conversion.
Condition 1.
For some $c_{f}\geq 1$ and $c_{g}\geq 1$, the surrogate loss function $g_{t}:\mathbb{R}^{d}\to\mathbb{R}$ satisfies that, for any $\mathbf{u}\in\mathcal{X}$, $\left\|\nabla g_{t}(\mathbf{y}_{t})\right\|_{2}\leq c_{g}G$ and $\nabla f_{t}(\mathbf{x}_{t})^{\mathsf{T}}(\mathbf{x}_{t}-\mathbf{u})\leq c_{f}\nabla g_{t}(\mathbf{y}_{t})^{\mathsf{T}}(\mathbf{y}_{t}-\mathbf{u})$.
We note that two types of such conversions have been proposed in the literature. The first, by Cutkosky and Orabona (2018), satisfies the condition with $c_{f}=2$ and $c_{g}=1$. Later, an improved conversion was proposed with $c_{f}=1$ and $c_{g}=1$ (Cutkosky, 2020). In this work, we use the latter, as it incurs smaller constants and leads to tighter regret bounds.
Lemma 3 (Theorem 2 of Cutkosky (2020)).
Under Assumptions˜1 and 2, let $\mathbf{x}_{t}=\Pi_{\mathcal{X}}[\mathbf{y}_{t}]$, the surrogate loss function $g_{t}:\mathbb{R}^{d}\to\mathbb{R}$ and its subgradient at $\mathbf{y}_{t}$ are defined as follows:
$g_{t}(\mathbf{y})$ $\triangleq\nabla f_{t}(\mathbf{x}_{t})^{\mathsf{T}}\mathbf{y}+\frac{[-\nabla f_{t}(\mathbf{x}_{t})^{\mathsf{T}}(\mathbf{y}_{t}-\mathbf{x}_{t})]_{+}}{\left\|\mathbf{y}_{t}-\mathbf{x}_{t}\right\|_{2}}\left\|\mathbf{y}-\Pi_{\mathcal{X}}[\mathbf{y}]\right\|_{2},$ (12a)
$\nabla g_{t}(\mathbf{y}_{t})$ $=\nabla f_{t}(\mathbf{x}_{t})+\frac{[-\nabla f_{t}(\mathbf{x}_{t})^{\mathsf{T}}(\mathbf{y}_{t}-\mathbf{x}_{t})]_{+}}{\left\|\mathbf{y}_{t}-\mathbf{x}_{t}\right\|_{2}^{2}}\left(\mathbf{y}_{t}-\mathbf{x}_{t}\right),$ (12b)
then for any $\mathbf{u}\in\mathcal{X}$, $\left\|\nabla g_{t}(\mathbf{y}_{t})\right\|_{2}\leq\left\|\nabla f_{t}(\mathbf{x}_{t})\right\|_{2}\leq G$, and $\nabla f_{t}(\mathbf{x}_{t})^{\mathsf{T}}(\mathbf{x}_{t}-\mathbf{u})\leq\nabla g_{t}(\mathbf{y}_{t})^{\mathsf{T}}(\mathbf{y}_{t}-\mathbf{u})$.
Algorithm 3 LightONS
0: preconditioner coefficient $\epsilon$, hysteresis coefficient $k$.
1: Initialize $\gamma^{\prime}=\frac{1}{2}\min\left\{\frac{1}{c_{f}c_{g}DG},\frac{4}{c_{f}c_{g}(k+1)DG},\alpha\right\}$; $A_{0}=\epsilon I$; $\mathbf{x}_{1}=\mathbf{y}_{1}=\bm{0}$.
2:for $t=1,\dots,T$ do
3:  Observe $\nabla f_{t}(\mathbf{x}_{t})$.
4:  Construct $\nabla g_{t}(\mathbf{y}_{t})$, where $g_{t}$ satisfies Condition˜1.
5:  $A_{t}=A_{t-1}+\nabla g_{t}(\mathbf{y}_{t})\nabla g_{t}(\mathbf{y}_{t})^{\mathsf{T}}$.
6:  $\widehat{\mathbf{y}}_{t+1}=\mathbf{y}_{t}-\frac{1}{\gamma^{\prime}}A_{t}^{-1}\nabla g_{t}(\mathbf{y}_{t})$.
7:  $\mathbf{y}_{t+1}=\begin{cases}\widehat{\mathbf{y}}_{t+1}&\text{if}~\left\|\widehat{\mathbf{y}}_{t+1}\right\|_{2}\leq kD/2\\ \Pi^{A_{t}}_{\mathcal{B}(D/2)}[\widehat{\mathbf{y}}_{t+1}]&\text{otherwise}\end{cases}$.
8:  $\mathbf{x}_{t+1}=\Pi_{\mathcal{X}}[\mathbf{y}_{t+1}]$.
9:end for
We remark that the computational overhead of this conversion is negligible. Constructing the surrogate gradient $\nabla g_{t}(\mathbf{y}_{t})$ as Eq.˜12b takes only $O(d)$ arithmetic operations per round apart from the Euclidean projection, veiled by the $O(d^{2})$ cost of updating and inverting the Hessin-related matrix $A_{t}$. It is worth noting that several prior works have also employed this domain-conversion technique to address projection-related issues, but for different purposes, such as in non-stationary online learning (Zhao et al., 2025) and universal online learning (Yang et al., 2024).
One caveat is that the surrogate loss function in Lemma˜3 is not exp-concave. Fortunately, it can still exploit the exp-concavity of the original loss function $f_{t}$ using the following lemma, which shows that the surrogate loss $g_{t}$ exhibits a curvature property similar to Lemma˜1. The proof of this lemma is provided in Section˜B.3.
Lemma 4.
Under Assumptions˜1, 2, 3 and 1, for any $\mathbf{u}\in\mathcal{X}$,
$f_{t}(\mathbf{x}_{t})-f_{t}(\mathbf{u})$ $\leq\nabla f_{t}(\mathbf{x}_{t})^{\mathsf{T}}(\mathbf{x}_{t}-\mathbf{u})-\frac{\gamma_{0}}{2}\left(\nabla f_{t}(\mathbf{x}_{t})^{\mathsf{T}}(\mathbf{x}_{t}-\mathbf{u})\right)^{2}$ (13)
$\leq\nabla g_{t}(\mathbf{y}_{t})^{\mathsf{T}}(\mathbf{y}_{t}-\mathbf{u})-\frac{\gamma^{\prime}}{2}\left(\nabla g_{t}(\mathbf{y}_{t})^{\mathsf{T}}(\mathbf{y}_{t}-\mathbf{u})\right)^{2}$
where $\gamma^{\prime}=\frac{1}{2}\min\left\{\frac{1}{c_{f}c_{g}DG},\frac{4}{(k+1)c_{f}c_{g}DG},\alpha\right\}$ as in LightONS (Algorithm˜3).
Lemma˜4 shows that, when the hysteresis coefficient $k$ is small (specifically $k\leq 3$), the curvature parameter $\gamma^{\prime}$ is unimpaired relative to $\gamma_{0}$ in Lemma˜1, with the domain conversion in Lemma˜3 ensuring $c_{f}=c_{g}=1$.
The following theorem establishes the optimality and efficiency of our proposed LightONS. The proof of this theorem is provided in Section˜B.4.
Theorem 2.
Under Assumptions˜1, 2 and 3, LightONS (Algorithm˜3) satisfies that, for any $\mathbf{u}\in\mathcal{X}$,
$\textsc{Reg}_{T}(\mathbf{u})\leq\frac{d}{2\gamma^{\prime}}\log\left(1+\frac{c_{g}^{2}G^{2}}{d\epsilon}T\right)+\frac{\gamma^{\prime}\epsilon D^{2}}{8},$ (14)
where $\gamma^{\prime}=\frac{1}{2}\min\left\{\frac{1}{c_{f}c_{g}DG},\frac{4}{c_{f}c_{g}(k+1)DG},\alpha\right\}$. The total runtime of LightONS over $T$ rounds is
$O\left(\left(\texttt{EP}_{\mathcal{X}}+d^{2}\right)T+(k-1)^{-1}d^{\omega+0.5}\sqrt{T/\epsilon}\log T\right).$ (15)
The preceding theorem immediately yields the following corollary.
Corollary 1.
Under Assumptions˜1, 2 and 3, LightONS (Algorithm˜3) with $k=2$ and the surrogate loss function from Lemma˜3 satisfies that, for any $\mathbf{u}\in\mathcal{X}$,
$\textsc{Reg}_{T}(\mathbf{u})\leq\frac{d}{2\gamma_{0}}\log\left(1+\frac{G^{2}}{d\epsilon}T\right)+\frac{\gamma_{0}\epsilon D^{2}}{8},$ (16)
where $\gamma_{0}=\frac{1}{2}\min\left\{\frac{1}{DG},\alpha\right\}$. Furthermore, with $\epsilon=d\log T$, the total runtime of LightONS over $T$ rounds is:
$O\left(\left(\texttt{EP}_{\mathcal{X}}+d^{2}\right)T+d^{\omega}\sqrt{T\log T}\right).$ (17)
Remark 1.
We note that the domain conversion slightly worsens the constants, as seen in the second inequality of Eq.˜13. Nonetheless, this degradation is dominated by the problem-dependent parameters of OXO, namely $T$, $d$, $D$, $G$, $\alpha$. In their dependence on these parameters, LightONS’s regret in Eq.˜16 exactly matches ONS’s regret in Eq.˜1.
LightONSpreserves the infrequent-projection property of LightONS.Core, yielding a total runtime better than both ONS and OQNS. More importantly, LightONS retains the flexible mirror-descent structure of ONS. In fact, LightONS differs from ONS in two key aspects: (i) the projection-hysteresis mechanism (introduced in LightONS.Core to enhance efficiency) and (ii) the domain conversion (introduced here to ensure proper learning), both of which are largely orthogonal to the mirror-descent update in the ONS algorithm. Consequently, LightONS applies to a wide range of scenarios where ONS is essential (particularly when its mirror-descent update plays a critical role), such as gradient-norm adaptive regret, parametric stochastic bandits, and memory-efficient online learning. We illustrate these applications in Section˜5.
Algorithm 4 FastProj onto Euclidean ball $\mathcal{B}(R)$ with Mahalanobis norm $\left\|\cdot\right\|_{A}$
0: point $\mathbf{u}\notin\mathcal{B}(R)$, error tolerance $\zeta$, range of AA’s eigenvalues $\underline{\lambda},\overline{\lambda}]$.
0: approximate Mahalanobis projection $\mathbf{v}\approx\Pi^{A}_{\mathcal{B}(R)}[\mathbf{u}]$.
1: (Choice 1.) $\mathbf{p}=A\mathbf{u}$, let $\rho(\mu)=\|\left(A+\mu I\right)^{-1}\mathbf{p}\|_{2}^{2}-R^{2}$.
2: (Choice 2.) Tridiagonalize $QCQ^{\mathsf{T}}=A$, $\mathbf{q}=CQ^{\mathsf{T}}\mathbf{u}$, let $\rho(\mu)=\|\left(C+\mu I\right)^{-1}\mathbf{q}\|_{2}^{2}-R^{2}$.
[…]
4:for $i=1,\dots,n$ do
[…]
6:end for
[…]
3.3 Numerical Implementation
The overall efficiency of LightONS hinges on two key operations: the matrix inversion $A_{t}^{-1}$ and the (infrequent) Mahalanobis projection $\Pi^{A_{t}}_{\mathcal{B}(D/2)}[\widehat{\mathbf{y}}_{t+1}]$. In this subsection, we detail efficient numerical approaches for the two operations respectively.
The per-round update of the matrix $A_{t}$ is a rank-one update. Instead inverting from scratch at a cost of $O(d^{\omega})$, we can update $A_{t-1}^{-1}$ to obtain $A_{t}^{-1}$ in only $O(d^{2})$ time with the Sherman-Morrison-Woodbury formula. The following equation ensures $V_{t}=A_{t}^{-1}$ for any $t\in[T]$.
$V_{0}=\frac{1}{\epsilon}I;\quad V_{t+1}=V_{t}-\frac{1}{\left\|\nabla g_{t}(\mathbf{y}_{t})\right\|_{V_{t}^{-1}}^{2}}V_{t}\nabla g_{t}(\mathbf{y}_{t})\nabla g_{t}(\mathbf{y}_{t})^{\mathsf{T}}V_{t}.$ (18)
In LightONS, all Mahalanobis projections are onto the Euclidean ball $\mathcal{B}(D/2)$ rather than the potentially complex domain $\mathcal{X}$. This geometry enables customized numerical approaches faster than generic solvers (Lee et al., 2015; Jiang et al., 2020). Specifically, the dual problem of the Mahalanobis projection reduces to a one-dimensional root-finding problem (Exercise 4.22 of Boyd and Vandenberghe (2004)), solvable via bisection. We propose FastProj in Algorithm˜4, with its theoretical guarantees given in the lemma below. The proof of this theorem is provided in Section˜B.5.
Theorem 3.
Let $\mathbf{v}$ denote the output of FastProj (Algorithm˜4), and $\mathbf{v}^{*}=\Pi^{A}_{\mathcal{B}(R)}[\mathbf{u}]$ denote the exact Mahalanobis projection, then $\mathbf{v}\in\mathcal{B}(R)$ and $\left\|\mathbf{v}-\mathbf{v}^{*}\right\|_{2}\leq\zeta$.
With choice 1, the total runtime is $O(d^{\omega}n)$, and with choice 2, the total runtime is $O(d^{3}+dn)$. The number of bisection iterations $n=O(\log(\frac{1}{\zeta}\frac{~\overline{\lambda}~}{~\underline{\lambda}~}))$. In the context of LightONS (Algorithm˜3), $\zeta_{t}=O(1/t^{2})$, $\underline{\lambda}_{t}=\epsilon$ and $\overline{\lambda}_{t}=\epsilon+c_{g}^{2}G^{2}t$, thus $n_{t}=O(\log t)$.
We present two choices for different purposes. Choice 1 attains superior theoretical dependence on $d$ by exploiting fast matrix-multiplication with exponent $\omega<2.3714$. Choice 2 offers stronger practical performance via tridiagonalization (Parlett, 1998; Golub and Van Loan, 2013), which is preferable when adopting the practical setting $\omega=3$.
The accumulated truncation error contributes only a negligible additive constant to the regret of LightONS. Full details are deferred to Section˜B.4.
(a) Instantaneous regret.
(b) Regret.
(c) Norm.
(d) Instantaneous regret.
(e) Regret.
(f) Norm.
Figure 1: Experimental results with $\mathcal{X}=\mathcal{B}(1)$, $d=10$, $T=10^{4}$. The first row shows linear regression with $G=\frac{1}{10}$, $\alpha=5$; The second row shows logistic regression with $G=\frac{1}{10}$, $\alpha=\mathrm{e}^{-1/5}$.
3.4 Empirical Validation
We conduct numerical experiments to validate the theoretical guarantees of LightONS, especially its negligible statistical gap from ONS and its non-asymptotic statistical advantages over OQNS. Three algorithms, namely LightONS, ONS, and OQNS, are configured with their theoretically optimal parameters. The implementation of ONS and LightONS follows Algorithms˜1 and 3, respectively; While OQNS is implemented in accordance with Algorithm 3 of Mhammedi and Gatmiry (2023). We evaluate on two fundamental tasks: linear regression and logistic regression. The loss functions are defined as:
$\ell^{\text{linear}}_{t}(\mathbf{w})=\frac{1}{2}\left(\mathbf{x}_{t}^{\mathsf{T}}\mathbf{w}+y_{t}\right)^{2},\quad\ell^{\text{logistic}}_{t}(\mathbf{w})=\log\left(1+\exp\left(\mathbf{x}_{t}^{\mathsf{T}}\mathbf{w}\right)\right),\quad(\mathbf{x}_{t},y_{t})\sim\left|\mathcal{N}(\bm{0},I)\right|.$
Our logistic regression setup slightly differs from the standard formulation, as we omit the binary label $y_{t}\in\left\{0,1\right\}$ for simplicity, which does not affect exp-concavity. Each entry of $\mathbf{x}_{t}$ and $y_{t}$ are i.i.d. sampled from a folded Gaussian distribution, i.e., the absolute value of a standard Gaussian random variable. All results are averaged over 5 independent runs with the same seeds. In Figure˜1, the averaged performance is shown as a dark line, while individual runs are in transparent lines.
Our experiments confirm that computational gains of LightONS incur negligible loss in statistical performance. As shown in regret plots, curves of LightONS (red dotted) overlap with those of ONS (green solid). This provides strong empirical evidence that LightONS retains the sharp regret constants of the ONS. In contrast, OQNS (blue dashed) shows noticeably higher regret and slower convergence in both regret and decision-norm plots, indicating that its theoretical bounds may carry greater constants.
The key advantage of LightONS, its computational efficiency, is evident in decisions-norm plots. After an initial convergence phase (e.g., in Figure˜1(f), the first $10^{2}$ rounds), both ONS and LightONS stabilize near the optimal solution. Then ONS frequently pushes decisions outside the domain, triggering costly Mahalanobis projections. In contrast, LightONS consistently outputs decisions with $\left\|\mathbf{y}_{t}\right\|\leq kD/2=2$ and avoids Mahalanobis projections.
4 Answering a COLT’13 Open Problem
In this section, we highlight the broader impact of our method by extending it to a related setting. Via the online-to-batch conversion, our method applies to Stochastic eXp-concave Optimization (SXO), where the optimal OXO regret translates into the (near) optimal SXO sample complexity while substantially reducing the computational cost. This extension answers a COLT’13 open problem posed by Koren (2013), demonstrating our method’s significance beyond OXO.
4.1 Restatement of the Open Problem
SXO seeks to minimize an exp-concave function $F:\mathcal{X}\to\mathbb{R}$ over a convex domain $\mathcal{X}\subseteq\mathbb{R}^{d}$, where the learner has access to $F$ only through some stochastic oracle. In this section, we consider the stochastic gradient oracle, which returns unbiased estimates of the gradient of $F$, and we are interested in the sample complexity and total runtime required to find an $\epsilon$-optimal solution $\mathbf{x}_{\epsilon}$, i.e.,
$F(\mathbf{x}_{\epsilon})-\min_{\mathbf{x}\in\mathcal{X}}F(\mathbf{x})\leq\epsilon,\quad\text{where}~F(\mathbf{x})=\mathop{\mathbb{E}}_{\xi\sim\mathcal{D}_{\xi}}\left[f(\mathbf{x};\xi)\right].$
The regularity of the stochastic functions is formally stated below (Koren, 2013).
Assumption 4.
The loss function $F:\mathcal{X}\to\mathbb{R}$ is the expectation of a random function $f:\mathcal{X}\times\Xi\to\mathbb{R}$ over an unknown distribution $\mathcal{D}_{\xi}$ on $\Xi$, i.e., $F(\mathbf{x})=\mathop{\mathbb{E}}_{\xi\sim\mathcal{D}_{\xi}}\left[f(\mathbf{x};\xi)\right]$. For any $\xi\in\Xi$, the stochastic loss function $f(\cdot;\xi)$ is $\alpha$-exp-concave, differentiable and $G$-Lipschitz over $\mathcal{X}$. For any query point $\mathbf{x}\in\mathcal{X}$, the stochastic gradient oracle returns $\nabla f(\mathbf{x};\xi)$ with $\xi$ i.i.d. drawn from $\mathcal{D}_{\xi}$.
[…]
Open problem (Koren, 2013).
Under Assumption˜4 and $\mathcal{X}=\mathcal{B}(1)$, (a) Is it possible to find an SXO algorithm that attains the sample complexity of $\widetilde{O}(d/\epsilon)$ with only linear-in-$d$ runtime per iteration, i.e., $\widetilde{O}(d^{2}/\epsilon)$ runtime overall? (b) Is it possible to perform any better than $\widetilde{O}(d^{4}/\epsilon)$ runtime overall?
The first part of the open problem remains open. In particular, Mahdavi et al. (2015) shows an information-theoretic sample complexity lower bound of $\Omega(d/\epsilon)$ for SXO, which implies a runtime lower bound of $\Omega(d^{2}/\epsilon)$ since each gradient query costs $\Omega(d)$ time.
Section˜4.2 answers the second part in the affirmative by combining LightONS with the online-to-batch conversion, obtaining total runtime $\widetilde{O}(d^{3}/\epsilon)$ while maintaining the optimal $\widetilde{O}(d/\epsilon)$ sample complexity. In Section˜4.3, we present evidence that this $\widetilde{O}(d^{3}/\epsilon)$ runtime is likely unimprovable. If confirmed, it would refute the first part of the open problem and fully settle it.
4.2 Answering the Open Problem with LightONS
By applying the online-to-batch conversion for exp-concave functions from (Mehta, 2017), LightONS provides both high-probability and in-expectation convergence rate guarantees for SXO, as stated in the following theorem. The proof of the this theorem is provided in Section˜C.1.
Theorem 4.
Under Assumption˜4 and $\mathcal{X}=\mathcal{B}(1)$, let LightONS (Algorithm˜3) run for $T$ rounds with $T$ queries to the stochastic gradient oracle $\left\{\nabla f(\mathbf{x}_{t};\xi_{t})\right\}_{t=1}^{T}$, where $\left\{\mathbf{x}_{t}\right\}_{t=1}^{T}$ are the online decisions of LightONS, and let the aggregated solution be $\bar{\mathbf{x}}_{T}=\frac{1}{T}\sum_{t=1}^{T}\mathbf{x}_{t}$, for any $\delta\in(0,1)$, then the aggregated solution satisfies that, with probability at least $1-\delta$, with $T=\Theta\left(\frac{d}{\epsilon}\log\frac{d}{\epsilon}\log\frac{1}{\delta}\right)$,
$F(\bar{\mathbf{x}}_{T})-\min_{\mathbf{x}\in\mathcal{X}}F(\mathbf{x})\leq\frac{1}{T}\left(\textsc{Reg}_{T}+4\sqrt{\frac{\textsc{Reg}_{T}}{2\gamma_{0}}\left(\log\frac{4\log T}{\delta}\right)}+\frac{8}{\gamma_{0}}\log\frac{4\log T}{\delta}\right)=O(\epsilon),$ (19)
where $\textsc{Reg}_{T}\leq\frac{d}{2\gamma_{0}}\log\left(1+\frac{G^{2}}{d\epsilon}T\right)+\frac{\gamma_{0}\epsilon D^{2}}{8}$ as in Theorem˜2. The total runtime of the algorithm is $O\left(\frac{d^{3}}{\epsilon}\log\frac{d}{\epsilon}\log\frac{1}{\delta}+\frac{d^{3.5}}{\sqrt{\epsilon}}\log\frac{d}{\epsilon}\log\frac{1}{\delta}\right)=\widetilde{O}\left(\frac{d^{3}}{\epsilon}\right)$. With $\frac{1}{\delta}=T^{\prime}=\Theta\left(\frac{d}{\epsilon}\log\frac{d}{\epsilon}\right)$, in expectation,
$\mathop{\mathbb{E}}\left[F(\bar{\mathbf{x}}_{T^{\prime}})-\min_{\mathbf{x}\in\mathcal{X}}F(\mathbf{x})\right]\leq O(\epsilon).$ (20)
The total runtime of the algorithm is $O\left(\frac{d^{3}}{\epsilon}\log\frac{d}{\epsilon}+\frac{d^{3.5}}{\sqrt{\epsilon}}\log\frac{d}{\epsilon}\right)=\widetilde{O}\left(\frac{d^{3}}{\epsilon}\right)$.
4.3 Discussions on SXO
The LightONS-based SXO method of Theorem˜4 matches the best known SXO runtime up to polylogarithmic factors. An online-to-batch conversion of OQNS yields the same asymptotical runtime $\widetilde{O}(d^{3}/\epsilon)$, but, as discussed in Section˜3, its regret bound carries large constants, which slows convergence in stochastic settings. Our experiments in Section˜3.4, conducted in SXO settings, confirm this difference.
We conjecture that no SXO algorithm can asymptotically beat total runtime $\widetilde{O}(d^{3}/\epsilon)$. Two observations support this view:
  • Fast matrix-multiplication. Linear regression with random design, a special case of SXO, reduces to solving a linear system with $d$ variables and $\widetilde{O}(d/\epsilon)$ equations. Although fast matrix-multiplication accelerates this to $\widetilde{O}(d^{\omega}/\epsilon)$ time (Ibarra et al., 1982), such techniques do not plausibly extend beyond linear models. In practice, the runtime effectively reverts to $\widetilde{O}(d^{3}/\epsilon)$.
  • Offline convex optimization. SXO methods based on Empirical Risk Minimization (ERM) (Koren and Levy, 2015; Mehta, 2017) likewise fail to surpass the $\widetilde{O}(d^{3}/\epsilon)$ barrier even with cutting-edge offline solvers (Lee et al., 2015; Jiang et al., 2020), as detailed in Section˜C.2. Moreover, such methods require $\Omega(d^{2}+d/\epsilon)$ working memory to store all samples and well-roundedness of the domain, whereas the OXO-based algorithms typically require only $O(d^{2})$ memory and no such geometric assumptions.
5 Applications to Various Problems
ONS radiates influence well beyond the classical settings of OXO and SXO. By preserving the elegant algorithmic structure of ONS, LightONS can be seamlessly integrated into various applications where ONS serves as a computational core to substantially improve efficiency. In this section, we elaborate on three representative applications of ONS where LightONS fit in without compromising statistical advantages or requiring much new analytical efforts.
  1. (i)
    Gradient-norm adaptivity. ONS achieves a regret bound scaling with the accumulated squared gradient norms, which further implies small-loss adaptivity in OXO and comparator-norm adaptivity in OCO.
  2. (ii)
    Parametric stochastic bandits. Instead of directly minimizing the regret, ONS serves as the parameter estimator to handle customized local norms that is crucial for the exploration-exploitation trade-off.
  3. (iii)
    Memory-efficient OXO. When the online gradients exhibit low-rank structure, combining ONS with matrix sketching can achieve linear-in-$d$ runtime and working memory.
We note that in these scenarios, OQNS can hardly be applied due to its significant algorithmic deviations from disciplined online learning frameworks, especially its log-barrier regularization.
5.1 Gradient-Norm Adaptivity
In this subsection, we target OXO regret bounds that scale with the accumulated squared gradient norms $G_{T}$ instead of the time horizon $T$, i.e.,
$\textsc{Reg}_{T}(\mathbf{u})=O\left(d\log G_{T}\right),\quad\text{where}~G_{T}\triangleq\sum_{t=1}^{T}\left\|\nabla f_{t}(\mathbf{x}_{t})\right\|_{2}^{2}.$ (21)
In benign environments, $G_{T}$ can be $O(1)$, for example, when the loss is nearly fixed, decisions converge quickly, yielding regret far below worst-case bounds that scale with $T$. Yet, since $G_{T}\leq G^{2}T$, the gradient-norm adaptive bound safeguards the minimax-optimality in $T$. Prior works have leveraged Eq.˜21 to achieve such stronger types of adaptivity, including small-loss adaptivity for OXO with smoothness and comparator-norm adaptivity for unbounded OCO, as discussed below.
OXO with smoothness.
Orabona et al. (2012) demonstrate that ONS directly implies the desired regret bound in Eq.˜21 for OXO. Under smoothness assumptions, the gradient-norm adaptivity can further be improved to small-loss adaptivity (Srebro et al., 2010; Zhao et al., 2020), where the regret scales with the cumulative loss of the best comparator in hindsight, regardless of online decisions. This is formalized as follows:
$\textsc{Reg}_{T}=O\left(d\log L_{T}\right),\quad\text{where}~L_{T}\triangleq\min_{\mathbf{u}\in\mathcal{X}}\sum_{t=1}^{T}\left(f_{t}(\mathbf{u})-\min_{\mathbf{x}\in\mathcal{X}}f_{t}(\mathbf{x})\right),$ (22)
when the loss function $f_{t}:\mathcal{X}\to\mathbb{R}$ is $H$-smooth, i.e., for any $t\in[T]$ and any pair $(\mathbf{x},\mathbf{y})\in\mathcal{X}^{2}$, $\left\|\nabla f(\mathbf{x})-\nabla f(\mathbf{y})\right\|_{2}\leq H\left\|\mathbf{x}-\mathbf{y}\right\|_{2}$.
Unbounded OCO.
Cutkosky and Orabona (2018) extend the gradient-norm adaptive OXO regret in Eq.˜21 to a comparator-norm adaptive OCO regret, i.e.,
$\textsc{Reg}_{T}(\mathbf{u})=\widetilde{O}\left(\left\|\mathbf{u}\right\|_{2}\sqrt{dG_{T}}\right),\quad\text{for any (unbounded) comparator}~\mathbf{u}\in\mathbb{R}^{d},$ (23)
without any prior knowledge of $\mathbf{u}$. Such guarantees for unbounded comparators are unachievable via classical OCO algorithms that explicitly depend on the domain diameter $D$, for example, OGD demands an explicit $D$ to obtain $O(DG\sqrt{T})$ regret (Zinkevich, 2003; Abernethy et al., 2008).
Nevertheless, both (Orabona et al., 2012; Cutkosky and Orabona, 2018) incur the worst-case $\widetilde{O}(d^{\omega}T)$ runtime bottleneck of ONS. Below, we show that LightONS can replace ONS in these algorithms to substantially improve efficiency without compromising statistical guarantees.
Improvements by LightONS.
LightONSrecovers ONS’s gradient-norm adaptive bound in Eq.˜21, thereby leading to small-loss adaptivity for OXO under smoothness assumptions, as well as comparator-norm adaptivity for unbounded OCO. We state the theoretical guarantees based on LightONS and omit those of the original ONS-based methods as they are identical except for the runtime. The proofs are deferred to Section˜D.1.
Theorem 5 (LightONS’s improvement for OXO with smoothness).
Under Assumptions˜1, 2 and 3, and that $f_{t}$ is $H$-smooth for any $t\in[T]$, LightONS (Algorithm˜3) satisfies that,
$\textsc{Reg}_{T}\leq\frac{d}{2\gamma_{0}}\log\left(\frac{8H}{d\epsilon}L_{T}+\frac{4H}{\gamma_{0}\epsilon}\log\frac{4H}{\mathrm{e}\gamma_{0}\epsilon}+\frac{\gamma_{0}D^{2}H}{d}+2\right)+\frac{\gamma_{0}\epsilon D^{2}}{8},$
where $L_{T}$ is defined in Eq.˜22. The total runtime of the algorithm is $O((\texttt{EP}_{\mathcal{X}}+d^{2})T+d^{\omega}\sqrt{T}\log T)$.
Theorem 6 (LightONS’s improvement for unbounded OCO).
If the loss function $f_{t}:\mathbb{R}^{d}\to\mathbb{R}$ is convex and $1$-Lipschitz for any $t\in[T]$, then there exists an algorithm satisfying that, for any $\mathbf{u}\in\mathbb{R}^{d}$,
$\textsc{Reg}_{T}(\mathbf{u})\leq\widetilde{O}\left(\sqrt{d\sum_{t=1}^{T}\left(\nabla f_{t}(\mathbf{x}_{t})^{\mathsf{T}}\mathbf{u}\right)^{2}}\right)\leq\widetilde{O}\left(\left\|\mathbf{u}\right\|_{2}\sqrt{dG_{T}}\right).$
The total runtime of the algorithm is $O(d^{2}T+d^{\omega}\sqrt{T}\log T)$.
Specifically, Theorem˜5 recovers Theorem 1 of Orabona et al. (2012), and Theorem˜6 recovers Theorem 8 of Cutkosky and Orabona (2018). The computational gains from replacing ONS with LightONS, discussed earlier, apply here.
Comparison with OQNS.
We remark that OQNS can hardly achieve full gradient-norm adaptivity, as the log-barrier regularization introduces an unavoidable $O(\log T)$ bias term independent of the gradient norm. From the perspective of OMD, the regret can be decomposed into a stability term and a bias term. Adopted from (Hazan, 2016), the regret decomposition of a typical OMD algorithm is as follows:
$\textsc{Reg}_{T}(\mathbf{u})\leq\underbrace{\frac{\eta}{2}\sum_{t=1}^{T}\left\|\nabla f_{t}(\mathbf{x}_{t})\right\|_{\ast}^{2}}_{\text{stability}}+\underbrace{\frac{\left\|\mathbf{u}-\mathbf{x}_{1}\right\|^{2}}{2\eta}}_{\text{bias}}.$ (24)
While ONS and LightONS, as instances of OMD, can flexibly balance stability and bias, OQNS sacrifices such flexibility for computational efficiency. OQNS employs a highly stable log-barrier, which suppresses the stability term but inflates the bias in a manner resistant to small gradient norms. Specifically, the log-barrier contributes a term like $\log(D^{2}-\left\|\mathbf{u}\right\|_{2}^{2})$ in the regret bound for any comparator $\mathbf{u}\in\mathcal{B}(D)$. As $\left\|\mathbf{u}\right\|_{2}$ approaches $D$, this term diverges. This issue is mitigated using the fixed-share trick (Mhammedi and Gatmiry, 2023), an illustration of which is as follows:
$\sum_{t=1}^{T}\left(f_{t}(\mathbf{x}_{t})-f_{t}(\mathbf{u})\right)\leq\underbrace{\sum_{t=1}^{T}\left(f_{t}(\mathbf{x}_{t})-f_{t}(\mathbf{v})\right)}_{\text{fixed-share regret}}+\underbrace{\sum_{t=1}^{T}\left(f_{t}(\mathbf{v})-f_{t}(\mathbf{u})\right)}_{\text{fixed-share margin}},\quad\text{where}~\mathbf{v}=\left(1-\frac{1}{T}\right)\mathbf{u}.$
In the fixed-share margin term, Lipschitzness yields $f_{t}(\mathbf{v})-f_{t}(\mathbf{u})\leq O(GD/T)$, summing to $O(1)$ over $T$ rounds. In the fixed-share regret term, since $1-\left\|\mathbf{v}\right\|_{2}^{2}\geq\Omega(D^{2}/T)$, the term $\log(D^{2}-\left\|\mathbf{v}\right\|_{2}^{2})$ contributes $O(\log T)$ to the regret, which is independent of $G_{T}$ and prevents full gradient-norm adaptivity.
5.2 Logistic Bandits
Generalized linear bandits (GLB) constitutes an important class of parametric stochastic bandits, where the expected loss depends on an unknown parameter through a known link function. This formulation introduces non-linearities into linear bandits, enhancing their expressivity while retaining tractable structures. We focus on logistic bandits as a representative instance of GLB to concretely illustrate the applicability of ONS and the integration of LightONS.
Problem setting.
Logistic bandits can be interpreted as interactions between a learner and an adversary, which unfolds as follows: At each round $t\in[T]$, the learner selects an arm $\mathbf{x}_{t}$ from an arm set $\mathcal{X}_{t}\subseteq\mathcal{B}(1)^{K}$ with $K$ arms, and suffers a stochastic loss $y_{t}\in\left\{0,1\right\}$, where $\mathop{\mathbb{P}}\left(y_{t}=1\,|\,\mathbf{x}_{t}\right)=\sigma(\mathbf{x}_{t}^{\mathsf{T}}\mathbf{w}^{\star})$. The true parameter $\mathbf{w}^{\star}\in\mathcal{W}=\mathcal{B}(D/2)$ is unknown, and $\sigma(z)=1/(1+\exp(-z))$ denotes the sigmoid function. The performance of the learner is measured by its pseudo-regret, quantifying the cumulative expected loss against the optimal arms in hindsight, which is defined as
$\textsc{Reg}_{T}=\sum_{t=1}^{T}\left(\sigma(\mathbf{x}_{t}^{\mathsf{T}}\mathbf{w}^{\star})-\sigma({\mathbf{x}^{\star}_{t}}^{\mathsf{T}}\mathbf{w}^{\star})\right),$
where $\mathbf{x}^{\star}_{t}=\operatorname*{arg\,max}_{\mathbf{x}\in\mathcal{X}_{t}}\mathbf{x}^{\mathsf{T}}\mathbf{w}^{\star}$.
A challenge in GLB lies in its dependence on $\kappa=\max_{\mathbf{x}\in\mathcal{X},\mathbf{w}\in\mathcal{W}}1/\sigma^{\prime}(\mathbf{x}^{\mathsf{T}}\mathbf{w})$, a problem-dependent constant that may grow expoenentially with $D$. Directly applying ONS to GLB yields a pseudo-regret bound of $O(\kappa\sqrt{T})$, which becomes vacuous for $\kappa=\Omega(\sqrt{T})$ (Jun et al., 2017).
Current results.
Recently, Zhang et al. (2025) propose the first GLB algorithm that achieves $\kappa$-free-leading-term pseudo-regret with constant working memory and constant per-round time. Their pseudo-regret upper bound is $O(d\sqrt{T}\log T+\kappa(d\log T)^{2})$, which remains sublinear in $T$ for $\kappa=o(T)$.
Their key idea is to estimate the parameter $\mathbf{w}^{\star}$ with a “look-ahead” variant of ONS, whose analytical properties can remove the dependence on $\kappa$ in the pseudo-regret’s leading terms while retaining the constant-memory and constant-time efficiency of ONS. An illustration of their “look-ahead” ONS is provided as follows, where $\ell_{t}(\mathbf{w})=-y_{t}\log\sigma(\mathbf{x}_{t}^{\mathsf{T}}\mathbf{w})-(1-y_{t})\log(1-\sigma(\mathbf{x}_{t}^{\mathsf{T}}\mathbf{w}))$ denotes the logistic loss for parameter estimation at round $t$:
$\widetilde{H}_{t}=\epsilon I+\sum_{i=1}^{t-1}\nabla^{2}\ell_{i}(\mathbf{w}_{i+1})+\eta\nabla^{2}\ell_{t}(\mathbf{w}_{t}),~\widehat{\mathbf{w}}_{t+1}=\mathbf{w}_{t}-\frac{1}{\eta}\widetilde{H}_{t}^{-1}\nabla\ell_{t}(\mathbf{w}_{t}),~\mathbf{w}_{t+1}=\Pi^{\widetilde{H}_{t}}_{\mathcal{B}(D/2)}[\widehat{\mathbf{w}}_{t+1}].$ (25)
As a result, their algorithm incurs the worst-case $\widetilde{O}((d^{2}K+d^{\omega})T)$ runtime bottleneck of ONS, where $K$ is the number of arms. Below we show that plugging LightONS can improve the runtime while preserving the $\kappa$-free-leading-term pseudo-regret guarantee.
Improvements by LightONS.
The LightONS counterpart of the “look-ahead” ONS replaces the Mahalanobis projection in Eq.˜25 with the projection-hysteresis mechanism as follows:
$\mathbf{w}_{t+1}=\begin{cases}\widehat{\mathbf{w}}_{t+1}&\text{if}~\left\|\widehat{\mathbf{w}}_{t+1}\right\|_{2}\leq kD/2\\ \Pi^{\widetilde{H}_{t}}_{\mathcal{B}(D/2)}[\widehat{\mathbf{w}}_{t+1}]&\text{otherwise}\end{cases}.$ (26)
The replacement entails theoretical guarantees as follows, with proofs deferred to Section˜D.2.
Theorem 7 (LightONS’s improvement for (Zhang et al., 2025)).
If $\bigcup_{t=1}^{T}\mathcal{X}_{t}\subseteq\mathcal{B}(1)$, $\left\|\mathbf{w}^{\star}\right\|\leq D/2$ and $D$ is known to the learner, then there exists an algorithm for binary logistic bandits, whose pseudo-regret is $O(d\sqrt{T}\log T+\kappa(d\log T)^{2})$, and whose total runtime is $\widetilde{O}(d^{2}KT+d^{\omega}\cdot\min\{\sqrt{\kappa dT},T\})$.
The algorithm in Theorem˜7 reduces the total runtime from $\widetilde{O}((d^{2}K+d^{\omega})T)$ to $\widetilde{O}(d^{2}KT+d^{\omega}\cdot\min\{\sqrt{\kappa dT},T\})$, while preserving the $\kappa$-free-leading-term pseudo-regret guarantee. When $\kappa=o(T)$, as required for sublinear pseudo-regret, the improved algorithm is asymptotically faster than the original algorithm.
We remark that that Eq.˜26 does not suffer from improper learning concerns, because the algorithm’s decision is the arm $\mathbf{x}_{t}$, rather than the estimated parameter $\mathbf{w}_{t}$. Consequently, the replacement is free from the improper-to-proper conversion and resembles LightONS.Core.
Comparison with OQNS.
OQNS can hardly be integrated into $\kappa$-free-leading-term GLB algorithms, as it is tailored to the OXO protocol and lacks the flexibility to handle customized local norms beyond those of ONS. Specifically, the local norm $\widetilde{H}_{t}$ in (Zhang et al., 2025), i.e., Eq.˜25, is not a simple accumulation of gradient outer products, unlike the OQNS’s local norm in Eq.˜7b. In contrast, LightONS retains the structural flexibility of ONS, enabling integration into GLB with minimal modifications.
5.3 Memory-Efficient OXO
In addition to the projection issue, another challenge of ONS is its $O(d^{2})$ working memory, in contrast to OGD’s $O(d)$ working memory. Luo et al. (2016) propose Sketched Online Newton Step (SON), which mitigates this issue by incorporating matrix sketching into ONS.
Current results.
SON achieves runtime and working memory linear in $d$ when the sketched dimension $d^{\prime}\ll d$. The value of $d^{\prime}$ typically depends on the problem’s intrinsic dimensionality, such as the number of non-zero eigenvalues in the Hessian-related matrix. The theoretical guarantees of SON are summarized in the following proposition.
Proposition 3 (Theorem 3 of Luo et al. (2016)).
Under Assumption˜2 and additional assumptions described in Eqs.˜30 and 31, SON (Algorithms 1 and 6 of Luo et al. (2016)) satisfies that,
$\textsc{Reg}_{T}(\mathbf{u})\leq\frac{d^{\prime}}{\gamma}\log\left(1+\frac{G^{2}}{2d^{\prime}\epsilon}T\right)+\frac{\gamma\epsilon D^{2}}{8}+\frac{\Delta_{1:T}}{2\gamma},$ (27)
where $\mathbf{u}$ and $\gamma$ are restricted by Eq.˜31, $\epsilon$ is the preconditioner coefficient similar to that in ONS and LightONS, and $\Delta_{1:T}$ is the cumulative sketching error, which can be bounded as
$\Delta_{1:T}\leq\min_{j\in[d^{\prime}]}\frac{2d^{\prime}}{(d^{\prime}-j+1)\epsilon}\sum_{i=j}^{d}\lambda_{i}\left(\sum_{t=1}^{T}\nabla f_{t}(\mathbf{x}_{t})\nabla f_{t}(\mathbf{x}_{t})^{\mathsf{T}}\right),$ (28)
and $\lambda_{i}$ is the $i$-th greatest eigenvalue. The working memory is $O(d^{\prime}d)$ and the total runtime is  [^3] The $\log T$ term in Eq. 29 arises from the eigendecomposition underlying sketching. Although Theorem 3 of Luo et al. (2016) ignores it, we include it for comparison with LightONS and OQNS.
$O\left(d^{\prime}dT\log T\right).$ (29)
However, SON demands additional assumptions. Its domain $\mathcal{X}_{t}$ must be an intersection of two parallel half-spaces, onto which Mahalanobis projections admit closed-form with $O(d^{2})$ time. For bounded convex domains in OXO, SON loses its computational advantage and reverts to the high computational cost of ONS. The domain restriction further imposes a stronger assumption on loss functions. Under Assumptions˜1, 2 and 3, the curvature parameter $\gamma_{0}=\frac{1}{2}\min\{\frac{1}{DG},\alpha\}$ in Lemma˜1 collapses to zero for the unbounded intersection-of-parallel-half-spaces domain, reducing exp-concavity to convexity. Consequently, SON assumes an explicit quadratic property analogous to Lemma˜1, rather than the standard Assumption˜3. We list these additional assumptions as follows.
  • Additional assumption on domains. For any $t\in[T]$, the domain at the $t$-th round is the intersection of two parallel half-spaces, i.e.,
    $\mathcal{X}_{t}=\{\mathbf{x}\,|\,|\mathbf{w}_{t}^{\mathsf{T}}\mathbf{x}|<D/2\},\quad\text{where}~\mathbf{w}_{t}\in\mathbb{R}^{d}~\text{is known and}~\left\|\mathbf{w}_{t}\right\|=1.$ (30)
  • Additional assumption on loss functions. For any $t\in[T]$, the loss function $f_{t}:\mathcal{X}\to\mathbb{R}$, the trajectory $\{\mathbf{x}_{t}\}_{t=1}^{T}$ and the comparator $\mathbf{u}\in\bigcap_{t=1}^{T}\mathcal{X}_{t}$ satisfies that
    $f_{t}(\mathbf{x})-f_{t}(\mathbf{u})\leq\nabla f_{t}(\mathbf{x})^{\mathsf{T}}(\mathbf{x}-\mathbf{u})-\frac{\gamma}{2}\left(\nabla f_{t}(\mathbf{x})^{\mathsf{T}}(\mathbf{x}-\mathbf{u})\right)^{2}.$ (31)
Improvements by LightONS.
Replacing all accesses to the Hessian-related matrix $A_{t}$ with the sketching primitives as in SON yields a hybrid method that combines the projection efficiency of LightONS with the memory efficiency of SON. This hybrid method extends linear-in-$d$ runtime and working memory to the standard OXO setting in Section˜2.1, while retaining the regret $O(d^{\prime}\log T)$ of SON. Theoretical guarantees of the hybrid method are summarized in the following theorem, with algorithms and proofs deferred to Section˜D.3.
Theorem 8 (LightONS’s improvement for memory-efficient OXO).
Under Assumptions˜1, 2 and 3, there exists an algorithm, which satisfies that, for any $\mathbf{u}\in\mathcal{X}$,
$\textsc{Reg}_{T}(\mathbf{u})\leq\frac{d^{\prime}}{\gamma_{0}}\log\left(1+\frac{G^{2}}{2d^{\prime}\epsilon}T\right)+\frac{\gamma_{0}\epsilon D^{2}}{8}+\frac{\Delta_{1:T}}{2\gamma_{0}},$ (32)
where $\gamma_{0}=\frac{1}{2}\min\left\{\frac{1}{DG},\alpha\right\}$, $\epsilon$ is the preconditioner coefficient similar to that in LightONS, $\Delta_{1:T}$ is the cumulative sketching error, which can be bounded as
$\Delta_{1:T}\leq\min_{j\in[d^{\prime}]}\frac{2d^{\prime}}{(d^{\prime}-j+1)\epsilon}\sum_{i=j}^{d}\lambda_{i}\left(\sum_{t=1}^{T}\nabla g_{t}(\mathbf{y}_{t})\nabla g_{t}(\mathbf{y}_{t})^{\mathsf{T}}\right),$ (33)
$\nabla g_{t}(\mathbf{y}_{t})$ is the surrogate gradient as in Lemma˜3, and $\lambda_{i}$ is the $i$-th greatest eigenvalue. The working memory is $O(d^{\prime}d)$ and the total runtime is
$O\left(\left(\texttt{EP}_{\mathcal{X}}+d^{\prime}d\log T\right)T+d^{\omega}\sqrt{(d+\Delta_{1:T})T/\epsilon}\log T\right).$ (34)
Comparison with OQNS.
Mhammedi and Gatmiry (2023) mention the possibility of combining OQNS with sketching but provide neither algorithms nor analysis. Incorporating sketching further complicates OQNS’s already intricate analysis, as the sketching error interacts with the log-barrier, Hessian approximation, and decisions’ update, as shown in Eqs.˜7b and 7a. In contrast, within the OMD framework, sketching errors naturally appear as an additive term in the regret bound via OMD’s stability term, illustrated in Eq.˜24, since OMD’s regret decomposition is somewhat orthogonal to sketching.
6 Conclusion
For online exp-concave optimization (OXO), we propose LightONS, a simple yet powerful variant of ONS. By combining the projection-hysteresis mechanism with the improper-to-proper conversion, LightONS achieves significant reductions in worst-case total runtime while preserving the regret optimality of ONS. These gains extend to the stochastic optimization setting, answering a COLT’13 open problem on efficient and optimal SXO (Koren, 2013). Moreover, due to its fidelity to the mirror-descent udpate of ONS, LightONS serves as a drop-in replacement across diverse applications, including regret bounds adaptive to gradient norms, parametric stochastic bandits, and memory-efficient settings, all without compromising statistical guarantees.
Several important directions remain open. First, while LightONS adapts to gradient norms, achieving adaptivity to gradient variation for OXO within $\widetilde{O}(d^{2}T)$ time remains an open challenge (Chiang et al., 2012; Zhao et al., 2024). Second, applying the LightONS technique to other preconditioned online learning algorithms may not improve their (asymptotic) runtime if the bottleneck lies in other operations, such as explicit matrix factorization rather than Mahalanobis projections (e.g., in AdaGrad (Duchi et al., 2011)). Accelerating these algorithms remains unclear. These future directions motivate the development of broader, general-purpose acceleration techniques for Hessian-related online learning algorithms.
Appendix A Technical Lemmas
In this section, we present technical lemmas for the analysis of the proposed OXO algorithms.
A.1 Elliptical Potential Lemmas
We begin with two lemmas addressing Hessian-related matrices, which are essential for both regret and runtime analysis. Lemma˜5 is standard in the ONS literature (e.g., [Hazan et al., 2007, Luo et al., 2016]). The proof of Lemma˜6 mirrors that of Lemma˜5, differing only in the potential function: $F(A)=-\operatorname{tr}(A)$ versus $F(A)=\log\operatorname{det}(A)$.
Lemma 5 (elliptical potential lemma).
Let $A_{i}=\lambda I+\sum_{j=1}^{i}\mathbf{v}_{j}\mathbf{v}_{j}^{\mathsf{T}}$, if $\left\|\mathbf{v}_{i}\right\|_{2}\leq L$ for any $i\in[n]$, then
$\sum_{i=1}^{n}\mathbf{v}_{i}^{\mathsf{T}}A_{i}^{-1}\mathbf{v}_{i}\leq\log\operatorname{det}(A_{n})-\log\operatorname{det}(A_{0})\leq d\log\left(1+\frac{L^{2}}{d\lambda}n\right).$ (35)
Lemma 6.
Let $A_{i}=\lambda I+\sum_{j=1}^{i}\mathbf{v}_{j}\mathbf{v}_{j}^{\mathsf{T}}$, if $\left\|\mathbf{v}_{i}\right\|_{2}\leq L$ for any $i\in[n]$, then
$\sum_{i=1}^{n}\mathbf{v}_{i}^{\mathsf{T}}A_{i}^{-2}\mathbf{v}_{i}\leq\operatorname{tr}(A_{0}^{-1})-\operatorname{tr}(A_{n}^{-1})\leq\frac{d}{\lambda}.$ (36)
Proof of Lemmas˜5 and 6.
Let $\mathbb{S}_{++}$ denote the set of $d$-dimensional positive-definite and symmetric matrices, and let $\left\langle\cdot,\cdot\right\rangle_{\operatorname{F}}$ denote the inner product induced by the Frobenius matrix norm. Then for any concave function $F:\mathbb{S}_{++}\to\mathbb{R}$, we have
$\sum_{i=1}^{n}\mathbf{v}_{i}^{\mathsf{T}}\nabla{F}(A_{i})\mathbf{v}_{i}=\sum_{i=1}^{n}\left\langle\nabla{F}(A_{i}),A_{i}-A_{i-1}\right\rangle_{\operatorname{F}}\leq\sum_{i=1}^{n}\left(F(A_{i})-F(A_{i-1})\right)=F(A_{n})-F(A_{0}).$ (37)
The first equality follows from $A_{i}-A_{i-1}=\mathbf{v}_{i}\mathbf{v}_{i}^{\mathsf{T}}$ (by the definition of $A_{i}$) and $\mathbf{u}^{\mathsf{T}}A\mathbf{v}=\left\langle A,\mathbf{v}\mathbf{u}^{\mathsf{T}}\right\rangle_{\operatorname{F}}$ (by the definition of the $\left\langle\cdot,\cdot\right\rangle_{\operatorname{F}}$); The inequality from the concavity of $F$; The final equality from telescoping. Substituting the respective potentials into Eq.˜37 yields Lemmas˜5 and 6. Concretely,
  • For Lemma˜5, $F(A)=\log\operatorname{det}(A)$ and $\nabla{F}(A)=A^{-1}$.
  • For Lemma˜6, $F(A)=-\operatorname{tr}(A^{-1})$ and $\nabla{F}(A)=A^{-2}$.
Concavity of $F$ is standard [Boyd and Vandenberghe, 2004] and gradients of $F$ follow from matrix calculus [Petersen and Petersen, 2012]. ∎
A.2 Proof of Lemma˜1
Proof.
Since $g(\mathbf{x})=\mathrm{e}^{-\alpha f(\mathbf{x})}$ is concave and positive, then $h(\mathbf{x})=\mathrm{e}^{-2\gamma_{0}f(\mathbf{x})}$ is also concave because $2\gamma_{0}/\alpha\leq 1$ and, for any $\lambda\in[0,1]$ and any $(\mathbf{x},\mathbf{y})\in\mathcal{X}^{2}$,
$h(\lambda\mathbf{x}+(1-\lambda)\mathbf{y})$ $=g(\lambda\mathbf{x}+(1-\lambda)\mathbf{y})^{2\gamma_{0}/\alpha}\geq\left(\lambda g(\mathbf{x})+(1-\lambda)g(\mathbf{y})\right)^{2\gamma_{0}/\alpha}$
$\geq\lambda g(\mathbf{x})^{2\gamma_{0}/\alpha}+(1-\lambda)g(\mathbf{y})^{2\gamma_{0}/\alpha}=\lambda h(\mathbf{x})+(1-\lambda)h(\mathbf{y})$
where the first inequality follows from the concavity of $g$ and the second from the concavity of $a\mapsto a^{2\gamma_{0}/\alpha}$. Then, by the concavity of $h$, for any $(\mathbf{x},\mathbf{u})\in\mathcal{X}^{2}$,
$\mathrm{e}^{-2\gamma_{0}f(\mathbf{u})}\leq\mathrm{e}^{-2\gamma_{0}f(\mathbf{x})}+\left(-2\gamma_{0}\mathrm{e}^{-2\gamma_{0}f(\mathbf{x})}\nabla f(\mathbf{x})\right)^{\mathsf{T}}(\mathbf{u}-\mathbf{x}).$
Rearranging the preceding inequality yields
$f(\mathbf{x})-f(\mathbf{u})\leq\frac{1}{2\gamma_{0}}\log\left(1-2\gamma_{0}\nabla f(\mathbf{x})^{\mathsf{T}}(\mathbf{u}-\mathbf{x})\right).$
To prove Eq.˜3, it suffices to apply that $\log(1+a)\leq a-a^{2}/4$ (which holds for any $\left|a\right|\leq 1$) to the right-hand side of the preceding inequality. We note that $\left|2\gamma_{0}\nabla f(\mathbf{x})^{\mathsf{T}}(\mathbf{x}-\mathbf{u})\right|\leq 1$ due to the selection of $\gamma_{0}=\frac{1}{2}\min\left\{\frac{1}{DG},\alpha\right\}$. ∎
Appendix B Proofs for Section˜3
In this section, we prove the guarantees of LightONS.
B.1 Proof of Lemma˜2
Proof.
Let $\Phi_{T}$ denote the sum of squared norms of updates, as in the left-hand side of Eq.˜36. Note that $A_{t}=\epsilon I+\sum_{i=1}^{t}\nabla f_{t}(\mathbf{x}_{t})\nabla f_{t}(\mathbf{x}_{t})^{\mathsf{T}}$ in Algorithm˜1 and $\left\|\nabla f_{t}(\mathbf{x}_{t})\right\|_{2}\leq G$ in Assumption˜2. Then by Lemma˜6 we have
$\Phi_{T}\triangleq\sum_{t=1}^{T}\left\|\frac{1}{\gamma}A_{t}^{-1}\nabla f_{t}(\mathbf{x}_{t})\right\|_{2}^{2}\leq\frac{1}{\gamma^{2}}\frac{d}{\epsilon}.$
Let $\left\{\tau_{i}\right\}_{i=1}^{N}\subseteq[T]$ denote all the rounds when LightONS.Core performs the Mahalanobis projection. With $\tau_{0}=1$, we have that for any $i\in[N]$,
$\sum_{t=\tau_{i-1}}^{\tau_{i}-1}\frac{1}{\gamma}A_{t}^{-1}\nabla f_{t}(\mathbf{x}_{t})=\widehat{\mathbf{x}}_{\tau_{i}}-\mathbf{x}_{\tau_{i-1}}.$
Since $\left\|\widehat{\mathbf{x}}_{\tau_{i}+1}\right\|_{2}>kD/2$ and $\left\|\mathbf{x}_{\tau_{i}+1}\right\|_{2}=\left\|\Pi^{A_{\tau_{i}}}_{\mathcal{X}}[\widehat{\mathbf{x}}_{\tau_{i}+1}]\right\|_{2}\leq D/2$, we have
$\left\|\sum_{t=\tau_{i-1}}^{\tau_{i}-1}\frac{1}{\gamma}A_{t}^{-1}\nabla f_{t}(\mathbf{x}_{t})\right\|_{2}=\left\|\widehat{\mathbf{x}}_{\tau_{i}}-\mathbf{x}_{\tau_{i-1}}\right\|_{2}>(k-1)\frac{D}{2}.$
By applying the inequality $n\sum_{i=1}^{n}\left\|\mathbf{v}_{i}\right\|_{2}^{2}\geq\left\|\sum_{i=1}^{n}\mathbf{v}_{i}\right\|_{2}^{2}$, we obtain
$(\tau_{i}-\tau_{i-1})\frac{1}{\gamma^{2}}\sum_{t=\tau_{i-1}}^{\tau_{i}-1}\left\|A_{t}^{-1}\nabla f_{t}(\mathbf{x}_{t})\right\|_{2}^{2}\geq\left\|\sum_{t=\tau_{i-1}}^{\tau_{i}-1}\frac{1}{\gamma}A_{t}^{-1}\nabla f_{t}(\mathbf{x}_{t})\right\|_{2}^{2}>(k-1)^{2}\frac{D^{2}}{4}.$
Rearranging the preceding inequality yields
$\frac{1}{\tau_{i}-\tau_{i-1}}<\frac{4}{(k-1)^{2}D^{2}\gamma^{2}}\sum_{t=\tau_{i-1}}^{\tau_{i}-1}\left\|A_{t}^{-1}\nabla f_{t}(\mathbf{x}_{t})\right\|_{2}^{2}.$
By applying the inequality $\frac{n}{\sum_{i=1}^{n}\frac{1}{a_{i}}}\leq\frac{\sum_{i=1}^{n}a_{i}}{n}$ and the fact that $\tau_{N}\leq T$, we obtain
$\frac{N^{2}}{T}\leq\frac{N^{2}}{\tau_{N}-1}\leq\sum_{i=1}^{N}\frac{1}{\tau_{i}-\tau_{i-1}}<\frac{4}{(k-1)^{2}D^{2}\gamma^{2}}\sum_{t=1}^{\tau_{N}}\left\|A_{t}^{-1}\nabla f_{t}(\mathbf{x}_{t})\right\|_{2}^{2}\leq\frac{4}{(k-1)^{2}D^{2}\gamma^{2}}\Phi_{T}.$
Finally, rearranging the preceding inequality yields the desired result $N\leq\frac{2}{D\gamma}\sqrt{\frac{d}{\epsilon}T}$. ∎
We remark that Lemma˜2 is tight in terms of $T$ up to logarithmic factors. Consider $d=1$ and $\nabla f_{t}(\mathbf{x}_{t})=1/\sqrt{T}$, then $A_{t}^{-1}\nabla f_{t}(\mathbf{x}_{t})=\Omega(1/\sqrt{T})$ and $\|\sum_{t=1}^{T}A_{t}^{-1}\nabla f_{t}(\mathbf{x}_{t})\|_{2}=\Omega(\sqrt{T})$.
B.2 Proof of Theorem˜1
Proof.
The proof of Theorem˜1 follows as a specialization of the proof of Theorem˜2 given in Section˜B.4. The only substantive difference is that, in the present setting, the analysis is carried out directly on the original loss functions $f_{t}$, rather than on surrogate losses $g_{t}$.
We therefore omit the repetitive details and highlight only the key intermediate arguments.
  • Choice of the expanded curvature parameter $\gamma$ in Algorithm˜2. The expression $\gamma=\frac{1}{2}\min\{\frac{2}{(k+1)DG},\alpha\}$ directly follows from Lemma˜1.
  • OMD regret analysis with selective projections. The regret analysis follows the same structure as Lemma˜8, which accounts for projections triggered only when the iterate exits the expanded domain $\widetilde{\mathcal{X}}$. Details are given in Section˜B.4.
  • Runtime of LightONS.Core. The runtime bound is derived in the same manner as in the proof of Theorem˜2 in Section˜B.4.
B.3 Proof of Lemma˜4
Proof.
By the definition of curvature parameters, namely $\gamma_{0}=\frac{1}{2}\min\left\{\frac{1}{DG},\alpha\right\}$ in Lemma˜1 and $\gamma^{\prime}=\frac{1}{2}\min\left\{\frac{1}{c_{f}c_{g}DG},\frac{4}{(k+1)c_{f}c_{g}DG},\alpha\right\}$ in Algorithm˜3, we first have $\gamma^{\prime}\leq\gamma_{0}$. Then by relaxing Eq.˜3 in Lemma˜1 we have
$f_{t}(\mathbf{x}_{t})-f_{t}(\mathbf{u})$ $\leq\nabla f_{t}(\mathbf{x}_{t})^{\mathsf{T}}(\mathbf{x}_{t}-\mathbf{u})-\frac{\gamma_{0}}{2}\left(\nabla f_{t}(\mathbf{x}_{t})^{\mathsf{T}}(\mathbf{x}_{t}-\mathbf{u})\right)^{2}$
$\leq\nabla f_{t}(\mathbf{x}_{t})^{\mathsf{T}}(\mathbf{x}_{t}-\mathbf{u})-\frac{\gamma^{\prime}}{2}\left(\nabla f_{t}(\mathbf{x}_{t})^{\mathsf{T}}(\mathbf{x}_{t}-\mathbf{u})\right)^{2}$
Let $U(a)=a-\frac{\gamma^{\prime}}{2}a^{2}$. To finish the proof, it suffices to show that
$U\left(\nabla f_{t}(\mathbf{x}_{t})^{\mathsf{T}}(\mathbf{x}_{t}-\mathbf{u})\right)\leq U\left(\nabla g_{t}(\mathbf{y}_{t})^{\mathsf{T}}(\mathbf{y}_{t}-\mathbf{u})\right).$
Note that $U$ is monotonically increasing on $(-\infty,\frac{1}{\gamma^{\prime}}]$. Therefore, it suffices to show that
$\nabla f_{t}(\mathbf{x}_{t})^{\mathsf{T}}(\mathbf{x}_{t}-\mathbf{u})\leq c_{f}\nabla g_{t}(\mathbf{y}_{t})^{\mathsf{T}}(\mathbf{y}_{t}-\mathbf{u})\leq c_{f}G\cdot\max\left\{\frac{k+1}{2}c_{g}D,2c_{g}D\right\}\leq\frac{1}{\gamma^{\prime}}.$
The first inequality follows from Condition˜1; The second inequality from the Cauchy-Schwarz inequality $\mathbf{u}^{\mathsf{T}}\mathbf{v}\leq\left\|\mathbf{u}\right\|_{2}\left\|\mathbf{v}\right\|_{2}$; The last inequality from the definition of $\gamma^{\prime}$. ∎
B.4 Proof of Theorem˜2
We first introduce a property of the Mahalanobis projection. This lemma is often referred to as the Pythagorean theorem in Banach space or the non-expansiveness of projections.
Lemma 7 (Lemma 8 of Hazan et al. [2007]).
If $A\in\mathbb{R}^{d\times d}$ is positive-definite and symmetric matrix, then for any compact convex set $\mathcal{X}\subseteq\mathbb{R}^{d}$, any point $\mathbf{y}\in\mathbb{R}^{d}$ and any point $\mathbf{u}\in\mathcal{X}$, $\left\|\Pi^{A}_{\mathcal{X}}[\mathbf{y}]-\mathbf{u}\right\|_{A}\leq\left\|\mathbf{y}-\mathbf{u}\right\|_{A}$.
Based on Lemma˜7, we can decompose the regret of LightONS as the following lemma shows.
Lemma 8.
Ignoring the truncation error of Mahalanobis projections, under the same assumptions as Theorem˜2, in Algorithm˜3, for any $t\in[T]$ and all $\mathbf{u}\in\mathcal{X}$, it holds that
$2\nabla g_{t}(\mathbf{y}_{t})^{\mathsf{T}}(\mathbf{y}_{t}-\mathbf{u})\leq\frac{1}{\gamma^{\prime}}\left\|\nabla g_{t}(\mathbf{y}_{t})\right\|_{A_{t}^{-1}}^{2}+\gamma^{\prime}\left\|\mathbf{y}_{t}-\mathbf{u}\right\|_{A_{t}}^{2}-\gamma^{\prime}\left\|\mathbf{y}_{t+1}-\mathbf{u}\right\|_{A_{t}}^{2}.$ (38)
Proof of Lemma˜8.
When $\left\|\widehat{\mathbf{y}}_{t+1}\right\|_{2}\leq kD/2$, no Mahalanobis projection is performed, i.e., $\mathbf{y}_{t+1}=\widehat{\mathbf{y}}_{t+1}=\mathbf{y}_{t}-\frac{1}{\gamma^{\prime}}A_{t}^{-1}\nabla g_{t}(\mathbf{y}_{t})$, then for any $\mathbf{u}\in\mathbb{R}^{d}$, we have
$\left\|\mathbf{y}_{t+1}-\mathbf{u}\right\|_{A_{t}}^{2}=\left\|\mathbf{y}_{t}-\frac{1}{\gamma^{\prime}}A_{t}^{-1}\nabla g_{t}(\mathbf{y}_{t})-\mathbf{u}\right\|_{A_{t}}^{2}.$
Otherwise, when $\left\|\widehat{\mathbf{y}}_{t+1}\right\|_{2}>kD/2$, a Mahalanobis projection is performed, by Lemma˜7, for any $\mathbf{u}\in\mathcal{X}$, we have
$\left\|\mathbf{y}_{t+1}-\mathbf{u}\right\|_{A_{t}}^{2}=\left\|\Pi_{\mathcal{B}(D/2)}^{A_{t}}\left[\mathbf{y}_{t}-\frac{1}{\gamma^{\prime}}A_{t}^{-1}\nabla g_{t}(\mathbf{y}_{t})\right]-\mathbf{u}\right\|_{A_{t}}^{2}\leq\left\|\mathbf{y}_{t}-\frac{1}{\gamma^{\prime}}A_{t}^{-1}\nabla g_{t}(\mathbf{y}_{t})-\mathbf{u}\right\|_{A_{t}}^{2}.$
Combining both cases, we conclude that, for any $t\in[T]$ and any $\mathbf{u}\in\mathcal{X}$,
$\left\|\mathbf{y}_{t+1}-\mathbf{u}\right\|_{A_{t}}^{2}\leq\left\|\left(\mathbf{y}_{t}-\mathbf{u}\right)-\frac{1}{\gamma^{\prime}}A_{t}^{-1}\nabla g_{t}(\mathbf{y}_{t})\right\|_{A_{t}}^{2}.$ (39)
Rearranging Eq.˜39 yields the desired result of Eq.˜38. ∎
Note that Lemma˜8 ignores the truncation error of the Mahalanobis projections. The next lemma complements the analysis by showing that, thanks to FastProj (Algorithm˜4), the truncation error only incurs a negligible additive $O(1/t^{2})$ term in the regret decomposition, which can be safely ignored in the final regret bound of LightONS.
Lemma 9.
Let the Mahalanobis projection of Algorithm˜3 is implemented with Algorithm˜4 with $A=A_{t}$, $R=D/2$, $\mathbf{u}=\widehat{\mathbf{y}}_{t+1}$, $\zeta=\zeta_{t}$, $\underline{\lambda}=\epsilon$, $\overline{\lambda}=\epsilon+G^{2}t$, where $\zeta_{t}$ is defined as
$\zeta_{t}=\min\left\{\frac{\gamma^{\prime}}{2kD(c_{g}^{2}G^{2}t+\epsilon)t^{2}},\frac{1}{t}\sqrt{\frac{\gamma^{\prime}}{2(c_{g}^{2}G^{2}t+\epsilon)t}}\right\}.$ (40)
Then under the same assumptions as Theorem˜2, in Algorithm˜3, for any $t\in[T]$ and all $\mathbf{u}\in\mathcal{X}$,
$2\nabla g_{t}(\mathbf{y}_{t})^{\mathsf{T}}(\mathbf{y}_{t}-\mathbf{u})\leq\frac{1}{\gamma^{\prime}}\left\|\nabla g_{t}(\mathbf{y}_{t})\right\|_{A_{t}^{-1}}^{2}+\gamma^{\prime}\left\|\mathbf{y}_{t}-\mathbf{u}\right\|_{A_{t}}^{2}-\gamma^{\prime}\left\|\mathbf{y}_{t+1}-\mathbf{u}\right\|_{A_{t}}^{2}+\frac{1}{t^{2}}.$ (41)
Proof of Lemma˜9.
It suffices to prove the case when $\left\|\widehat{\mathbf{y}}_{t+1}\right\|_{2}>kD/2$. Let $\mathbf{y}^{\ast}_{t+1}=\Pi_{\mathcal{B}(D/2)}^{A_{t}}\left[\widehat{\mathbf{y}}_{t+1}\right]$ denote the exact Mahalanobis projection without truncation error, and $\bm{\delta}_{t}=\mathbf{y}^{\ast}_{t+1}-\mathbf{y}_{t+1}$ denote the truncation error. By Lemma˜11, $\left\|\bm{\delta}_{t}\right\|_{2}=\left\|\mathbf{y}_{t+1}-\mathbf{y}^{\ast}_{t+1}\right\|_{2}\leq\zeta_{t}$; By Lemma˜7, for any $\mathbf{u}\in\mathcal{X}$,
$\left\|\mathbf{y}_{t+1}-\mathbf{u}\right\|_{A_{t}}^{2}$ $=\left\|\mathbf{y}^{\ast}_{t+1}-\mathbf{u}\right\|_{A_{t}}^{2}+2\left(\mathbf{y}^{\ast}_{t+1}-\mathbf{u}\right)^{\mathsf{T}}A_{t}\bm{\delta}_{t}+\left\|\bm{\delta}_{t}\right\|_{A_{t}}^{2}$
$\leq\left\|\mathbf{y}^{\ast}_{t+1}-\mathbf{u}\right\|_{A_{t}}^{2}+2\cdot 2D\cdot(c_{g}^{2}G^{2}T+\epsilon)\cdot\left\|\bm{\delta}_{t}\right\|_{2}+(c_{g}^{2}G^{2}T+\epsilon)\cdot\left\|\bm{\delta}_{t}\right\|_{2}^{2}$
$\leq\left\|\mathbf{y}_{t}-\frac{1}{\gamma^{\prime}}A_{t}^{-1}\nabla g_{t}(\mathbf{y}_{t})-\mathbf{u}\right\|_{A_{t}}^{2}+\frac{2\gamma^{\prime}}{t^{2}}$
The first inequality considers the operator norm of $A_{t}$ and uses the fact that $\left\|\nabla g_{t}(\mathbf{y}_{t})\right\|_{2}\leq c_{g}G$ in Condition˜1; The second inequality comes from the selection of $\zeta_{t}$ in Eq.˜40. ∎
For conciseness, we ignore the truncation error of the Mahalanobis projections in the proof of Theorem˜2, since it only incurs an additive $O(\sum_{t=1}^{T}1/t^{2})=O(1)$ term in the final regret bound.
Proof of Theorem˜2.
Substituting Eq.˜38 into Eq.˜13 yields the following inequalities:
[…]
$=$ $\frac{1}{2\gamma^{\prime}}\left\|\nabla g_{t}(\mathbf{y}_{t})\right\|_{A_{t}^{-1}}^{2}+\frac{\gamma^{\prime}}{2}\left\|\mathbf{y}_{t}-\mathbf{u}\right\|_{A_{t-1}}^{2}-\frac{\gamma^{\prime}}{2}\left\|\mathbf{y}_{t+1}-\mathbf{u}\right\|_{A_{t}}^{2}$
Telescoping the preceding inequality establishes the desired regret bound Eq.˜14:
$\sum_{t=1}^{T}\left(f_{t}(\mathbf{x}_{t})-f_{t}(\mathbf{u})\right)$ $\leq\sum_{t=1}^{T}\left(\frac{1}{2\gamma^{\prime}}\left\|\nabla g_{t}(\mathbf{y}_{t})\right\|_{A_{t}^{-1}}^{2}+\frac{\gamma^{\prime}}{2}\left\|\mathbf{y}_{t}-\mathbf{u}\right\|_{A_{t-1}}^{2}-\frac{\gamma^{\prime}}{2}\left\|\mathbf{y}_{t+1}-\mathbf{u}\right\|_{A_{t}}^{2}\right)$
$=\left(\frac{1}{2\gamma^{\prime}}\sum_{t=1}^{T}\left\|\nabla g_{t}(\mathbf{y}_{t})\right\|_{A_{t}^{-1}}^{2}\right)+\frac{\gamma^{\prime}}{2}\left\|\mathbf{y}_{1}-\mathbf{u}\right\|_{A_{0}}^{2}-\frac{\gamma^{\prime}}{2}\left\|\mathbf{y}_{T+1}-\mathbf{u}\right\|_{A_{T}}^{2}$
[…]
Finally, the desired runtime Eq.˜15 follows from the following two parts:
  • Runtime aside from FastProj. In each round, the domain conversion of Lemma˜3 takes $O(\texttt{EP}_{\mathcal{X}}+d)$ time, updating and inverting $A_{t}$ as Eq.˜18 takes $O(d^{2})$ time, and other operations, such as computing the surrogate gradient as Eq.˜12b, take only $O(d)$ time. Overall runtime aside from FastProj is $O((\texttt{EP}_{\mathcal{X}}+d^{2})T)$.
  • Runtime of FastProj. In each call of FastProj, by Theorem˜3, the iteration number of bisections $n_{t}=O(\log t)=O(\log T)$ and each bisection takes $O(d^{\omega})$ time with Choice 1. Since LightONS.Core is a subroutine in LightONS, the number of calls to FastProj is at most $O((k-1)^{-1}d^{0.5}\sqrt{T/\epsilon})$ by Lemma˜2. Overall runtime of FastProj is $O((k-1)^{-1}d^{0.5}\sqrt{T/\epsilon}\cdot d^{\omega}\log T)$.
B.5 Proof of Theorem˜3
First we reduce the Mahalanobis projection onto an Euclidean ball to a one-dimensional root-finding problem as the following lemma shows.
Lemma 10.
$\Pi^{A}_{\mathcal{B}(R)}[\mathbf{y}]=(A+\mu^{\ast}I)^{-1}A\mathbf{y}$, where $\mu^{\ast}$ is the only positive zero of the following function
$0=\rho(\mu)=\left\|\left(A+\mu I\right)^{-1}A\mathbf{y}\right\|_{2}^{2}-R^{2}=\sum_{i=1}^{d}\frac{v_{i}^{2}}{\left(1+\frac{\mu}{\lambda_{i}}\right)^{2}}-R^{2}.$ (42)
Here $v_{i}=\mathbf{e}_{i}^{\mathsf{T}}Q^{\mathsf{T}}\mathbf{y}$, where $\mathbf{e}_{i}$ is the $i$-th standard basis vector, and $A=Q\Lambda Q^{\mathsf{T}}$ is the eigendecomposition, $\Lambda=\operatorname{diag}\left(\lambda_{1},\dots,\lambda_{d}\right)$ with $\lambda_{1}\geq\dots\geq\lambda_{d}>0$.
Besides, the positive zero of $\rho$, denoted as $\mu^{\ast}$, satisfies that
$\left(\frac{\left\|\mathbf{y}\right\|_{2}}{R}-1\right)\lambda_{d}<\mu^{\ast}<\left(\frac{\left\|\mathbf{y}\right\|_{2}}{R}-1\right)\lambda_{1}.$
Proof of Lemma˜10.
The Mahalanobis projection onto an Euclidean ball is formulated as
$\min_{\mathbf{x}\in\mathbb{R}^{d}}$ $(\mathbf{x}-\mathbf{y})^{\mathsf{T}}A(\mathbf{x}-\mathbf{y})$
s.t. $\mathbf{x}^{\mathsf{T}}\mathbf{x}\leq R^{2}$
The Lagrangian of this quadratic program is $\mathcal{L}=(\mathbf{x}-\mathbf{y})^{\mathsf{T}}A(\mathbf{x}-\mathbf{y})+\mu(\mathbf{x}^{\mathsf{T}}\mathbf{x}-R^{2})$. According to the KKT conditions, $\nabla\mathcal{L}=2A(\mathbf{x}-\mathbf{y})+2\mu\mathbf{x}=\bm{0}$ with $\mu>0$. (Otherwise, $\mu=0$ implies $\mathbf{y}\in\mathcal{B}(R)$ and the projection is trivial.) Rearranging $\nabla\mathcal{L}=\bm{0}$ yields $\mathbf{x}=(A+\mu I)^{-1}A\mathbf{y}$, thus $\mathbf{x}^{\mathsf{T}}\mathbf{x}=R^{2}$ is equivalent to $\rho(\mu)=0$, which proves Eq.˜42.
Note that $\sum_{i=1}^{d}v_{i}^{2}=\left\|Q\mathbf{y}\right\|_{2}^{2}=\left\|\mathbf{y}\right\|_{2}^{2}$ due to the orthogonality of $Q$. Therefore,
$\frac{\left\|\mathbf{y}\right\|_{2}^{2}}{(1+\frac{\mu}{\lambda_{d}})^{2}}-R^{2}\leq\rho(\mu)=\sum_{i=1}^{d}\frac{v_{i}^{2}}{(1+\frac{\mu}{\lambda_{i}})^{2}}-R^{2}\leq\frac{\left\|\mathbf{y}\right\|_{2}^{2}}{(1+\frac{\mu}{\lambda_{1}})^{2}}-R^{2}.$
It is straightforward that $\rho((\frac{\left\|\mathbf{y}\right\|_{2}}{R}-1)\lambda_{d})\geq 0$ and $\rho((\frac{\left\|\mathbf{y}\right\|_{2}}{R}-1)\lambda_{1})\leq 0$. Then, since $\rho$ monotonically decreases on $0,\infty)$, we conclude that $\mu^{\ast}\in((\frac{\left\|\mathbf{y}\right\|_{2}}{R}-1)\lambda_{d},(\frac{\left\|\mathbf{y}\right\|_{2}}{R}-1)\lambda_{1})$ and $\mu^{\ast}$ is unique on $0,\infty)$. ∎
The next lemma reveals the relationship between the truncation error of the root-finding and the truncation error of the Mahalanobis projection.
Lemma 11.
Let $\mathbf{x}^{\ast}=(A+\mu^{\ast}I)^{-1}A\mathbf{y}$, where $\rho(\mu^{\ast})=0$ as in Lemma˜10. If $\widetilde{\mu}\geq 0$ and $\left|\widetilde{\mu}-\mu^{\ast}\right|<\frac{\lambda_{d}}{\left\|\mathbf{y}\right\|_{2}}\zeta$, then $\mathbf{x}\in\mathcal{B}(R)$ and $\left\|\mathbf{x}-\mathbf{x}^{\ast}\right\|_{2}\leq\zeta$, where $\mathbf{x}$ is constructed from $\widetilde{\mu}$ as
$\mathbf{x}=\Pi_{\mathcal{B}(R)}\left[\widetilde{\mathbf{x}}\right]=\frac{R}{\left\|\widetilde{\mathbf{x}}\right\|_{2}}\cdot\widetilde{\mathbf{x}},\quad\widetilde{\mathbf{x}}=\left(A+\widetilde{\mu}I\right)^{-1}A\mathbf{y}.$ (43)
Proof of Lemma˜11.
Eq.˜43 immediately implies $\mathbf{x}\in\mathcal{B}(R)$.
By Lemma˜7, we have $\left\|\mathbf{x}-\mathbf{x}^{\ast}\right\|_{2}\leq\left\|\widetilde{\mathbf{x}}-\mathbf{x}^{\ast}\right\|_{2}$. Then it suffices to prove $\left\|\widetilde{\mathbf{x}}-\mathbf{x}^{\ast}\right\|_{2}\leq\zeta$:
$\left\|\widetilde{\mathbf{x}}-\mathbf{x}^{\ast}\right\|_{2}^{2}$ $=\left\|\left(A+\widetilde{\mu}I\right)^{-1}A\mathbf{y}-\left(A+\mu^{\ast}I\right)^{-1}A\mathbf{y}\right\|_{2}^{2}=\sum_{i=1}^{d}v_{i}^{2}\left(\frac{1}{1+\frac{\widetilde{\mu}}{\lambda_{i}}}-\frac{1}{1+\frac{\mu^{\ast}}{\lambda_{i}}}\right)^{2}$
$\leq\sum_{i=1}^{d}v_{i}^{2}\left(1-\frac{1}{1+\frac{\left|\widetilde{\mu}-\mu^{\ast}\right|}{\lambda_{i}}}\right)^{2}\leq\sum_{i=1}^{d}v_{i}^{2}\left(\frac{\widetilde{\mu}-\mu^{\ast}}{\lambda_{i}}\right)^{2}\leq\left\|\mathbf{y}\right\|_{2}^{2}\left(\frac{\widetilde{\mu}-\mu^{\ast}}{\lambda_{d}}\right)^{2}<\zeta^{2}$
The first inequality is because $\lambda_{i}>0$, $\widetilde{\mu}\geq 0$ and $\mu^{\ast}\geq 0$; The second inequality employs the following inequality, $(1-\frac{1}{1+a})^{2}=\frac{a^{2}}{(a+1)^{2}}\leq a^{2}$ for any $a\geq 0$; The third inequality is uses $\sum_{i=1}^{d}v_{i}^{2}=\left\|\mathbf{y}\right\|_{2}^{2}$; The last inequality substitutes $|\widetilde{\mu}-\mu^{\ast}|<\frac{\lambda_{d}}{\left\|\mathbf{y}\right\|_{2}}\zeta$. ∎
Proof of Theorem˜3.
First, we justify the bisection. By Lemma˜10, the wanted zero satisfies $\mu^{\ast}\in((\frac{\left\|\mathbf{y}\right\|_{2}}{R}-1)\lambda_{d},(\frac{\left\|\mathbf{y}\right\|_{2}}{R}-1)\lambda_{1})=(a_{1},b_{1})$, which implies the initial interval of bisection (Line 3 in Algorithm˜4). Furthermore, $\rho$ monotonically decreases on $0,\infty)$, which implies the selection of $a_{t+1}$ and $b_{t+1}$ (Lines 4–6 in Algorithm˜4).
Next, we show the convergence of the bisection. With the interval length halving each iteration, based on the value of $T$ (Line 3 in Algorithm˜4), we have
$|\mu_{T+1}-\mu^{\ast}|\leq\frac{b_{1}-a_{1}}{2^{T}}=\frac{(\frac{\left\|\mathbf{y}\right\|_{2}}{R}-1)(\lambda_{1}-\lambda_{d})}{2^{T}}\leq\frac{(\frac{\left\|\mathbf{y}\right\|_{2}}{R}-1)(\overline{\lambda}-\underline{\lambda})}{2^{T}}\leq\frac{\underline{\lambda}}{\left\|\mathbf{y}\right\|_{2}}\zeta\leq\frac{\lambda_{d}}{\left\|\mathbf{y}\right\|_{2}}\zeta.$
Then by Lemma˜11, Algorithm˜4 achieves an error $\left\|\mathbf{x}-\mathbf{x}^{\ast}\right\|_{2}\leq\zeta$.
Finally, we bound the number of arithmetic operations. Note that choice 1 and choice 2 are equivalent as
$\left\|\left(A+\mu I\right)^{-1}A\mathbf{y}\right\|_{2}^{2}=\left\|P\left(C+\mu I\right)^{-1}P^{\mathsf{T}}PCP^{\mathsf{T}}\mathbf{y}\right\|_{2}^{2}=\left\|\left(C+\mu I\right)^{-1}\mathbf{w}\right\|_{2}^{2}.$
With choice 1, each iteration requires $d^{\omega}$ arithmetic operations, where the computational bottleneck lies in the matrix inversion $(A+\mu I)^{-1}$, resulting in a runtime of $O(d^{\omega}n)$. With choice 2, the tridiagonalization of $A$ requires $O(d^{3})$ arithmetic operations [Golub and Van Loan, 2013], and each iteration can solve $(C+\mu I)^{-1}\mathbf{w}$ using the Thomas algorithm with only $O(d)$ arithmetic operations [Golub and Van Loan, 2013], resulting in a runtime of $O(d^{3}+dn)$. ∎
Appendix C Proofs for Section˜4
In Section˜C.1, we prove Theorem˜4 based on Corollary 2 of Mehta [2017]. In Section˜C.2, we provide proofs for the ERM-based SXO methods discussed in Section˜4.3.
C.1 Proof of Theorem˜4
Proof.
First, we prove Eq.˜19. Without loss of generality, we consider $T\geq 3$. Then the high-probability excess risk bound of Eq.˜19 directly follows from Corollary 2 of Mehta [2017] by substituting LightONS’s regret in Theorem˜2.
Then, we verify that the choice of $T=\Theta\left(\frac{d}{\epsilon}\log\frac{d}{\epsilon}\log\frac{1}{\delta}\right)$ yields a high-probability excess risk of $O(\epsilon)$. Let $\Gamma_{\delta}=\frac{4\log T}{\delta}=O\left(\frac{1}{\delta}\log\frac{d}{\epsilon}+\frac{1}{\delta}\log\log\frac{1}{\delta}\right)$, we have
$F(\bar{\mathbf{x}}_{T})-\min_{\mathbf{x}\in\mathcal{X}}F(\mathbf{x})\leq O\left(\frac{\textsc{Reg}_{T}+\sqrt{\textsc{Reg}_{T}\cdot\log\Gamma_{\delta}}+\log\Gamma_{\delta}}{T}\right)\leq O\left(\frac{\textsc{Reg}_{T}+\log\Gamma_{\delta}}{T}\right).$
The first inequality is essentially Corollary 2 of Mehta [2017], and the second inequality follows from the fact that $\sqrt{ab}\leq\frac{a+b}{2}=O(a+b)$ for any positive terms $a$ and $b$. Recalling that $\textsc{Reg}_{T}=O(d\log T)=O\left(d\log\frac{d}{\epsilon}+d\log\log\frac{1}{\delta}\right)$, we have
$O\left(\frac{\textsc{Reg}_{T}}{T}\right)=O\left(\frac{d\left(\log\frac{d}{\epsilon}+\log\log\frac{1}{\delta}\right)}{\frac{d}{\epsilon}\log\frac{d}{\epsilon}\log\frac{1}{\delta}}\right)=O(\epsilon),$
and
$O\left(\frac{\log\Gamma_{\delta}}{T}\right)=O\left(\frac{\log\frac{1}{\delta}+\log\left(\log\frac{d}{\epsilon}+\log\log\frac{1}{\delta}\right)}{\frac{d}{\epsilon}\log\frac{d}{\epsilon}\log\frac{1}{\delta}}\right)=O\left(\frac{\epsilon}{d}\right)=O(\epsilon).$
Next, we prove Eq.˜20 from Eq.˜19. By the definition of expectation, we have
$\mathop{\mathbb{E}}\left[F(\bar{\mathbf{x}}_{T^{\prime}})-\min_{\mathbf{x}\in\mathcal{X}}F(\mathbf{x})\right]$ $\leq O\left(\frac{\textsc{Reg}_{T^{\prime}}+\sqrt{\textsc{Reg}_{T^{\prime}}\cdot\log\Gamma_{1/T^{\prime}}}+\Gamma_{1/T^{\prime}}}{T^{\prime}}\right)+\frac{\max\limits_{\mathbf{y}\in\mathcal{X}}F(\mathbf{y})-\min\limits_{\mathbf{x}\in\mathcal{X}}F(\mathbf{x})}{T^{\prime}}$
$\leq O\left(\frac{\textsc{Reg}_{T^{\prime}}+\Gamma_{1/T^{\prime}}}{T^{\prime}}\right)+\frac{DG}{T^{\prime}}\leq O\left(\frac{\textsc{Reg}_{T^{\prime}}+\Gamma_{1/T^{\prime}}}{T^{\prime}}\right)$
The above inequalities use the Lipschitzness of $F$ and the boundedness of $\mathcal{X}$. Similarly, we verify that the choice of $T^{\prime}=\Theta\left(\frac{d}{\epsilon}\log\frac{d}{\epsilon}\right)$ yields an in-expectation excess risk of $O(\epsilon)$ with $\frac{1}{\delta}=T^{\prime}=\Theta\left(\frac{d}{\epsilon}\log\frac{d}{\epsilon}\right)$. It suffices to note that $\Gamma_{1/T^{\prime}}=4T^{\prime}\log T^{\prime}=O\left(\frac{d}{\epsilon}\log^{2}\frac{d}{\epsilon}\right)$ and that
$O\left(\frac{\log\Gamma_{1/T^{\prime}}}{T^{\prime}}\right)=O\left(\frac{\log\frac{d}{\epsilon}}{\frac{d}{\epsilon}\log\frac{d}{\epsilon}}\right)=O\left(\frac{\epsilon}{d}\right)=O(\epsilon).$
Finally, the total runtime follows from Theorem˜2 by substituting $T=\Theta\left(\frac{d}{\epsilon}\log\frac{d}{\epsilon}\log\frac{1}{\delta}\right)$. ∎
C.2 Proofs for ERM-based SXO Methods
In this part, we construct a scenario showing that ERM-based SXO methods cannot break the $\widetilde{O}(d^{3}/\epsilon)$ runtime barrier, even when equipped with state-of-the-art offline convex optimization solvers.
Proposition 4.
Under Assumption˜4 and $\mathcal{X}=\mathcal{B}(1)$, let the stochastic functions take the form $f(\mathbf{x};\xi)=\phi\left(\mathbf{w}(\xi)^{\mathsf{T}}\mathbf{x}\right)$, where $\phi:\mathbb{R}\to\mathbb{R}$ is a black-box function, and $\mathbf{w}:\Xi\to\mathbb{R}^{d}$ is a fixed mapping. by [Koren and Levy, 2015, Mehta, 2017], obtaining an $\epsilon$-optimal solution for SXO reduces to obtaining an $O(\epsilon)$-optimal solution to the offline ($\alpha$-exp-concave) objective
$\widehat{F}(\mathbf{x})=\frac{1}{T}\sum_{t=1}^{T}\phi\left(\mathbf{w}(\xi_{t})^{\mathsf{T}}\mathbf{x}\right),$
where $T=\widetilde{O}(d/\epsilon)$ is the necessary sample size. Cutting-Plane Methods (CPM) of Lee et al. [2015], Jiang et al. [2020] solves this offline problem to $O(\epsilon)$-accuracy in $\widetilde{O}(d^{3}/\epsilon)$ time.
Proof of Proposition˜4.
CPM of Lee et al. [2015], Jiang et al. [2020] queries the gradient of $\widehat{F}$ for $\widetilde{O}(d)$ times, and each gradient query costs $O(dT)$ time due to the finite-sum structure of $\widehat{F}$ and the black-box nature of $\phi$. Therefore, the total runtime is $\widetilde{O}(d^{2}T+d^{3})=\widetilde{O}(d^{3}/\epsilon)$. ∎
Appendix D Proofs for Section˜5
In this section, we provide technical details on integrating LightONS into the applications in Section˜5 and their analysis.
D.1 Proofs for Gradient-Norm Adaptivity
The proof of Theorems˜6 and 5 closely follows that of their ONS counterparts with minor modifications.
Proof of Theorem˜5.
Based on the proof of Theorem 1 of Orabona et al. [2012], it suffices to bound the gradient norms of the surrogate loss by those of the original loss. We note that Appendix B.4.2 of Yang et al. [2024] also discusses the small-loss bounds of ONS with domain conversion, although with a different domain conversion. Since $\left\|\nabla g_{t}(\mathbf{y}_{t})\right\|_{2}\leq\left\|\nabla f_{t}(\mathbf{x}_{t})\right\|_{2}$ by Lemma˜3, we have
$G_{T,g}\triangleq\sum_{t=1}^{T}\left\|\nabla g_{t}(\mathbf{y}_{t})\right\|_{2}^{2}\leq\sum_{t=1}^{T}\left\|\nabla f_{t}(\mathbf{x}_{t})\right\|_{2}^{2}\triangleq G_{T,f}.$
Theorem˜2 implies the runtime with $\epsilon=d$ and the following regret bound:
$\textsc{Reg}_{T}(\mathbf{u})\leq\frac{d}{2\gamma_{0}}\log\left(1+\frac{G_{T,g}}{d\epsilon}\right)+\frac{\gamma_{0}\epsilon D^{2}}{8}\leq\frac{d}{2\gamma_{0}}\log\left(1+\frac{G_{T,f}}{d\epsilon}\right)+\frac{\gamma_{0}\epsilon D^{2}}{8}.$
The preceding inequality uses Jensen’s inequality and $\operatorname{tr}(A_{T})=d\epsilon+G_{T,g}$, i.e.,
$\log\operatorname{det}(A_{T})-\log\operatorname{det}(A_{0})\leq d\log\frac{\operatorname{tr}(A_{T})}{d}-\log\operatorname{det}(A_{0})=d\log\left(1+\frac{G_{T,g}}{d\epsilon}\right).$ (44)
Then Corollary 5 of Orabona et al. [2012] converts the gradient-norm adaptive bound to the small-loss bound with respect to $L_{T}$ as defined Eq.˜22 and completes the proof. ∎
Proof of Theorem˜6.
Theorem 8 of Cutkosky and Orabona [2018] hinges on their Algorithm 7, Lemmas 16 and 17, apart from the coin-betting framework. To prove Theorem˜6, we show how LightONS adapts their analysis with minimal changes.
  • Modifications to their Algorithm 7. Since the decisions of their ONS are intermediate decisions to maximize the “wealth” in the coin-betting framework instead of true decisions to minimize regret, we can ignore the improper-to-proper conversion and replace their ONS with LightONS.Core. Specifically, their ONS runs on $\mathcal{X}=\mathcal{B}(1/2)$ while LightONS.Core runs on $\mathcal{Y}=\mathcal{B}(3/4)$ with the hysteresis coefficient $k=3/2$.
  • Modifications to their Lemma 16. Their ONS’s domain $\mathcal{X}=\mathcal{B}(1/2)$ implies a curvature parameter $\gamma=\frac{2-\log 3}{2}$ while LightONS.Core’s domain $\mathcal{Y}=\mathcal{B}(3/4)$ implies $\gamma=\frac{6-\log 7}{18}$.
  • Modifications to their Lemma 17. Relaxing the radius from $1/2$ to $3/4$ enlarges constants in the regret. Nonetheless, the numerical constants in their Lemma 17 are loose enough to accommodate our changes, greatly simplifying our analysis. Following their proof, let $\mathbf{z}_{t}$ denote the gradient that LightONS.Core receives at time $t$, Theorem˜2 and Eq.˜44 imply
    $\textsc{Reg}_{T}\leq\frac{d}{2\gamma}\log\left(1+\frac{\sum_{t=1}^{T}\left\|\mathbf{z}_{t}\right\|_{2}^{2}}{d\epsilon}\right)+\frac{\gamma\epsilon D^{2}}{8}.$
    Plugging $D=1$, $\gamma=\frac{6-\log 7}{18}\in(0.225,0.226)$, $\epsilon=d$, and $\left\|\mathbf{z}_{t}\right\|_{2}^{2}\leq 16\left\|\nabla f_{t}(\mathbf{x}_{t})\right\|_{2}^{2}$ yields
    $\textsc{Reg}_{T}\leq d\left(\frac{5}{2}\log\left(1+\frac{16}{d^{2}}\sum_{t=1}^{T}\left\|\nabla f_{t}(\mathbf{x}_{t})\right\|_{2}^{2}\right)+\frac{1}{35}\right).$
    The preceding regret bound fully recovers their Lemma 17 when $d\geq 2$.
The runtime follows from Theorem˜2 with $\epsilon=d$. ∎
D.2 Proofs for Logistic Bandits
We propose Algorithm˜5 to replace the ONS-like subroutine in the original work. We emphasize that the learning of $\mathbf{w}_{t}$ does not suffer from improper learning concerns. Because the algorithm’s final output decision is the arm $\mathbf{x}_{t}$, rather than the estimated parameter $\mathbf{w}_{t}$. Consequently, Algorithm˜5 is free from the improper-to-proper conversion and resembles LightONS.Core.
Before proving Theorem˜7, we first present two critical lemmas that facilitate migrating the original ONS-like subroutine to Algorithm˜5.
Let $\ell_{t}(\mathbf{w})=-y_{t}\log\sigma(\mathbf{x}_{t}^{\mathsf{T}}\mathbf{w})-(1-y_{t})\log(1-\sigma(\mathbf{x}_{t}^{\mathsf{T}}\mathbf{w}))$ be the logistic loss function, then $\nabla\ell_{t}(\mathbf{w})=\left(\sigma(\mathbf{x}_{t}^{\mathsf{T}}\mathbf{w})-y_{t}\right)\mathbf{x}_{t}$ and $\nabla^{2}\ell_{t}(\mathbf{w})=\sigma^{\prime}(\mathbf{x}_{t}^{\mathsf{T}}\mathbf{w})\mathbf{x}_{t}\mathbf{x}_{t}^{\mathsf{T}}$. For consistency, we use the notation $\mathcal{W}=\mathcal{B}(D/2)$ as in the main text, instead of $\mathcal{W}=\mathcal{B}(S)$ as in the original work.
The following lemma resembles Lemma 1 of Zhang et al. [2025], showing that Algorithm˜5 implies an OMD-like regret decomposition form similar to that of original ONS-like subroutine.
Lemma 12.
Algorithm˜5 satisfies that, for any $\mathbf{u}\in\mathcal{B}(D/2)$
$\left\|\mathbf{w}_{t+1}-\mathbf{u}\right\|_{H_{t-1}}^{2}\leq 2\eta\nabla\widetilde{\ell}_{t}(\mathbf{w}_{t+1})^{\mathsf{T}}(\mathbf{u}-\mathbf{w}_{t+1})+\left\|\mathbf{w}_{t}-\mathbf{u}\right\|_{H_{t-1}}^{2}-\left\|\mathbf{w}_{t}-\mathbf{w}_{t+1}\right\|_{H_{t-1}}^{2},$
where $\widetilde{\ell}_{t}$ is the second-order Taylor expansion of $\ell_{t}$ at $\mathbf{w}_{t}$, i.e.,
$\widetilde{\ell}_{t}(\mathbf{w})\triangleq\ell_{t}(\mathbf{w}_{t})+\nabla\ell_{t}(\mathbf{w}_{t})^{\mathsf{T}}(\mathbf{w}-\mathbf{w}_{t})+\frac{1}{2}\left\|\mathbf{w}-\mathbf{w}_{t}\right\|_{\nabla^{2}\ell_{t}(\mathbf{w}_{t})}^{2}.$
Proof of Lemma˜12.
Before proceeding, we note that the descend-and-project update in Algorithm˜5 is equivalent to the OMD update, as Appendix D of Zhang et al. [2025] has shown. Specifically,
$\mathbf{w}_{t+1}$ $=\Pi^{\widetilde{H}_{t}}_{\mathcal{B}(D/2)}[\widehat{\mathbf{w}}_{t+1}]=\Pi^{\widetilde{H}_{t}}_{\mathcal{B}(D/2)}\left[\mathbf{w}_{t}-\frac{1}{\eta}\widetilde{H}_{t}^{-1}\nabla\ell_{t}(\mathbf{w}_{t})\right]$ (45)
$\iff\mathbf{w}_{t+1}$ $=\operatorname*{arg\,min}_{\mathbf{w}\in\mathcal{B}(D/2)}\widetilde{\ell}_{t}(\mathbf{w})+\frac{1}{2\eta}\left\|\mathbf{w}-\mathbf{w}_{t}\right\|_{H_{t-1}}^{2}$
and
$\mathbf{w}_{t+1}$ $=\widehat{\mathbf{w}}_{t+1}=\mathbf{w}_{t}-\frac{1}{\eta}\widetilde{H}_{t}^{-1}\nabla\ell_{t}(\mathbf{w}_{t})$ (46)
$\iff\mathbf{w}_{t+1}$ $=\operatorname*{arg\,min}_{\mathbf{w}\in\mathbb{R}^{d}}\widetilde{\ell}_{t}(\mathbf{w})+\frac{1}{2\eta}\left\|\mathbf{w}-\mathbf{w}_{t}\right\|_{H_{t-1}}^{2}$
To recover Lemma 1 of Zhang et al. [2025], we examine whether the following inequality holds
$\nabla\widetilde{\ell}_{t}(\mathbf{w}_{t+1})^{\mathsf{T}}\left(\mathbf{w}_{t+1}-\mathbf{u}\right)\leq\frac{1}{2\eta}\left(\left\|\mathbf{w}_{t}-\mathbf{u}\right\|_{H_{t-1}}^{2}-\left\|\mathbf{w}_{t+1}-\mathbf{u}\right\|_{H_{t-1}}^{2}-\left\|\mathbf{w}_{t+1}-\mathbf{w}_{t}\right\|_{H_{t-1}}^{2}\right).$ (47)
When $\left\|\widehat{\mathbf{w}}_{t+1}\right\|_{2}\leq kD/2$ and the Mahalanobis projection is not performed, by Eq.˜45 we have
$\nabla_{\mathbf{w}=\mathbf{w}_{t+1}}\left(\widetilde{\ell}_{t}(\mathbf{w})+\frac{1}{2\eta}\left\|\mathbf{w}-\mathbf{w}_{t}\right\|_{H_{t-1}}^{2}\right)^{\mathsf{T}}(\mathbf{w}_{t+1}-\mathbf{u})=\bm{0}^{\mathsf{T}}(\mathbf{w}_{t+1}-\mathbf{u})=0.$
Rearranging terms, we have
$\nabla\widetilde{\ell}_{t}(\mathbf{w}_{t+1})^{\mathsf{T}}\left(\mathbf{w}_{t+1}-\mathbf{u}\right)$ $=-\frac{1}{\eta}(\mathbf{w}_{t+1}-\mathbf{w}_{t})^{\mathsf{T}}H_{t-1}(\mathbf{w}_{t+1}-\mathbf{u})$
$=\frac{1}{2\eta}\left(\left\|\mathbf{w}_{t}-\mathbf{u}\right\|_{H_{t-1}}^{2}-\left\|\mathbf{w}_{t+1}-\mathbf{u}\right\|_{H_{t-1}}^{2}-\left\|\mathbf{w}_{t+1}-\mathbf{w}_{t}\right\|_{H_{t-1}}^{2}\right)$
which means that Eq.˜47 holds with equality. When $\left\|\widehat{\mathbf{w}}_{t+1}\right\|_{2}>kD/2$ and the Mahalanobis projection is performed, by Eq.˜46, we have
$\nabla\left(\widetilde{\ell}_{t}(\mathbf{w}_{t+1})+\frac{1}{2\eta}\left\|\mathbf{w}_{t+1}-\mathbf{w}_{t}\right\|_{H_{t-1}}^{2}\right)^{\mathsf{T}}(\mathbf{w}_{t+1}-\mathbf{u})\leq 0.$
Rearranging terms, we have
$\nabla\widetilde{\ell}_{t}(\mathbf{w}_{t+1})^{\mathsf{T}}\left(\mathbf{w}_{t+1}-\mathbf{u}\right)$ $\leq-\frac{1}{\eta}(\mathbf{w}_{t+1}-\mathbf{w}_{t})^{\mathsf{T}}H_{t-1}(\mathbf{w}_{t+1}-\mathbf{u})$
$=\frac{1}{2\eta}\left(\left\|\mathbf{w}_{t}-\mathbf{u}\right\|_{H_{t-1}}^{2}-\left\|\mathbf{w}_{t+1}-\mathbf{u}\right\|_{H_{t-1}}^{2}-\left\|\mathbf{w}_{t+1}-\mathbf{w}_{t}\right\|_{H_{t-1}}^{2}\right)$
which means that Eq.˜47 holds.
Therefore, combining both cases, we have Eq.˜47 always holds. ∎
The runtime of Algorithm˜5 is given by the following lemma.
Lemma 13.
Algorithm˜5 has a runtime of $\widetilde{O}(d^{2}T+d^{\omega}\cdot\min\{\sqrt{\kappa dT},T\})$.
Proof of Lemma˜13.
Compared with the runtime of LightONS Theorem˜2, the only difference the dependence is that $\kappa$ appears in the number of Mahalanobis projections. It suffices to show how the logistic loss function affects the analysis of Lemma˜2. Specifically, we need to bound the quantity
$\Phi^{\prime}_{T}\triangleq\sum_{t=1}^{T}\left\|\frac{1}{\eta}\widetilde{H}_{t}^{-1}\nabla\ell_{t}(\mathbf{w}_{t})\right\|_{2}^{2}.$
By the update rule of Algorithm˜5 and that $\eta>1$ in [Zhang et al., 2025], we have
$\widetilde{H}_{t}=\epsilon I+\sum_{i=1}^{t-1}\sigma^{\prime}(\mathbf{x}_{i}^{\mathsf{T}}\mathbf{w}_{i+1})\mathbf{x}_{i}\mathbf{x}_{i}^{\mathsf{T}}+\eta\sigma^{\prime}(\mathbf{x}_{t}^{\mathsf{T}}\mathbf{w}_{t})\mathbf{x}_{t}\mathbf{x}_{t}^{\mathsf{T}}\succ\lambda I+\sum_{i=1}^{t}\frac{1}{\kappa}\mathbf{x}_{i}\mathbf{x}_{i}^{\mathsf{T}},$
where $\kappa=\max_{\mathbf{x}\in\mathcal{X},\mathbf{w}\in\mathcal{W}}1/\sigma^{\prime}(\mathbf{x}^{\mathsf{T}}\mathbf{w})$. Since $\left|\sigma(\mathbf{x}_{t}^{\mathsf{T}}\mathbf{w}_{t})-y_{t}\right|\leq 1$, we have
$\left\|\widetilde{H}_{t}^{-1}\nabla\ell_{t}(\mathbf{w}_{t})\right\|_{2}^{2}$ $=\left\|\widetilde{H}_{t}^{-1}(\sigma(\mathbf{x}_{t}^{\mathsf{T}}\mathbf{w}_{t})-y_{t})\mathbf{x}_{t}\right\|_{2}^{2}\leq\left\|\widetilde{H}_{t}^{-1}\mathbf{x}_{t}\right\|_{2}^{2}$
$<\left\|\left(\lambda I+\sum_{i=1}^{t}\frac{1}{\kappa}\mathbf{x}_{i}\mathbf{x}_{i}^{\mathsf{T}}\right)^{-1}\mathbf{x}_{t}\right\|_{2}^{2}=\kappa^{2}\left\|\left(\kappa\lambda I+\sum_{i=1}^{t}\mathbf{x}_{i}\mathbf{x}_{i}^{\mathsf{T}}\right)^{-1}\mathbf{x}_{t}\right\|_{2}^{2}$
Finally, summing up the preceding inequality with Lemma˜6, we have
$\Phi^{\prime}_{T}<\frac{\kappa^{2}}{\eta^{2}}\sum_{t=1}^{T}\left\|\left(\kappa\lambda I+\sum_{i=1}^{t}\mathbf{x}_{i}\mathbf{x}_{i}^{\mathsf{T}}\right)^{-1}\mathbf{x}_{t}\right\|_{2}^{2}\leq\frac{\kappa d}{\eta^{2}\lambda}.$
We note that $H_{t}$ and $\widetilde{H}_{t}$ admit rank-one updates similar to those in LightONS.
Therefore, by reusing the proof of Lemma˜2, we obtain that the runtime of Algorithm˜5 is $\widetilde{O}(d^{2}T+d^{\omega}\cdot\min\{\sqrt{\kappa dT},T\})$. ∎
Algorithm 5 LightONS for [Zhang et al., 2025]
0: domain $\mathcal{W}=\mathcal{B}(D/2)$, regularization coefficient $\lambda$, inverse step size $\eta$.
1: Initialize $H_{0}=\epsilon I$; $\mathbf{w}=\bm{0}$.
2:for $t=1,\dots,T$ do
3:  Update the lower confidence bound function as in Algorithm 1 of Zhang et al. [2025].
4:  Select the arm $\mathbf{x}_{t}$ as the Algorithm 1 of Zhang et al. [2025] and observe the loss $y_{t}$.
5:  $\widetilde{H}_{t}=H_{t-1}+\eta\nabla^{2}\ell_{t}(\mathbf{w}_{t})$.
6:  $\widehat{\mathbf{w}}_{t+1}=\mathbf{w}_{t}-\frac{1}{\eta}\widetilde{H}_{t}^{-1}\nabla\ell_{t}(\mathbf{w}_{t})$.
7:  $\mathbf{w}_{t+1}=\begin{cases}\widehat{\mathbf{w}}_{t+1}&\text{if}~\left\|\widehat{\mathbf{w}}_{t+1}\right\|_{2}\leq kD/2\\ \Pi^{\widetilde{H}_{t}}_{\mathcal{B}(D/2)}[\widehat{\mathbf{w}}_{t+1}]&\text{otherwise}\end{cases}$.
8:  $H_{t}=H_{t-1}+\nabla^{2}\ell_{t}(\mathbf{w}_{t+1})$.
9:end for
Based on the preceding lemmas, we are ready to prove Theorem˜7.
Proof of Theorem˜7.
The upper-confidence-bound-based pseudo-regret analysis in [Zhang et al., 2025] primarily relies on their Theorem 1, which constitutes their Lemmas 4, 5, and 6. To prove Theorem˜7, we examine how replacing the ONS-like subroutine with Algorithm˜5 affects these lemmas except the moderate expansion of the diameter from $D$ to $D^{\prime}=kD=2D$.
  • Modifications to their Lemma 4. The original Lemma 4 is supported by the original Lemma 1 and local relaxation of generalized linear models. Since the local relaxation is independent of the specific update rules of $\mathbf{w}_{t}$, we only need to verify that the original Lemma 1 in still holds when using LightONS as a replacement, as demonstrated in Lemma˜12.
  • Modifications to their Lemma 5 and 6. The original Lemma 5 and 6 depend solely on the structure of the covariance matrix, i.e., $H_{t}=\lambda I+\sum_{i=1}^{t}\sigma^{\prime}(\mathbf{x}_{i}^{\mathsf{T}}\mathbf{w}_{i+1})\mathbf{x}_{i}\mathbf{x}_{i}^{\mathsf{T}}$, and are independent of the specific update rules of $\mathbf{w}_{t}$. Therefore, plugging LightONS does not affect the original Lemma 5 and 6.
Finally, the runtime of Algorithm˜5 follows from Lemma˜13. With the overhead $O(d^{2}KT)$ for selecting arms as in [Zhang et al., 2025], the total runtime is $\widetilde{O}(d^{2}KT+d^{\omega}\cdot\min\{\sqrt{\kappa dT},T\})$. ∎
We remark that extending Algorithm˜5 from the binary logistic bandits to the generalized linear bandits directly follows from [Zhang et al., 2025].
D.3 Proofs for Memory-Efficient OXO
We propose LightONS.Sketch in Algorithm˜6, which integrates LightONS with sketching. The key difference between LightONS.Sketch and LightONS is the storage strategy. Instead of storing full matrices $A_{t}\in\mathbb{R}^{d\times d}$ and $V_{t}\in\mathbb{R}^{d\times d}$, LightONS.Sketch maintains compact sketches $S_{t}\in\mathbb{R}^{2d^{\prime}\times d}$ and $R_{t}\in\mathbb{R}^{2d^{\prime}\times 2d^{\prime}}$. Each row of $S_{t}$ stores a principal gradient component, while $R_{t}$ plays a role analogous to $V_{t}$ in Eq.˜18. The full matrix and its inverse are reconstructed as
$\widetilde{A}_{t}=\epsilon I+S_{t}^{\mathsf{T}}S_{t},\quad\widetilde{A}_{t}^{-1}=\frac{1}{\epsilon}\left(I-S_{t}^{\mathsf{T}}R_{t}S_{t}\right),\quad R_{t}=\left(\epsilon I+S_{t}S_{t}^{\mathsf{T}}\right)^{-1}.$
This relationship follows from the Sherman-Morrison-Woodbury formula:
$(A+BCD)^{-1}=A^{-1}-A^{-1}B(C^{-1}+DA^{-1}B)^{-1}DA^{-1}.$
Following [Luo et al., 2016], LightONS.Sketch uses Fast Frequent Directions [Ghashami et al., 2016] in Algorithm˜7 to update $S_{t}$ and $R_{t}$.
Before proving Theorem˜8, we first introduce three lemmas to characterize the error introduced by sketching. Specifically, Lemma˜14 bound the error for regret analysis with Lemma˜5, while Lemma˜15 bound the error for projection-count analysis with Lemma˜6. Lemma˜16 bounds the total error with the spectrum of the Hessian-related matrix.
The notation $\Delta_{t}$ in the lemmas denotes the error introduced by sketching at the $t$-th round. When SVD is performed $\Delta_{t}=\frac{2d^{\prime}}{\epsilon}\sigma_{d^{\prime}}(S_{t})^{2}$; Otherwise, $\Delta_{t}=0$. Here $\sigma_{i}(S_{t})$ is the $i$-th greatest singular value of $S_{t}$.
Lemma 14.
At the $t$-th round of Algorithm˜6, in Algorithm˜7, $\Delta_{t}$ satisfies:
$\left\|\nabla g_{t}(\mathbf{y}_{t})\right\|_{\widetilde{A}_{t}^{-1}}^{2}\leq\left\langle\widetilde{A}_{t}^{-1},\widetilde{A}_{t}-\widetilde{A}_{t-1}\right\rangle_{\operatorname{F}}+\Delta_{t}.$
Proof of Lemma˜14.
Let $\left\langle\cdot,\cdot\right\rangle_{\operatorname{F}}$ denote the inner product induced by the Frobenius matrix norm, let $(\cdot)_{:,i}$ denote the $i$-th column of the matrix, and let $\bar{U}_{t}\bar{\Sigma}_{t}\bar{V}_{t}^{\mathsf{T}}=S_{t}$ be the full SVD with all $2d^{\prime}$ singular values. When SVD is not performed, then $\widetilde{A}_{t}=\widetilde{A}_{t-1}+\nabla g_{t}(\mathbf{y}_{t})\nabla g_{t}(\mathbf{y}_{t})^{\mathsf{T}}$ and $\Delta_{t}=0$ trivially holds. When SVD is performed,
$\left\langle\widetilde{A}_{t}^{-1},\widetilde{A}_{t-1}+\nabla g_{t}(\mathbf{y}_{t})\nabla g_{t}(\mathbf{y}_{t})^{\mathsf{T}}-\widetilde{A}_{t}\right\rangle_{\operatorname{F}}=\sum_{i=1}^{2d^{\prime}}\min\left\{(\bar{\Sigma}_{t})_{i,i}^{2},(\bar{\Sigma}_{t})_{d^{\prime},d^{\prime}}^{2}\right\}\left\|(\bar{V}_{t})_{:,i}\right\|_{\widetilde{A}_{t}^{-1}}^{2}\leq\underbrace{\frac{2d^{\prime}}{\epsilon}(\bar{\Sigma}_{t})_{d^{\prime},d^{\prime}}^{2}}_{\Delta_{t}}.$
The inequality uses the fact that $\widetilde{A}_{t}\succeq\epsilon I$ and $\left\|(\bar{V}_{t})_{:,i}\right\|_{\widetilde{A}_{t}^{-1}}^{2}\leq\left\|(\bar{V}_{t})_{:,i}\right\|_{\left(\epsilon I\right)^{-1}}^{2}=\frac{1}{\epsilon}$. Substituting $(\bar{\Sigma}_{t})_{i,i}=\sigma_{i}(S_{t})$ completes the proof. ∎
Lemma 15.
At the $t$-th round of Algorithm˜6, in Algorithm˜7, $\Delta_{t}$ also satisfies:
$\left\|\nabla g_{t}(\mathbf{y}_{t})\right\|_{\widetilde{A}_{t}^{-2}}^{2}=\left\langle\widetilde{A}_{t}^{-2},\widetilde{A}_{t}-\widetilde{A}_{t-1}\right\rangle_{\operatorname{F}}+\frac{\Delta_{t}}{\epsilon}.$
We omit proof of Lemma˜15, as it directly reuses the proof of Lemma˜14.
Lemma 16 (Theorem 1.1 and Section 3 of Ghashami et al. [2016]).
In Algorithm˜6, the total error $\Delta_{1:T}$ satisfies that, with $\lambda_{i}$ is the $i$-th greatest eigenvalue, for any $j\in[d^{\prime}]$,
$\Delta_{1:T}\triangleq\sum_{t=1}^{T}\Delta_{t}\leq\frac{2d^{\prime}}{(d^{\prime}-j+1)\epsilon}\sum_{i=j}^{d}\lambda_{i}\left(\sum_{t=1}^{T}\nabla g_{t}(\mathbf{y}_{t})\nabla g_{t}(\mathbf{y}_{t})^{\mathsf{T}}\right).$
Algorithm 6 LightONS.Sketch
0: preconditioner coefficient $\epsilon$, hysteresis coefficient $k$, dimension to reduce to $d^{\prime}$.
1: Initialize $\gamma^{\prime}=\frac{1}{2}\min\left\{\frac{1}{c_{f}c_{g}DG},\frac{4}{c_{f}c_{g}(k+1)DG},\alpha\right\}$; $S_{0}=O_{2d^{\prime}\times d}$; $R_{0}=\frac{1}{\epsilon}I_{2d^{\prime}\times 2d^{\prime}}$; $\mathbf{x}_{1}=\mathbf{y}_{1}=\bm{0}$.
2:for $t=1,\dots,T$ do
3:  Observe $\nabla f_{t}(\mathbf{x}_{t})$.
4:  Construct $\nabla g_{t}(\mathbf{y}_{t})$ satisfying Condition˜1.
[…]
6:  $\widehat{\mathbf{y}}_{t+1}=\mathbf{y}_{t}-\frac{1}{\gamma^{\prime}}\widetilde{A}_{t}^{-1}\nabla g_{t}(\mathbf{y}_{t})$, where $\widetilde{A}_{t}^{-1}=\frac{1}{\epsilon}(I-S_{t}^{\mathsf{T}}R_{t}S_{t})$.
7:  $\mathbf{y}_{t+1}=\begin{cases}\widehat{\mathbf{y}}_{t+1}&\text{if}~\left\|\widehat{\mathbf{y}}_{t+1}\right\|_{2}\leq kD/2\\ \Pi^{\widetilde{A}_{t}}_{\mathcal{B}(D/2)}[\widehat{\mathbf{y}}_{t+1}]&\text{otherwise}\end{cases}$, where $\widetilde{A}_{t}=\epsilon I+S_{t}^{\mathsf{T}}S_{t}$.
8:  $\mathbf{x}_{t+1}=\Pi_{\mathcal{X}}[\mathbf{y}_{t+1}]$.
9:end for
Algorithm 7 Fast Frequent Directions in [Ghashami et al., 2016]
0: new gradient $\nabla g_{t}(\mathbf{y}_{t})$, frequent directions $S_{t-1}$, low-dimension inverse $R_{t-1}$.
0: updated frequent directions $S_{t}$, updated low-dimension inverse $R_{t}$.
[…]
2:if $S_{t}$ still has all-zero rows then
3:  $R_{t}=\left(\epsilon I+S_{t}S_{t}^{\mathsf{T}}\right)^{-1}$.
4:else
5:  $U_{t}\Sigma_{t}V_{t}^{\mathsf{T}}=S_{t}$, truncated SVD with top $d^{\prime}$ singular values.
[…]
7:  $R_{t}=\operatorname{diag}\left(\frac{1}{\epsilon+(\Sigma_{t})_{1,1}^{2}-(\Sigma_{t})_{d^{\prime},d^{\prime}}^{2}},\dots,\frac{1}{\epsilon+(\Sigma_{t})_{d^{\prime},d^{\prime}}^{2}-(\Sigma_{t})_{d^{\prime},d^{\prime}}^{2}},\frac{1}{\epsilon},\dots,\frac{1}{\epsilon}\right)$.
8:end if
With the help these lemmas, we can prove Theorem˜7 by bounding difference between LightONS and LightONS.Sketch.
Proof of Theorem˜8.
We note that Lemma˜8 holds for Algorithm˜6, as the proof of Lemma˜8 only requires $A_{t}$ to be positive-definite. Thus by Eqs.˜38 and 13 we obtain
[…]
$\leq$ $\frac{1}{2\gamma^{\prime}}\left\|\nabla g_{t}(\mathbf{y}_{t})\right\|_{\widetilde{A}_{t}^{-1}}^{2}+\frac{\gamma^{\prime}}{2}\left\|\mathbf{y}_{t}-\mathbf{u}\right\|_{\widetilde{A}_{t-1}}^{2}-\frac{\gamma^{\prime}}{2}\left\|\mathbf{y}_{t+1}-\mathbf{u}\right\|_{\widetilde{A}_{t}}^{2}$
where the last inequality uses the fact that $\widetilde{A}_{t}\preceq\widetilde{A}_{t-1}+\nabla g_{t}(\mathbf{y}_{t})\nabla g_{t}(\mathbf{y}_{t})^{\mathsf{T}}$, which is ensured by Algorithm˜7. Then plugging Lemma˜14 into the preceding inequality yields
$f_{t}(\mathbf{x}_{t})-f_{t}(\mathbf{u})$ $\leq\frac{1}{2\gamma^{\prime}}\left(\left\langle\widetilde{A}_{t}^{-1},\widetilde{A}_{t}-\widetilde{A}_{t-1}\right\rangle_{\operatorname{F}}+\Delta_{t}\right)+\frac{\gamma^{\prime}}{2}\left\|\mathbf{y}_{t}-\mathbf{u}\right\|_{\widetilde{A}_{t-1}}^{2}-\frac{\gamma^{\prime}}{2}\left\|\mathbf{y}_{t+1}-\mathbf{u}\right\|_{\widetilde{A}_{t}}^{2}$
$\leq\frac{1}{2\gamma^{\prime}}\left(\log\operatorname{det}(\widetilde{A}_{t})-\log\operatorname{det}(\widetilde{A}_{t-1})+\Delta_{t}\right)+\frac{\gamma^{\prime}}{2}\left\|\mathbf{y}_{t}-\mathbf{u}\right\|_{\widetilde{A}_{t-1}}^{2}-\frac{\gamma^{\prime}}{2}\left\|\mathbf{y}_{t+1}-\mathbf{u}\right\|_{\widetilde{A}_{t}}^{2}$
where the equality uses the fact that $\left\langle X^{-1},X-Y\right\rangle_{\operatorname{F}}\leq\log\operatorname{det}(X)-\log\operatorname{det}(Y)$, which comes from the proof of Lemmas˜5 and 6. Telescoping the preceding inequality and discarding the negative terms establishes the desired regret bound Eq.˜32:
$\sum_{t=1}^{T}\left(f_{t}(\mathbf{x}_{t})-f_{t}(\mathbf{u})\right)\leq\frac{1}{2\gamma^{\prime}}\log\frac{\operatorname{det}(\widetilde{A}_{T})}{\operatorname{det}(\widetilde{A}_{0})}+\frac{1}{2\gamma^{\prime}}\Delta_{1:T}+\frac{\gamma^{\prime}}{2}\left\|\mathbf{y}_{1}-\mathbf{u}\right\|_{\widetilde{A}_{0}}^{2}.$
The logarithmic term is further bounded with Jensen’s inequality, differing slightly from Lemma˜5 due to the number of non-zero eigenvalues:
$\log\frac{\operatorname{det}(\widetilde{A}_{T})}{\operatorname{det}(\widetilde{A}_{0})}=\sum_{i=1}^{2d^{\prime}}\log\left(1+\frac{\sigma_{i}^{2}(S_{T})}{\epsilon}\right)\leq 2d^{\prime}\log\left(1+\frac{\left\|S_{T}\right\|_{\operatorname{F}}^{2}}{2d^{\prime}\epsilon}\right)\leq 2d^{\prime}\log\left(1+\frac{G^{2}}{2d^{\prime}\epsilon}T\right).$
Finally, the runtime Eq.˜34 follows from the following two parts:
  • Runtime aside from Algorithm˜7. Following the same analysis of Lemmas˜2 and 13, overall runtime aside from Algorithm˜7 is $O((k-1)^{-1}\sqrt{(d+\Delta_{1:T})T/\epsilon}\cdot d^{\omega}\log T)$ due to FastProj. It suffices to verify the following inequality which follows from Lemma˜15:
    $\Phi^{\prime\prime}_{T}\triangleq\sum_{t=1}^{T}\left\|\frac{1}{\gamma^{\prime}}\widetilde{A}_{t}^{-1}\nabla g_{t}(\mathbf{y}_{t})\right\|_{2}^{2}\leq\frac{1}{\gamma^{\prime 2}}\frac{d+\Delta_{1:T}}{\epsilon}.$
  • Runtime of Algorithm˜7. After SVD is performed, the last $d^{\prime}+1$ rows of $S_{t}$ are all-zero, thus the SVD is performed at most $\lceil T/(d^{\prime}+1)\rceil$ times. The runtime of SVD is $O(d^{\prime 2}d\log T)$ [Golub and Van Loan, 2013] to achieve the desired accuracy that does not affect the regret bound.  [^4] The runtime of SVD can be improved to $O\big(d^{\prime 2}(d+\log T)\big)$ and the factor $O(\log T)$ can be independent of the minimal singular value gap of $S_{t}$ and only depends on the scale of singular values of $S_{t}$ [Parlett, 1998]. The runtime of updating $R_{t}$ is $O(d^{\prime}d)$, as this can be implemented with twice rank-one updates:
    $(A+\mathbf{u}\mathbf{v}^{\mathsf{T}})^{-1}=A^{-1}-\frac{1}{1+\mathbf{v}^{\mathsf{T}}A^{-1}\mathbf{u}}A^{-1}\mathbf{u}\mathbf{v}^{\mathsf{T}}A^{-1}.$
    Specifically,
    $\quad R_{t}^{-1}=\epsilon I+S_{t}S_{t}^{\mathsf{T}}=\epsilon I+\left(S_{t-1}+\mathbf{e}_{i_{t}}\nabla g_{t}(\mathbf{y}_{t})^{\mathsf{T}}\right)\left(S_{t-1}+\mathbf{e}_{i_{t}}\nabla g_{t}(\mathbf{y}_{t})^{\mathsf{T}}\right)^{\mathsf{T}}$
    $=\underbrace{\epsilon I+S_{t-1}S_{t-1}^{\mathsf{T}}}_{R_{t-1}^{-1}}+\mathbf{e}_{i_{t}}\underbrace{\left(S_{t-1}\nabla g_{t}(\mathbf{y}_{t})\right)^{\mathsf{T}}}_{\mathbf{a}_{t}^{\mathsf{T}}}+\underbrace{\left(S_{t-1}\nabla g_{t}(\mathbf{y}_{t})+\mathbf{e}_{i_{t}}\nabla g_{t}(\mathbf{y}_{t})\nabla g_{t}(\mathbf{y}_{t})^{\mathsf{T}}\right)}_{\mathbf{b}_{t}}\mathbf{e}_{i_{t}}^{\mathsf{T}}$
    where $\mathbf{e}_{i_{t}}$, $\mathbf{a}_{t}$ and $\mathbf{b}_{t}$ are $2d^{\prime}$-dimensional vectors. Therefore, the overall runtime of Algorithm˜7 is $O(d^{\prime 2}d(\log T)\lceil T/(d^{\prime}+1)\rceil+d^{\prime}dT)=O(d^{\prime}dT\log T)$.