### v2

# A Simple, Optimal and Efficient Algorithm for Online Exp-Concave Optimization

Abstract
Online eXp-concave Optimization (OXO) is a fundamental problem in online learning, where the goal is to minimize regret when loss functions are exponentially concave. The standard algorithm, Online Newton Step (ONS), guarantees an optimal $O(d\log T)$ regret, where $d$ is the dimension and $T$ is the time horizon. Despite its simplicity, ONS may face a computational bottleneck due to the Mahalanobis projection at each round. This step costs $\Omega(d^{\omega})$ arithmetic operations for bounded domains, even for simple domains such as the unit ball, where $\omega\in(2,3]$ is the matrix-multiplication exponent. As a result, the total runtime can reach $\widetilde{O}(d^{\omega}T)$, particularly when iterates frequently oscillate near the domain boundary. This paper proposes a simple variant of ONS, called LightONS, which reduces the total runtime to $O(d^{2}T+d^{\omega}\sqrt{T\log T})$ while preserving the optimal regret. Deploying LightONS with the online-to-batch conversion implies a method for stochastic exp-concave optimization with runtime $\widetilde{O}(d^{3}/\epsilon)$, thereby answering an open problem posed by Koren [2013]. The design leverages domain-conversion techniques from parameter-free online learning and defers expensive Mahalanobis projections until necessary, thereby preserving the elegant structure of ONS and enabling LightONS to act as an efficient plug-in replacement in broader scenarios, including gradient-norm adaptivity, parametric stochastic bandits, and memory-efficient OXO.
1 Introduction
Online Convex Optimization (OCO) provides a versatile framework for online learning, with deep connections to stochastic optimization, game theory, and information theory [Cesa-Bianchi and Lugosi, 2006; Hazan, 2016]. Online eXp-concave Optimization (OXO), is an important instance, where loss functions are exponentially concave (exp-concave), i.e., $\exp(-\alpha f(\cdot))$ is concave for some $\alpha>0$ with $f$ denoting the loss function. Exp-concavity naturally arises in many machine learning applications such as linear/logistic regression [Foster, 1991; Vovk, 1997; Foster et al., 2018], portfolio selection [Cover, 1991], linear-quadratic regulator control [Foster and Simchowitz, 2020], and so on. From a theoretical perspective, it introduces rich structures beyond convexity, allowing algorithms to exploit the local geometry of the loss landscape, often through local norms induced along the optimization trajectory. Such structures yield sharper statistical guarantees: Exp-concave losses admit a minimax-optimal regret of $O(d\log T)$ [Ordentlich and Cover, 1998], an exponential improvement over the $\Omega(\sqrt{T})$ lower bound for convex losses [Abernethy et al., 2008].
Online Newton Step (ONS) [Hazan et al., 2007] is the de facto standard for OXO, achieving an optimal regret bound of $O(d\log T)$ with constant time and space per round. ONS exhibits remarkable simplicity, which has driven many advances in diverse optimization settings [Orabona et al., 2012; Luo et al., 2016; Cutkosky and Orabona, 2018] and in machine learning applications even beyond regret minimization, such as generalized linear bandits [Zhang et al., 2016; Jun et al., 2017; Zhang et al., 2025]. The important insight of ONS is to leverage exp-concavity by maintaining a Hessian-related matrix that captures the local geometry of the optimization trajectory, which is crucial for a balance between statistical optimality and computational practicality. Indeed, all known optimal OXO algorithms that avoid ONS-like Hessian maintenance, such as Exponential Weight Online Optimization (EWOO) [Hazan et al., 2007], incur prohibitive time complexity $O(T^{24})$ due to integrating over the domain [Bubeck et al., 2018].
However, the time-varying Hessian-related matrix in ONS necessitates a Mahalanobis projection at each round to ensure feasibility, which introduces the computational bottleneck. Specifically, the Mahalanobis projection solves the quadratic program $\Pi_{\mathcal{X}}^{M}[\mathbf{y}]=\operatorname*{arg\,min}_{\mathbf{x}\in\mathcal{X}}(\mathbf{x}-\mathbf{y})^{\mathsf{T}}M(\mathbf{x}-\mathbf{y})$ for some positive-definite and symmetric matrix $M$ and compact convex domain $\mathcal{X}\subseteq\mathbb{R}^{d}$. When the domain $\mathcal{X}$ is bounded, best known implementations require $\Omega(d^{\omega})$ arithmetic operations due to matrix factorizations such as matrix square root [Golub and Van Loan, 2013], where $\omega\in(2,3]$ is the matrix-multiplication exponent.  [^1] The factorization underlying Mahalanobis projections is related to eigendecomposition, equivalent to finding roots of a $d$-degree polynomial, which is not exactly solvable by finitely many arithmetic operations when $d\geq 5$. Even for simple domains such as the unit ball and the probability simplex, Mahalanobis projection requires $\widetilde{O}(d^{\omega})$ time (see Section˜B.3) and $\widetilde{O}(d^{\omega+0.5})$ time (via interior-point methods [Nesterov and Nemirovskii, 1994]), respectively. Thus, the total runtime of ONS can reach $\widetilde{O}(d^{\omega}T)$, particularly when iterates frequently oscillate near the domain boundary. Although Alman et al. [2025] establish $\omega<2.3714$ theoretically, linear algebra libraries typically operate with $\omega=3$, yielding $\widetilde{O}(d^{3}T)$ time in practice. In contrast, for convex and strongly convex online optimization, Online Gradient Descent (OGD) [Zinkevich, 2003] requires only Euclidean projections and achieves minimax-optimal regret with a runtime of $O(dT)$ for these simple domains.
[…]
Related work.
The pursuit for computationally efficient OXO algorithms can be divided into two main research lines. (i) The first, which we focus on in this paper, aims to minimize runtime while preserving optimal regret. The most closely related work is by Mhammedi and Gatmiry [2023], who proposed OQNS (Online Quasi-Newton Steps), achieving the optimal regret $O(d\log T)$ with runtime $O(d^{2}T\log T+d^{\omega}\sqrt{T\log T})$. They employ a log-barrier to eliminate Mahalanobis projections, transferring the computational burden to Hessian-inverse evaluations under log-barrier regularization. However, their method departs from canonical algorithmic frameworks such as online mirror descent (OMD) and Follow-the-Regularized-Leader (FTRL) [Orabona, 2019], making it difficult to accommodate diverse local norms and thereby limiting its applicability beyond standard OXO. A concrete comparison of our method with OQNS is deferred to Section˜5. (ii) The second line aims to reduce regret within a time or space budget, for example via matrix sketching techniques [Luo et al., 2016]. These methods achieve linear-in-$d$ runtime and working memory, albeit under additional assumptions on the loss functions and domains. This line also includes projection-free methods that trade statistical optimality for computational gains, leading to suboptimal regret bounds like $O(T^{2/3})$ [Garber and Kretzu, 2023; Wan et al., 2022].
Table 1: Algorithmic upper bounds on regret and total runtime of OXO algorithms with respect to $d$ and $T$ over a simple domain of the unit ball. “ONS-like” indicates whether the method can be integrated into other settings where ONS serves as the backbone, including gradient-norm adaptivity, parametric stochastic bandits, and memory-efficient OXO.
Algorithm Regret Total Runtime ONS-like
OGD [Zinkevich, 2003] $\sqrt{T}$ dTdT
ONS [Hazan et al., 2007] $d\log{T}$ $d^{\omega}T\log T$
OQNS [Mhammedi and Gatmiry, 2023] $d\log{T}$ $d^{2}T\log T+d^{\omega}\sqrt{T\log T}$
LightONS (This Paper) $d\log{T}$ $d^{2}T+d^{\omega}\sqrt{T\log T}$
Contributions.
We propose Light Online Newton Step (LightONS), a simple variant of ONS that substantially reduces the runtime while retaining the optimal regret. Our method preserves the elegant structure of ONS and inherits its applicability to various scenarios. Our contributions include:
  • An optimal and efficient algorithm for OXO. As summarized in Table˜1, LightONS attains the best-known total runtime $O(d^{2}T+d^{\omega}\sqrt{T\log T})$ to achieve the minimax-optimal regret $O(d\log T)$ for OXO. In terms of regret, LightONS matches ONS’s dependence on all problem parameters in OXO ($T$, $d$, $D$, $G$, $\alpha$), whereas the prior method [Mhammedi and Gatmiry, 2023] suffers from large multiplicative constants. Empirical validations in Appendix˜A corroborate the theoretical superiority of our method.
  • An optimal and efficient algorithm for SXO. Equipped with the online-to-batch conversion, LightONS yields (up to poly-logarithmic factors) the optimal sample complexity $T=\widetilde{O}(d/\epsilon)$ for an excess risk of $\epsilon$, thus reducing the total runtime to $\widetilde{O}(d^{3}/\epsilon)$. Our result answers a COLT’13 open problem of Koren [2013]. We further provide evidence that the runtime $\widetilde{O}(d^{3}/\epsilon)$ is unlikely to be improved in practice.
  • Applicability across various scenarios. Importantly, LightONS preserves the online mirror descent (OMD) framework of ONS and inherits ONS’s structural flexibility, especially in accommodating diverse local norms. In Section˜5, we demonstrate LightONS smoothly replaces ONS and reduces runtime in various scenarios, including gradient-norm adaptivity, parametric stochastic bandits, and memory-efficient OXO. In contrast, the prior method [Mhammedi and Gatmiry, 2023], tailored for OXO, lacks the flexibility to fit for these scenarios.
Outline.
The remainder of this paper is organized as follows. Section˜2 presents preliminaries. Section˜3 presents our method and key analytical ingredients. Section˜4 discusses the implications of LightONS to SXO, which answers a COLT’13 open problem. Section˜5 demonstrates the broad applicability of LightONS inherited from ONS. Section˜6 concludes the paper and discusses future directions. Omitted technical details are deferred to the appendices.
2 Preliminaries
In this section, we introduce notations used throughout this paper, formalize the problem setting of Online eXp-concave Optimization (OXO), and review key relevant prior work.
Notations.
Let $a]_{+}=\max\{0,a\}$, $N]=\{1,\dots,N\}$, and $\mathcal{B}(R)=\{\mathbf{x}\,|\,\left\|\mathbf{x}\right\|_{2}\leq R\}$. For a positive-definite and symmetric matrix $M$, let $\lambda_{i}(M)$ be its $i$-th greatest eigenvalue and let $\left\|\mathbf{x}\right\|_{M}=\sqrt{\mathbf{x}^{\mathsf{T}}M\mathbf{x}}$ be the corresponding Mahalanobis norm. Let $\Pi_{\mathcal{X}}^{M}[\mathbf{y}]=\operatorname*{arg\,min}_{\mathbf{x}\in\mathcal{X}}\left\|\mathbf{x}-\mathbf{y}\right\|_{M}^{2}$ be the Mahalanobis projection, which exists and is unique [Boyd and Vandenberghe, 2004], and let $\Pi_{\mathcal{X}}[\mathbf{y}]=\Pi^{I}_{\mathcal{X}}[\mathbf{y}]=\operatorname*{arg\,min}_{\mathbf{x}\in\mathcal{X}}\left\|\mathbf{x}-\mathbf{y}\right\|_{2}^{2}$ be the Euclidean projection. For shorthand, we write $\texttt{EP}_{\mathcal{X}}$ and $\texttt{MP}_{\mathcal{X}}$ for the runtime of Euclidean and Mahalanobis projection onto $\mathcal{X}$, respectively.
2.1 Problem Setting
Online Convex Optimization (OCO) unfolds as a game between a learner and an environment over $T$ rounds. At each round $t\in[T]$, the learner selects a decision $\mathbf{x}_{t}$ from a compact convex domain $\mathcal{X}\subseteq\mathbb{R}^{d}$, and the environment simultaneously reveals a convex loss function $f_{t}:\mathcal{X}\to\mathbb{R}$; Then the learner incurs a loss $f_{t}(\mathbf{x}_{t})$ and observes a gradient $\nabla f_{t}(\mathbf{x}_{t})$ for updates. The performance of the learner is measured by its regret against some comparator $\mathbf{u}\in\mathcal{X}$, which is defined as:
$\textsc{Reg}_{T}(\mathbf{u})=\sum_{t=1}^{T}f_{t}(\mathbf{x}_{t})-\sum_{t=1}^{T}f_{t}(\mathbf{u}).$
The definition of exp-concavity [Kivinen and Warmuth, 1999; Cesa-Bianchi and Lugosi, 2006] and standard regularity assumptions of OXO [Hazan, 2016; Hazan et al., 2007; Mhammedi and Gatmiry, 2023] are formally stated below.
Assumption 1 (bounded domain).
The domain $\mathcal{X}\subseteq\mathbb{R}^{d}$ is compact and convex, and has a diameter of $D$, i.e., $\max_{(\mathbf{x},\mathbf{y})\in\mathcal{X}^{2}}\left\|\mathbf{x}-\mathbf{y}\right\|_{2}\leq D$. Without loss of generality, let the domain $\mathcal{X}$ be centered at the origin, i.e., $\max_{\mathbf{x}\in\mathcal{X}}\left\|\mathbf{x}\right\|_{2}\leq D/2$.
Assumption 2 (bounded gradient).
For any $t\in[T]$, the loss function $f_{t}:\mathcal{X}\to\mathbb{R}$ is differentiable and $G$-Lipschitz on $\mathcal{X}$, i.e., $\max_{\mathbf{x}\in\mathcal{X}}\left\|\nabla f_{t}(\mathbf{x})\right\|_{2}\leq G$.
Assumption 3 (exp-concave loss).
For any $t\in[T]$, the loss function $f_{t}:\mathcal{X}\to\mathbb{R}$ is $\alpha$-exp-concave on $\mathcal{X}$, i.e., $\exp\left(-\alpha f_{t}(\cdot)\right)$ is concave on $\mathcal{X}$.
2.2 Important Progress
In this subsection, we review two important OXO algorithms, namely Online Newton Step (ONS) [Hazan et al., 2007] and Online Quasi-Newton Step (OQNS) [Mhammedi and Gatmiry, 2023].
Online Newton Step (ONS).
Algorithm˜1 summarizes the update of ONS and enjoys the following theoretical guarantees on its regret and runtime.
Proposition 1 (Theorem 2 of Hazan et al. [2007]).
Under Assumptions˜1, 2 and 3, ONS (Algorithm˜1) with $\gamma_{0}=\frac{1}{2}\min\{\frac{1}{DG},\alpha\}$ satisfies that, for any $\mathbf{u}\in\mathcal{X}$,
$\textsc{Reg}_{T}(\mathbf{u})$ $\leq\frac{d}{2\gamma_{0}}\log\left(1+\frac{G^{2}}{d\epsilon}T\right)+\frac{\gamma_{0}\epsilon D^{2}}{8},$ (1a)
Runtime $\leq O\left(\left(\texttt{MP}_{\mathcal{X}}+d^{2}\right)T\right).$ (1b)
ONS follows the classical framework of online mirror descent (OMD) [Orabona, 2019], equipped with a time-varying Mahalanobis norm induced by the Hessian-related matrix $A_{t}$. The computational bottleneck of ONS lies in Line 6 in Algorithm˜1, the Mahalanobis projection $\mathbf{x}_{t+1}=\Pi^{A_{t}}_{\mathcal{X}}[\widehat{\mathbf{x}}_{t+1}]$. Once the decision $\widehat{\mathbf{x}}_{t+1}$ exits the domain $\mathcal{X}$, it must be projected back. Since each Mahalanobis projection is implemented with $\Omega(d^{\omega})$ arithmetic operations and ONS projects in $O(T)$ rounds in the worst case, the crippling total runtime $\widetilde{O}(d^{\omega}T)$ emerges, which evaluates to $\widetilde{O}(d^{3}T)$ in practice.
Hazan et al. [2007] introduce the following lemma to characterize the curvature induced by exp-concavity, which is crucial to ONS and other practical OXO algorithms. This lemma is essentially Lemma 3 of Hazan et al. [2007], with improved constants. Its proof is deferred to Section˜B.1.
Lemma 1.
If a function $f:\mathcal{X}\to\mathbb{R}$ is $\alpha$-exp-concave and differentiable, then for any $(\mathbf{x},\mathbf{u})\in\mathcal{X}^{2}$, $D\geq\left\|\mathbf{x}-\mathbf{u}\right\|_{2}$, $G\geq\left\|\nabla f(\mathbf{x})\right\|_{2}$, $\gamma\leq\gamma_{0}=\frac{1}{2}\min\{\frac{1}{DG},\alpha\}$, it holds that
$f(\mathbf{x})-f(\mathbf{u})\leq\nabla f(\mathbf{x})^{\mathsf{T}}(\mathbf{x}-\mathbf{u})-\frac{\gamma}{2}\left(\nabla f(\mathbf{x})^{\mathsf{T}}(\mathbf{x}-\mathbf{u})\right)^{2}.$ (2)
We remark that the preceding lemma necessitates the bounded domain assumption (Assumption˜1) to achieve the optimal regret $O(d\log T)$. In particular, a finite diameter $D$ is always required to obtain a non-zero curvature parameter $\gamma$, which reflects the curvature induced by exp-concavity.
Online Quasi-Newton Step (OQNS).
OQNS [Mhammedi and Gatmiry, 2023] effectively reduces the runtime to reach the optimal asymptotical regret $O(d\log T)$, as the following proposition states.
Proposition 2 (Theorem 9 of Mhammedi and Gatmiry [2023]).
Under Assumptions˜1, 2 and 3, OQNS (Algorithm 3 of Mhammedi and Gatmiry [2023]) with $\gamma=\frac{1}{2}\min\{\frac{1}{2DG},\alpha\}$ satisfies that, for any $\mathbf{u}\in\mathcal{X}$,
$\textsc{Reg}_{T}(\mathbf{u})$ $\leq\frac{5d}{\gamma}\log\left(d+T\right)+\frac{11DGd}{2}\log T+3DGd,$ (3a)
Runtime $\leq O\left(\left(\texttt{EP}_{\mathcal{X}}+d^{2}\log T\right)T+d^{\omega}\sqrt{T\log T}\right).$ (3b)
OQNS eliminates Mahalanobis projections with a log-barrier, such as $\log\frac{1}{1-\left\|\mathbf{x}\right\|_{2}^{2}}$ for $\mathcal{X}=\mathcal{B}(1)$, shifting the computational burden to Hessian-inverse evaluations. Key components of OQNS is illustrated below, with $\mathcal{X}=\mathcal{B}(1)$ and $G=1$ as in [Mhammedi and Gatmiry, 2023].
$\mathbf{x}_{t+1}$ $=\mathbf{x}_{t}-\textrm{Approx}\left(\left(\nabla^{2}\Phi_{t}(\mathbf{x}_{t})\right)^{-1}\nabla\Phi_{t}(\mathbf{x}_{t})\right),\penalty 10000\ \penalty 10000\ \text{where}$ (4)
$\Phi_{t}(\mathbf{x})$ $\triangleq\eta d\log\frac{1}{1-\left\|\mathbf{x}\right\|_{2}^{2}}+\frac{d+\eta}{2}\left\|\mathbf{x}\right\|_{2}^{2}+\sum_{s=1}^{t}\left(\nabla f_{s}(\mathbf{x}_{s})^{\mathsf{T}}\mathbf{x}+\frac{\gamma}{2}\left(\nabla f_{s}(\mathbf{x}_{s})^{\mathsf{T}}(\mathbf{x}-\mathbf{x}_{s})\right)^{2}\right).$
Evaluating the Hessian-inverse still takes $O(d^{\omega})$ time and OQNS mitigates this issue by approximating the Hessian-inverse gradient product with incremental updates. The approximation procedure, $\textrm{Approx}(\cdot)$, returns within $O(d^{2}\log T)$ time under proper conditions. OQNS controls the number of exact Hessian-inverse evaluations to $O(\sqrt{T\log T})$, leading to a total runtime of $O(d^{2}T\log T+d^{\omega}\sqrt{T\log T})$, while achieving the optimal regret $O(d\log T)$.
However, their improvement introduces large constant factors. Specifically, when $\alpha\geq\frac{1}{DG}$, the leading $\log T$ term of ONS carries coefficient $DGd$ in Eq.˜1a, whereas OQNS’s leading coefficient is $\frac{51}{2}DGd$ in Eq.˜3a. Moreover, OQNS departs from the OMD framework of ONS, limiting its adaptability to broader scenarios, which will be discussed in Section˜5.
Algorithm 1 ONS [Hazan et al., 2007]
0: preconditioner coefficient $\epsilon$.
1: Initialize $\gamma_{0}=\frac{1}{2}\min\{\frac{1}{DG},\alpha\}$, $A_{0}=\epsilon I$, $\mathbf{x}_{1}=\bm{0}$.
2:for $t=1,\dots,T$ do
3:  Observe $\nabla f_{t}(\mathbf{x}_{t})$.
4:  $A_{t}=A_{t-1}+\nabla f_{t}(\mathbf{x}_{t})\nabla f_{t}(\mathbf{x}_{t})^{\mathsf{T}}$.
5:  $\widehat{\mathbf{x}}_{t+1}=\mathbf{x}_{t}-\frac{1}{\gamma_{0}}A_{t}^{-1}\nabla f_{t}(\mathbf{x}_{t})$.
6:  $\mathbf{x}_{t+1}=\Pi^{A_{t}}_{\mathcal{X}}[\widehat{\mathbf{x}}_{t+1}]$.
7:end for
Algorithm 2 LightONS.Core
0: preconditioner coef. $\epsilon$, deferral coef. $k$.
1: Initialize $\gamma=\frac{1}{2}\min\{\frac{2}{(k+1)DG},\alpha\}$, $A_{0}=\epsilon I$, $\mathbf{x}_{1}=\bm{0}$.
2:for $t=1,\dots,T$ do
3:  Observe $\nabla f_{t}(\mathbf{x}_{t})$.
4:  $A_{t}=A_{t-1}+\nabla f_{t}(\mathbf{x}_{t})\nabla f_{t}(\mathbf{x}_{t})^{\mathsf{T}}$.
5:  $\widehat{\mathbf{x}}_{t+1}=\mathbf{x}_{t}-\frac{1}{\gamma}A_{t}^{-1}\nabla f_{t}(\mathbf{x}_{t})$.
6:  $\mathbf{x}_{t+1}=\begin{cases}\widehat{\mathbf{x}}_{t+1}&\text{if}\penalty 10000\ \left\|\widehat{\mathbf{x}}_{t+1}\right\|_{2}\leq kD/2\\ \Pi^{A_{t}}_{\mathcal{X}}[\widehat{\mathbf{x}}_{t+1}]&\text{otherwise}\end{cases}$.
7:end for
3 Our Algorithm: LightONS
Prior progress naturally raises a question: Can we retain the simplicity and elegance of ONS, while achieving a total runtime that is competitive with, or even superior to, the state-of-the-art? Our work is motivated by answering this question in the affirmative.
In this spirit, we present our algorithm LightONS. In Section˜3.1, to illustrate the key idea, we introduce the core algorithm that only differs from ONS by one line of code. However, the core algorithm is essentially improper learning. To address this issue, we introduce the improper-to-proper conversion in Section˜3.2, which yields the complete version of LightONS.
Furthermore, we empirically validate the superiority of our method, which corroborates theoretical guarantees. Details of the experiments are deferred to Appendix˜A.
3.1 Amortizing Deferred Projections
We first introduce the LightONS.Core in Algorithm˜2, which amortizes the costly Mahalanobis projections with a deferred-projection mechanism. LightONS.Core only differs from ONS by one line of code, as shown in Line 6 of Algorithms˜1 and 2. While ONS projects onto $\mathcal{X}$ immediately when the decision exits $\mathcal{X}$, LightONS.Core continues to update without projection outside $\mathcal{X}$, and projects only when the decision exits an expanded domain $\widetilde{\mathcal{X}}_{k}\subseteq\mathbb{R}^{d}$ defined as
$\widetilde{\mathcal{X}}_{k}\triangleq\mathcal{B}(kD/2)=\{\mathbf{x}\in\mathbb{R}^{d}\,|\,\left\|\mathbf{x}\right\|_{2}\leq kD/2\},$ (5)
where $k>1$ is the deferral coefficient. Essentially, $\widetilde{\mathcal{X}}_{k}$ is obtained by scaling by a factor of $k$ the minimal ball that contains $\mathcal{X}$.
Clearly, a larger $k$ implies lower runtime, as it invokes fewer Mahalanobis projections. The following lemma quantifies the relationship between the deferral coefficient $k$ and the number of Mahalanobis projections $N$. Its proof is deferred to Section˜C.1.
Lemma 2.
Under Assumption˜1, and that for any $t\in[T]$ the loss function $f_{t}$ is $\alpha$-exp-concave, differentiable and $G$-Lipschitz on $\widetilde{\mathcal{X}}_{k}$, let $N$ denote the number of Mahalanobis projections in LightONS.Core (Algorithm˜2) over $T$ rounds, then
$N\leq\left\lfloor\frac{2}{(k-1)D\gamma}\sqrt{\frac{d}{\epsilon}T}\right\rfloor.$ (6)
On the other hand, a larger $k$ exacerbates the deviation of LightONS.Core from ONS, potentially harming the regret guarantee. In particular, increasing $k$ degrades the curvature parameter $\gamma$, jeopardizing the curvature benefit of exp-concavity. In the extreme case where $k$ approaches infinity, $\gamma$ collapses to zero, and exp-concavity degenerates to mere convexity. The following theorem establishes the theoretical guarantees of LightONS.Core, revealing the trade-off between regret and runtime induced by the deferral coefficient $k$. Its proof is deferred to Section˜C.2.
Theorem 1.
Under Assumption˜1, and that for any $t\in[T]$ the loss function $f_{t}$ is $\alpha$-exp-concave, differentiable and $G$-Lipschitz on $\widetilde{\mathcal{X}}_{k}$, with $\gamma=\frac{1}{2}\min\{\frac{2}{(k+1)DG},\alpha\}$, LightONS.Core (Algorithm˜2) satisfies that, for any $\mathbf{u}\in\mathcal{X}$,
$\textsc{Reg}_{T}(\mathbf{u})$ $\leq\frac{d}{2\gamma}\log\left(1+\frac{G^{2}}{d\epsilon}T\right)+\frac{\gamma\epsilon D^{2}}{8},$ (7a)
Runtime $\leq O\left(d^{2}T+(k-1)^{-1}\sqrt{dT/\epsilon}\cdot\texttt{MP}_{\mathcal{X}}\right).$ (7b)
As $k$ increases, Eq.˜7a shows that the regret bound scales as $O(\frac{1}{\gamma})=O(k+1)$, whereas Eq.˜7b indicates that the total runtime spent on Mahalanobis projections decreases as $O(N)=O(\frac{1}{k-1})$. With $k=2$, LightONS.Core already achieves a significant runtime improvement over ONS without sacrificing the optimal regret. Specifically, the regret bound grows by at most a factor of $\frac{3}{2}$, while the number of projections is substantially reduced from $O(T)$ to $O(\sqrt{T})$.
Improper learning issue.
However, LightONS.Core falls in the scope of improper learning [Shalev-Shwartz and Ben-David, 2014], as the algorithm’s decisions $\mathbf{x}_{t}\in\widetilde{\mathcal{X}}_{k}\supset\mathcal{X}$ may reside beyond the domain while the comparator $\mathbf{u}\in\mathcal{X}$ is strictly constrained to the domain. Moreover, LightONS.Core requires additional assumptions that the Lipschitzness and exp-concavity of the loss functions extend to the expanded domain $\widetilde{\mathcal{X}}_{k}$. Vitally, improper learning suppresses the theoretical performance limits of its proper counterparts. A notable illustration is online logistic regression, where an improper learner achieves the regret bound of $O(d\log(GT))$ [Foster et al., 2018], while proper learners are limited to a regret lower bound of $\Omega(d\mathrm{e}^{G}\log T)$ [Hazan et al., 2014].
3.2 The Improper-to-Proper Conversion
Based on techniques from parameter-free online learning [Cutkosky and Orabona, 2018; Cutkosky, 2020], we incorporate a domain conversion into LightONS.Core, yielding LightONS, a fully proper algorithm. This incorporation is conceptually straightforward yet technically subtle, as it inherits the favorable regret-runtime trade-off of LightONS.Core while ensuring proper learning over the domain $\mathcal{X}$. LightONS is summarized in Algorithm˜3. A detailed discussion of the practical implementation of LightONS is deferred to Section˜B.3.
Algorithm 3 LightONS
0: preconditioner coefficient $\epsilon$, deferral coefficient $k$.
1: Initialize $\gamma^{\prime}=\frac{1}{2}\min\{\frac{1}{c_{f}c_{g}DG},\frac{4}{c_{f}c_{g}(k+1)DG},\alpha\}$, $A_{0}=\epsilon I$, $\mathbf{x}_{1}=\mathbf{y}_{1}=\bm{0}$.
2:for $t=1,\dots,T$ do
3:  Observe $\nabla f_{t}(\mathbf{x}_{t})$; and construct $\nabla g_{t}(\mathbf{y}_{t})$ satisfying Condition˜1.
4:  $A_{t}=A_{t-1}+\nabla g_{t}(\mathbf{y}_{t})\nabla g_{t}(\mathbf{y}_{t})^{\mathsf{T}}$.
5:  $\widehat{\mathbf{y}}_{t+1}=\mathbf{y}_{t}-\frac{1}{\gamma^{\prime}}A_{t}^{-1}\nabla g_{t}(\mathbf{y}_{t})$.
6:  $\mathbf{y}_{t+1}=\begin{cases}\widehat{\mathbf{y}}_{t+1}&\text{if}\penalty 10000\ \left\|\widehat{\mathbf{y}}_{t+1}\right\|_{2}\leq kD/2\\ \Pi^{A_{t}}_{\mathcal{B}(D/2)}[\widehat{\mathbf{y}}_{t+1}]&\text{otherwise}\end{cases}$.
7:  $\mathbf{x}_{t+1}=\Pi_{\mathcal{X}}[\mathbf{y}_{t+1}]$.
8:end for
The domain conversion proceeds by constructing surrogate loss functions and projecting surrogate decisions. At each round, LightONS constructs a surrogate loss $g_{t}:\mathbb{R}^{d}\to\mathbb{R}$ from the true loss $f_{t}:\mathcal{X}\to\mathbb{R}$ and supplies $g_{t}$ to to an underlying procedure of LightONS.Core; The underlying LightONS.Core outputs a surrogate decision $\mathbf{y}_{t}\in\widetilde{\mathcal{X}}_{k}$, which LightONS then maps to a (proper) true decision $\mathbf{x}_{t}\in\mathcal{X}$ via a Euclidean projection. For LightONS, any surrogate loss satisfying the condition below suffices to guarantee a valid conversion.
Condition 1.
For some $c_{f}\geq 1$ and $c_{g}\geq 1$, the surrogate loss function $g_{t}:\mathbb{R}^{d}\to\mathbb{R}$ satisfies that, for any $\mathbf{u}\in\mathcal{X}$, $\left\|\nabla g_{t}(\mathbf{y}_{t})\right\|_{2}\leq c_{g}\left\|\nabla f_{t}(\mathbf{x}_{t})\right\|_{2}$ and $\nabla f_{t}(\mathbf{x}_{t})^{\mathsf{T}}(\mathbf{x}_{t}-\mathbf{u})\leq c_{f}\nabla g_{t}(\mathbf{y}_{t})^{\mathsf{T}}(\mathbf{y}_{t}-\mathbf{u})$.
We note that two such conversions have been proposed in the literature. The first, by Cutkosky and Orabona [2018], achieves $c_{f}=2$ and $c_{g}=1$. Later, an improved conversion by Cutkosky [2020] achieves $c_{f}=c_{g}=1$. We prefer the latter which yields smaller constants and tighter regret.
Lemma 3 (Theorem 2 of Cutkosky [2020]).
Under Assumptions˜1 and 2, let $\mathbf{x}_{t}=\Pi_{\mathcal{X}}[\mathbf{y}_{t}]$, the surrogate loss function $g_{t}:\mathbb{R}^{d}\to\mathbb{R}$ and its subgradient at $\mathbf{y}_{t}$ are defined as follows:
$g_{t}(\mathbf{y})$ $\triangleq\nabla f_{t}(\mathbf{x}_{t})^{\mathsf{T}}\mathbf{y}+\frac{[-\nabla f_{t}(\mathbf{x}_{t})^{\mathsf{T}}(\mathbf{y}_{t}-\mathbf{x}_{t})]_{+}}{\left\|\mathbf{y}_{t}-\mathbf{x}_{t}\right\|_{2}}\left\|\mathbf{y}-\Pi_{\mathcal{X}}[\mathbf{y}]\right\|_{2},$ (8a)
$\nabla g_{t}(\mathbf{y}_{t})$ $=\nabla f_{t}(\mathbf{x}_{t})+\frac{[-\nabla f_{t}(\mathbf{x}_{t})^{\mathsf{T}}(\mathbf{y}_{t}-\mathbf{x}_{t})]_{+}}{\left\|\mathbf{y}_{t}-\mathbf{x}_{t}\right\|_{2}^{2}}\left(\mathbf{y}_{t}-\mathbf{x}_{t}\right),$ (8b)
then for any $\mathbf{u}\in\mathcal{X}$, $\left\|\nabla g_{t}(\mathbf{y}_{t})\right\|_{2}\leq\left\|\nabla f_{t}(\mathbf{x}_{t})\right\|_{2}$, and $\nabla f_{t}(\mathbf{x}_{t})^{\mathsf{T}}(\mathbf{x}_{t}-\mathbf{u})\leq\nabla g_{t}(\mathbf{y}_{t})^{\mathsf{T}}(\mathbf{y}_{t}-\mathbf{u})$.
The computational overhead of this conversion is negligible. Constructing the surrogate gradient $\nabla g_{t}(\mathbf{y}_{t})$ as Eq.˜8b takes only $O(d)$ arithmetic operations per round apart from the Euclidean projection, veiled by the $O(d^{2})$ cost of updating and inverting the Hessian-related matrix $A_{t}$. It is worth noting that prior work has also employed this domain-conversion technique to address projection-related issues, but for different purposes, such as in non-stationary online learning [Zhao et al., 2025] and universal online learning [Yang et al., 2024].
One caveat is that the surrogate loss $g_{t}$ in Lemma˜3 is actually not exp-concave, seemingly precluding the curvature benefit. Fortunately, the following lemma shows that the surrogate loss $g_{t}$ inherits the curvature in a form closely mirroring Lemma˜1, with proof in Section˜C.3.
Lemma 4.
Under Assumptions˜1, 2 and 3, with $\gamma^{\prime}=\frac{1}{2}\min\{\frac{1}{c_{f}c_{g}DG},\frac{4}{(k+1)c_{f}c_{g}DG},\alpha\}$, for any $\mathbf{u}\in\mathcal{X}$ and any surrogate loss function $g_{t}$ satisfying Condition˜1,
$f_{t}(\mathbf{x}_{t})-f_{t}(\mathbf{u})$ $\leq\nabla f_{t}(\mathbf{x}_{t})^{\mathsf{T}}(\mathbf{x}_{t}-\mathbf{u})-\frac{\gamma_{0}}{2}\left(\nabla f_{t}(\mathbf{x}_{t})^{\mathsf{T}}(\mathbf{x}_{t}-\mathbf{u})\right)^{2}$ (9)
$\leq\nabla g_{t}(\mathbf{y}_{t})^{\mathsf{T}}(\mathbf{y}_{t}-\mathbf{u})-\frac{\gamma^{\prime}}{2}\left(\nabla g_{t}(\mathbf{y}_{t})^{\mathsf{T}}(\mathbf{y}_{t}-\mathbf{u})\right)^{2}.$
We remark that, when the deferral coefficient $k\leq 3$ and $c_{f}=c_{g}=1$ (e.g., in Lemma˜3), the surrogate curvature parameter $\gamma^{\prime}$ in Lemma˜4 is unimpaired relative to the original $\gamma_{0}$ in Lemma˜1.
The following theorem establishes the theoretical guarantees of LightONS, and an immediate corollary specifies its regret and runtime by instantiating deferral coefficient $k=2$ and the surrogate loss in Lemma˜3. Its proof is deferred to Section˜C.4.
Theorem 2.
Under Assumptions˜1, 2 and 3, with $\gamma^{\prime}=\frac{1}{2}\min\{\frac{1}{c_{f}c_{g}DG},\frac{4}{c_{f}c_{g}(k+1)DG},\alpha\}$, LightONS (Algorithm˜3) satisfies that, for any $\mathbf{u}\in\mathcal{X}$,
$\textsc{Reg}_{T}(\mathbf{u})$ $\leq\frac{d}{2\gamma^{\prime}}\log\left(1+\frac{c_{g}^{2}G^{2}}{d\epsilon}T\right)+\frac{\gamma^{\prime}\epsilon D^{2}}{8},$ (10a)
Runtime $\leq O\left(\left(\texttt{EP}_{\mathcal{X}}+d^{2}\right)T+(k-1)^{-1}d^{\omega+0.5}\sqrt{T/\epsilon}\log T\right).$ (10b)
Corollary 1.
Under Assumptions˜1, 2 and 3, by setting the surrogate loss function as Lemma˜3, $k=2$ and $\epsilon=d\log T$, with $\gamma_{0}=\frac{1}{2}\min\{\frac{1}{DG},\alpha\}$, LightONS (Algorithm˜3) satisfies that, for any $\mathbf{u}\in\mathcal{X}$,
$\textsc{Reg}_{T}(\mathbf{u})$ $\leq\frac{d}{2\gamma_{0}}\log\left(1+\frac{G^{2}}{d\epsilon}T\right)+\frac{\gamma_{0}\epsilon D^{2}}{8},$ (11a)
Runtime $\leq O\left(\left(\texttt{EP}_{\mathcal{X}}+d^{2}\right)T+d^{\omega}\sqrt{T\log T}\right).$ (11b)
Remark 1.
LightONSachieves computational efficiency at the cost of a slightly worse regret than ONS, as seen in the gap of the second inequality in Eq.˜9. Nonetheless, this degradation is dominated by the problem-dependent parameters of OXO, namely $T$, $d$, $D$, $G$, $\alpha$. In their dependence on these parameters, LightONS’s regret in Eq.˜11a exactly matches ONS’s regret in Eq.˜1a.
LightONSachieves a total runtime better than both ONS and OQNS while matching the regret of ONS. More importantly, LightONS retains the flexible OMD framework of ONS. In fact, LightONS differs from ONS in two key aspects: (i) the deferred-projection mechanism (introduced in LightONS.Core to boost efficiency) and (ii) the domain conversion (introduced here to ensure proper learning), both of which are largely orthogonal to the mirror-descent update in the standard ONS. Consequently, LightONS applies to various scenarios where ONS is essential, particularly when its mirror-descent update plays a critical role. We illustrate these applications in Section˜5.
4 Answering a COLT’13 Open Problem
Via the online-to-batch conversion, our method applies to Stochastic eXp-concave Optimization (SXO), where the optimal OXO regret translates into the (near) optimal SXO sample complexity while substantially reducing the computational cost. This extension answers a COLT’13 open problem posed by Koren [2013], demonstrating our method’s significance beyond OXO.
4.1 Restatement of the Open Problem
SXO seeks to minimize an exp-concave function $F:\mathcal{X}\to\mathbb{R}$ over a convex domain $\mathcal{X}\subseteq\mathbb{R}^{d}$, and the learner has access to $F$ only through some stochastic oracle. In this section, we consider the case where $F$ is the expectation of a random function $f:\mathcal{X}\times\Xi\to\mathbb{R}$ over a distribution on $\Xi$, and the learner has access to the gradient of $f(\cdot;\xi)$ with $\xi$ drawn from $\mathcal{D}_{\xi}$. We are interested in the sample complexity and total runtime required to find an $\epsilon$-optimal solution $\mathbf{x}_{\epsilon}$, i.e.,
$F(\mathbf{x}_{\epsilon})-\min_{\mathbf{x}\in\mathcal{X}}F(\mathbf{x})\leq\epsilon,\quad\text{where}\penalty 10000\ F(\mathbf{x})=\mathop{\mathbb{E}}_{\xi\sim\mathcal{D}_{\xi}}\left[f(\mathbf{x};\xi)\right].$
The regularity of the random function is formally stated below.
Assumption 4.
The loss function $F:\mathcal{X}\to\mathbb{R}$ is the expectation of a random function $f:\mathcal{X}\times\Xi\to\mathbb{R}$ over an unknown distribution $\mathcal{D}_{\xi}$ on $\Xi$, i.e., $F(\mathbf{x})=\mathop{\mathbb{E}}_{\xi\sim\mathcal{D}_{\xi}}\left[f(\mathbf{x};\xi)\right]$. For any $\xi\in\Xi$, the random loss function $f(\cdot;\xi)$ is $\alpha$-exp-concave, differentiable and $G$-Lipschitz over $\mathcal{X}$. For any query point $\mathbf{x}\in\mathcal{X}$, the stochastic gradient oracle returns $\nabla f(\mathbf{x};\xi)$ with $\xi$ i.i.d. drawn from $\mathcal{D}_{\xi}$.
[…]
Open problem [Koren, 2013].
Under Assumption˜4 and that $\mathcal{X}=\mathcal{B}(1)$, (a) Is it possible to find an SXO algorithm that attains the sample complexity of $\widetilde{O}(d/\epsilon)$ with only linear-in-$d$ runtime per iteration, i.e., $\widetilde{O}(d^{2}/\epsilon)$ runtime overall? (b) Is it possible to perform any better than $\widetilde{O}(d^{4}/\epsilon)$ runtime overall?
The first part of the open problem remains open. In particular, Mahdavi et al. [2015] establishes an information-theoretic sample complexity lower bound of $\Omega(d/\epsilon)$ for SXO, which implies a runtime lower bound of $\Omega(d^{2}/\epsilon)$ since each gradient query costs $\Omega(d)$ time.
Section˜4.2 answers the second part in the affirmative by combining LightONS with the online-to-batch conversion, obtaining total runtime $\widetilde{O}(d^{3}/\epsilon)$ while maintaining the optimal $\widetilde{O}(d/\epsilon)$ sample complexity. We further present evidence that this $\widetilde{O}(d^{3}/\epsilon)$ runtime is likely unimprovable. If confirmed, it would refute the first part and settle the full open problem.
4.2 Answering the Open Problem with LightONS
Built upon the online-to-batch conversion of Mehta [2017], the LightONS-based SXO method reduces the total runtime to $\widetilde{O}(d^{3}/\epsilon)$, with (near) optimal convergence rate both in high probability and in expectation, as the following theorem shows. Its proof is deferred to Section˜D.1.
Theorem 3.
Under Assumption˜4 and that $\mathcal{X}=\mathcal{B}(1)$, let LightONS (Algorithm˜3) run for $T$ rounds with gradients $\{\nabla f(\mathbf{x}_{t};\xi_{t})\}_{t=1}^{T}$ where $\{\mathbf{x}_{t}\}_{t=1}^{T}$ are decisions, let $\bar{\mathbf{x}}_{T}=\frac{1}{T}\sum_{t=1}^{T}\mathbf{x}_{t}$, then
  • For any $\delta\in(0,1)$, with probability at least $1-\delta$, when $T=\Theta\left(\frac{d}{\epsilon}\log\frac{d}{\epsilon}\log\frac{1}{\delta}\right)$,
    $F(\bar{\mathbf{x}}_{T})-\min_{\mathbf{x}\in\mathcal{X}}F(\mathbf{x})\leq O(\epsilon)$ and $\textsc{Runtime}\leq\widetilde{O}\left(\frac{d^{3}}{\epsilon}\right)$.
  • When $T^{\prime}=\Theta\left(\frac{d}{\epsilon}\log\frac{d}{\epsilon}\right)$, $\mathop{\mathbb{E}}\left[F(\bar{\mathbf{x}}_{T^{\prime}})-\min_{\mathbf{x}\in\mathcal{X}}F(\mathbf{x})\right]\leq O(\epsilon)$ and $\textsc{Runtime}\leq\widetilde{O}\left(\frac{d^{3}}{\epsilon}\right)$.
The LightONS-based SXO method in Theorem˜3 matches the best known SXO runtime up to poly-logarithmic factors. An online-to-batch conversion of OQNS also yields a total runtime of $\widetilde{O}(d^{3}/\epsilon)$ but carries significant practical overheads due to its large constant factors, as discussed in Section˜3. Our experiments in Appendix˜A, conducted under SXO settings, confirm this difference.
Conjecture on the optimality of $\widetilde{O}(d^{3}/\epsilon)$.
We conjecture that no SXO algorithm can asymptotically improve upon the total runtime of $\widetilde{O}(d^{3}/\epsilon)$. A natural alternative approach to SXO is empirical risk minimization (ERM) [Koren and Levy, 2015; Mehta, 2017], which reduces SXO to offline exp-concave optimization and admits offline acceleration techniques. Notably, even in the simplest least-squares setting with full access to the data, no $\widetilde{O}(d^{2}/\epsilon)$-time algorithm is known. A detailed discussion on offline acceleration techniques, including fast matrix multiplication and cutting-plane methods, is deferred to Section˜D.2. Moreover, ERM-based SXO methods require $\Omega(d^{2}+d/\epsilon)$ memory to store all samples, whereas the OXO-based methods require only $O(d^{2})$ memory.
Table 2: Regret and total runtime comparison between ONS and LightONS on three benchmark applications. “OQNS?” indicates whether OQNS supports the same application.
Application Regret (ONS) Runtime (ONS) Regret (LightONS) Runtime (LightONS) OQNS?
Gradient-norm adaptivity $O(d\log G_{T})$ $\widetilde{O}(d^{\omega}T)$ same $O(d^{2}T+d^{\omega}\sqrt{T}\log T)$ hardly
Generalized linear bandits $\widetilde{O}(d\sqrt{T}+\kappa d^{2})$ $\widetilde{O}(d^{2}KT\,+$ $d^{\omega}T)$ same order $\widetilde{O}(d^{2}KT\,+$ $d^{\omega}\cdot\min\{\sqrt{\kappa dT},T\})$ hardly
Memory-efficient OXO $O(d^{\prime}\log T)$ $\widetilde{O}(d^{\omega}T)$ same $O((d^{2}+d^{\prime}d\log T)T)$ intricate
5 Applications to Various Problems
ONS extends its influence far beyond the classical settings of OXO and SXO. Since LightONS is designed to retain core updates of ONS with minimal modifications, it preserves its elegant structure. This enables seamless integration into a range of applications where ONS serves as the computational core, not only maintaining its statistical advantages but also significantly enhancing efficiency.
We highlight three representative applications of ONS (gradient-norm adaptivity, parameteric stochastic bandits, and memory-efficient OXO) where LightONS seamlessly fits, preserving statistical benefits while requiring minimal additional analytical effort. In contrast, deploying OQNS in these settings is either infeasible or would require substantial and non-trivial modifications. The main results are summarized in Table˜2, comparing ONS with LightONS and OQNS. We defer further details and proofs to Appendix˜E.
5.1 Gradient-Norm Adaptivity
ONS naturally achieves a problem-dependent regret bound that scales with the accumulated squared gradient norms $G_{T}$ instead of the time horizon $T$:
$\textsc{Reg}_{T}(\mathbf{u})=O\left(d\log G_{T}\right),\quad\text{where}\penalty 10000\ G_{T}\triangleq\sum_{t=1}^{T}\left\|\nabla f_{t}(\mathbf{x}_{t})\right\|_{2}^{2},$ (12)
where $G_{T}$ can be $o(T)$ in benign environments and is upper bounded by $G^{2}T$, safeguarding the optimal regret $O(d\log T)$ against worst cases. Prior work has leveraged this gradient-norm adaptivity to achieve strong guarantees in various settings. For OXO with smooth loss functions, Orabona et al. [2012] achieve a small-loss adaptive regret bound $\textsc{Reg}_{T}=O(d\log L_{T})$, where $L_{T}\triangleq\min_{\mathbf{u}\in\mathcal{X}}\sum_{t=1}^{T}(f_{t}(\mathbf{u})-\min_{\mathbf{x}\in\mathcal{X}}f_{t}(\mathbf{x}))$. For OCO with unbounded domain, Cutkosky and Orabona [2018] achieve a comparator-norm adaptive regret bound without prior knowledge of the comparator norm $\left\|\mathbf{u}\right\|_{2}$, i.e., $\textsc{Reg}_{T}(\mathbf{u})=\widetilde{O}(\left\|\mathbf{u}\right\|_{2}\sqrt{dG_{T}})$ for any $\mathbf{u}\in\mathbb{R}^{d}$. Based on ONS, preceding results incur the worst-case runtime bottleneck $\widetilde{O}(d^{\omega}T)$. By incorporating LightONS, the runtime improves to $O(d^{2}T+d^{\omega}\sqrt{T}\log T)$ while preserving the same regret.
OQNS can hardly achieve gradient-norm adaptivity. Its computational efficiency relies on the log-barrier that stabilizes iterates for efficient matrix approximations. However, this log-barrier incurs an unavoidable $O(\log T)$ term which precludes adaptation to $G_{T}$. It is known that the OMD framework permits a flexible trade-off between stability and bias in the regret decomposition. While LightONS retains this flexibility, the log-barrier in OQNS overly suppresses the stability term, leading to a large bias term that does not scale with the gradient norms. Details are deferred to Section˜E.1.
5.2 Parametric Stochastic Bandits
Parametric stochastic bandits model decision-making problems with partial feedback, and generalized linear bandits (GLB) is a fundamental instance with non-linearly parameterized losses. A notable challenge in GLB is its dependence on the condition number $\kappa\propto\exp(D)$. Directly applying ONS to GLB yields an $\widetilde{O}(\kappa d\sqrt{T})$ regret [Jun et al., 2017]. Recently, Zhang et al. [2025] propose the first jointly efficient GLB algorithm, which achieves an $\widetilde{O}(d\sqrt{T}+\kappa d^{2})$ regret (statistically efficient) and is one-pass (computationally efficient). Their key technique is an ONS-based subroutine with non-monotonic local norms, which enjoys the one-pass efficiency of ONS and eliminates $\kappa$ from leading terms in regret. By replacing with a LightONS-based counterpart, we retain the same regret bound and improves the runtime from $\widetilde{O}((d^{2}K+d^{\omega})T)$ to $\widetilde{O}(d^{2}KT+d^{\omega}\cdot\min\{\sqrt{\kappa dT},T\})$, where $K$ is the number of arms for bandits. On condition that the regret is sublinear in $T$, i.e., $\kappa=o(T)$, our improvement yields a strictly lower runtime order.
OQNS can hardly be adapted to the method of Zhang et al. [2025]. The reasons are two-fold: (i) The GLB analysis treats the ONS-based subroutine as a white-box, exploiting negative terms in the OMD analysis, while OQNS fails to provide such analytical properties due to its deviation from OMD; (ii) The GLB method relies on non-monotonic local norms, while OQNS is tailored for monotonic local norms as in Eq.˜4, with which OQNS’s analysis is deeply coupled. On the contrary, the key ingredients of LightONS (deferred projection and domain conversion) are largely orthogonal to the core mirror-descent updates, seamlessly accommodating customized local norms and preserving necessary analytical properties for GLB. Details are deferred to Section˜E.2.
5.3 Memory-Efficient OXO
Another computational challenge of ONS is its $O(d^{2})$ working memory, in contrast to OGD’s $O(d)$ working memory. Luo et al. [2016] propose Sketched Online Newton Step (SON), which achieves linear-in-$d$ runtime and memory with matrix sketching when the gradient feedbacks exhibit low-dimensional structures. However, SON demands additional assumptions on the domain and loss functions and reverts to the $\widetilde{O}(d^{\omega}T)$ runtime under the standard OXO setting (Assumptions˜1, 2 and 3). Equipping LightONS with sketching, we propose a hybrid algorithm which combines the projection efficiency of LightONS with the memory efficiency of SON, extending the linear-in-$d$ runtime and memory to the standard OXO setting.
Although Mhammedi and Gatmiry [2023] suggest integrating sketching into OQNS, their algorithm’s considerable structural complexity may pose significant obstacles to such integration. Details are deferred to Section˜E.3.
6 Conclusion
For online exp-concave optimization (OXO), we propose LightONS, a simple yet powerful variant of ONS. By combining the deferred-projection mechanism with the improper-to-proper conversion, LightONS achieves significant reductions in worst-case total runtime while preserving the regret optimality of ONS. These gains extend to the stochastic optimization setting, answering a COLT’13 open problem on efficient and optimal SXO [Koren, 2013]. Moreover, due to its fidelity to the mirror-descent udpate of ONS, LightONS serves as an efficient drop-in replacement of ONS across diverse applications, including gradient-norm adaptivity, parametric stochastic bandits, and memory-efficient settings, all without compromising statistical guarantees.
Several important directions remain open. First, while LightONS adapts to gradient norms, achieving adaptivity to gradient variation for OXO within $\widetilde{O}(d^{2}T)$ runtime remains an open challenge [Chiang et al., 2012; Zhao et al., 2024]. Second, applying the LightONS technique to other preconditioned online learning algorithms may not improve their (asymptotic) runtime if the bottleneck lies in other operations, such as explicit matrix factorization rather than Mahalanobis projections (e.g., in AdaGrad [Duchi et al., 2011]). Accelerating these algorithms remains unclear. These future directions motivate the development of broader, general-purpose acceleration techniques for Hessian-related online learning algorithms.
Appendix A Empirical Validation
We conduct numerical experiments to validate the theoretical guarantees of LightONS, especially its negligible statistical gap from ONS and its non-asymptotic statistical advantages over OQNS.
Setup.
We evaluate on two fundamental tasks, linear and logistic regression, over the Euclidean ball domain $\mathcal{X}=\mathcal{B}(D/2)\subset\mathbb{R}^{d}$ over a time horizon of $T$. With $\{(\mathbf{x}_{t},y_{t})\}_{t=1}^{T}$ being i.i.d. sampled from the folded standard Gaussian distribution, i.e., each entry of $\mathbf{x}_{t}$ and $y_{t}$ is the absolute value of a standard Gaussian random variable, the (online) loss functions are designed as
$\ell^{\text{linear}}_{t}(\mathbf{w})=\frac{1}{2}\left(\sqrt{\frac{G}{D}}\mathbf{x}_{t}^{\mathsf{T}}\mathbf{w}+\frac{\sqrt{DG}}{2}y_{t}\right)^{2},\quad\ell^{\text{logistic}}_{t}(\mathbf{w})=\log\left(1+\exp\left(G\mathbf{x}_{t}^{\mathsf{T}}\mathbf{w}\right)\right).$
Both loss functions are $G$-Lipschitz and $\alpha$-exp-concave with $\alpha=1/(DG)$ and $\alpha=\exp(-DG)$, respectively. Our logistic regression setup slightly differs from the standard formulation, as we omit the binary label $y_{t}\in\{0,1\}$ for simplicity, which does not affect exp-concavity.
Three algorithms, LightONS, ONS, and OQNS, are configured with their theoretically optimal parameters. The implementation of ONS and LightONS follows Algorithms˜1 and 3, respectively; OQNS is implemented in accordance with Algorithm 3 of Mhammedi and Gatmiry [2023].
Results.
Figure˜1 displays all results averaged over 5 independent runs with the same seeds, where the averaged performance is shown as a dark line, while individual runs are in transparent lines.
Our experiments confirm that computational gains of LightONS incur negligible loss in statistical performance. As shown in Figures˜1(a), 1(d), 1(b) and 1(e), curves of LightONS (red dotted) overlap with those of ONS (green solid). This provides empirical evidence that LightONS retains the sharp regret constants of the ONS. In contrast, OQNS (blue dashed) shows noticeably higher regret, indicating that its theoretical bounds may carry greater constants.
The key advantage of LightONS, its computational efficiency, is evident in Figures˜1(c) and 1(f). After an initial convergence phase (e.g., the first $10^{2}$ rounds in Figure˜1(f)), both ONS and LightONS stabilize near the offline optimal solution. Then ONS frequently pushes decisions outside the domain, triggering costly Mahalanobis projections. In contrast, LightONS consistently outputs decisions with $\left\|\mathbf{y}_{t}\right\|\leq kD/2=2$ and avoids Mahalanobis projections.
Appendix B Technical Lemmas
In this section, we present technical lemmas used in the subsequent analysis.
B.1 Proof of Lemma˜1
Proof.
Since $g(\mathbf{x})=\mathrm{e}^{-\alpha f(\mathbf{x})}$ is concave and positive, we have that $h(\mathbf{x})=\mathrm{e}^{-2\gamma_{0}f(\mathbf{x})}$ is also concave. Specifically, for any $\lambda\in[0,1]$ and any $(\mathbf{x},\mathbf{y})\in\mathcal{X}^{2}$, it holds that
$h(\lambda\mathbf{x}+(1-\lambda)\mathbf{y})$ $=g(\lambda\mathbf{x}+(1-\lambda)\mathbf{y})^{2\gamma_{0}/\alpha}\geq\left(\lambda g(\mathbf{x})+(1-\lambda)g(\mathbf{y})\right)^{2\gamma_{0}/\alpha}$
$\geq\lambda g(\mathbf{x})^{2\gamma_{0}/\alpha}+(1-\lambda)g(\mathbf{y})^{2\gamma_{0}/\alpha}=\lambda h(\mathbf{x})+(1-\lambda)h(\mathbf{y}),$
where the first inequality follows from the concavity of $g$ and the second from the concavity of $a\mapsto a^{2\gamma_{0}/\alpha}$ (where $2\gamma_{0}/\alpha\leq 1$). Next, by the concavity of $h$, for any $(\mathbf{x},\mathbf{u})\in\mathcal{X}^{2}$, we have
$\mathrm{e}^{-2\gamma_{0}f(\mathbf{u})}\leq\mathrm{e}^{-2\gamma_{0}f(\mathbf{x})}+\left(-2\gamma_{0}\mathrm{e}^{-2\gamma_{0}f(\mathbf{x})}\nabla f(\mathbf{x})\right)^{\mathsf{T}}(\mathbf{u}-\mathbf{x}).$
Rearranging the preceding inequality yields
$f(\mathbf{x})-f(\mathbf{u})\leq\frac{1}{2\gamma_{0}}\log\left(1-2\gamma_{0}\nabla f(\mathbf{x})^{\mathsf{T}}(\mathbf{u}-\mathbf{x})\right).$
To prove Eq.˜2, it suffices to apply that $\log(1+a)\leq a-a^{2}/4$ (which holds for any $\left|a\right|\leq 1$) to the right-hand side of the preceding inequality. We note that $\left|2\gamma_{0}\nabla f(\mathbf{x})^{\mathsf{T}}(\mathbf{x}-\mathbf{u})\right|\leq 1$ due to the selection of the curvature parameter $\gamma_{0}=\frac{1}{2}\min\{\frac{1}{DG},\alpha\}$. ∎
B.2 Elliptical Potential Lemmas
We introduce two lemmas tackling Hessian-related matrices, which are essential for both regret and runtime analysis. Lemma˜5 is standard in the ONS literature (e.g., [Hazan et al., 2007, Luo et al., 2016]). The proof of Lemma˜6 mirrors that of Lemma˜5, differing only in the potential function: $F(A)=-\operatorname{tr}(A)$ versus $F(A)=\log\operatorname{det}(A)$.
Lemma 5.
Let $A_{i}=\lambda I+\sum_{j=1}^{i}\mathbf{v}_{j}\mathbf{v}_{j}^{\mathsf{T}}$, if $\left\|\mathbf{v}_{i}\right\|_{2}\leq L$ for any $i\in[n]$, then
$\sum_{i=1}^{n}\mathbf{v}_{i}^{\mathsf{T}}A_{i}^{-1}\mathbf{v}_{i}\leq\log\operatorname{det}(A_{n})-\log\operatorname{det}(A_{0})\leq d\log\left(1+\frac{L^{2}}{d\lambda}n\right).$ (13)
Lemma 6.
Let $A_{i}=\lambda I+\sum_{j=1}^{i}\mathbf{v}_{j}\mathbf{v}_{j}^{\mathsf{T}}$, if $\left\|\mathbf{v}_{i}\right\|_{2}\leq L$ for any $i\in[n]$, then
$\sum_{i=1}^{n}\mathbf{v}_{i}^{\mathsf{T}}A_{i}^{-2}\mathbf{v}_{i}\leq\operatorname{tr}(A_{0}^{-1})-\operatorname{tr}(A_{n}^{-1})\leq\frac{d}{\lambda}.$ (14)
Proof of Lemmas˜5 and 6.
Let $\mathbb{S}_{++}$ denote the set of $d$-dimensional positive-definite and symmetric matrices, and let $\left\langle\cdot,\cdot\right\rangle_{\operatorname{F}}$ denote the inner product induced by the Frobenius matrix norm. Then for any concave function $F:\mathbb{S}_{++}\to\mathbb{R}$, we have
$\sum_{i=1}^{n}\mathbf{v}_{i}^{\mathsf{T}}\nabla{F}(A_{i})\mathbf{v}_{i}=\sum_{i=1}^{n}\left\langle\nabla{F}(A_{i}),A_{i}-A_{i-1}\right\rangle_{\operatorname{F}}\leq\sum_{i=1}^{n}\left(F(A_{i})-F(A_{i-1})\right)=F(A_{n})-F(A_{0}).$ (15)
The first equality follows from $A_{i}-A_{i-1}=\mathbf{v}_{i}\mathbf{v}_{i}^{\mathsf{T}}$ (by the definition of $A_{i}$) and $\mathbf{u}^{\mathsf{T}}A\mathbf{v}=\left\langle A,\mathbf{v}\mathbf{u}^{\mathsf{T}}\right\rangle_{\operatorname{F}}$ (by the definition of the $\left\langle\cdot,\cdot\right\rangle_{\operatorname{F}}$); The second inequality from the concavity of $F$; The last equality from telescoping. Substituting the respective potentials into Eq.˜15 yields Lemmas˜5 and 6. Concretely,
  • For Lemma˜5, $F(A)=\log\operatorname{det}(A)$ and $\nabla{F}(A)=A^{-1}$.
  • For Lemma˜6, $F(A)=-\operatorname{tr}(A^{-1})$ and $\nabla{F}(A)=A^{-2}$.
Concavity of $F$ is standard [Boyd and Vandenberghe, 2004] and gradients of $F$ follow from matrix calculus [Petersen and Petersen, 2012]. ∎
Algorithm 4 FastProj onto Euclidean ball $\mathcal{B}(R)$ with Mahalanobis norm $\left\|\cdot\right\|_{A}$
0: point $\mathbf{u}\notin\mathcal{B}(R)$, error tolerance $\zeta$, range of AA’s eigenvalues $\underline{\lambda},\overline{\lambda}]$.
0: approximate Mahalanobis projection $\mathbf{v}\approx\Pi^{A}_{\mathcal{B}(R)}[\mathbf{u}]$.
1: (Choice 1.) $\mathbf{p}=A\mathbf{u}$, let $\rho(\mu)=\|\left(A+\mu I\right)^{-1}\mathbf{p}\|_{2}^{2}-R^{2}$.
2: (Choice 2.) Tridiagonalize $QCQ^{\mathsf{T}}=A$, $\mathbf{q}=CQ^{\mathsf{T}}\mathbf{u}$, let $\rho(\mu)=\|\left(C+\mu I\right)^{-1}\mathbf{q}\|_{2}^{2}-R^{2}$.
[…]
4:for $i=1,\dots,n$ do
[…]
6:end for
[…]
B.3 Numerical Lemmas
The overall efficiency of LightONS hinges on two key operations, namely the matrix inversion $A_{t}^{-1}$ and the (infrequent) Mahalanobis projection $\Pi^{A_{t}}_{\mathcal{B}(D/2)}[\widehat{\mathbf{y}}_{t+1}]$. In this subsection, we detail efficient numerical approaches for the two operations respectively.
Matrix inversion.
The per-round update of the matrix $A_{t}$ is a rank-one update. Instead inverting from scratch at a cost of $O(d^{\omega})$, we can update $A_{t-1}^{-1}$ to obtain $A_{t}^{-1}$ in only $O(d^{2})$ time with the Sherman-Morrison-Woodbury formula. The following equation ensures $V_{t}=A_{t}^{-1}$ for any $t\in[T]$.
$V_{0}=\frac{1}{\epsilon}I;\quad V_{t+1}=V_{t}-\frac{1}{\left\|\nabla g_{t}(\mathbf{y}_{t})\right\|_{V_{t}^{-1}}^{2}}V_{t}\nabla g_{t}(\mathbf{y}_{t})\nabla g_{t}(\mathbf{y}_{t})^{\mathsf{T}}V_{t}.$ (16)
Mahalanobis projection onto Euclidean ball.
In LightONS, all Mahalanobis projections are onto the Euclidean ball $\widetilde{\mathcal{X}}_{k}=\mathcal{B}(kD/2)$ rather than the potentially complex domain $\mathcal{X}$. This geometry enables customized numerical approaches faster than generic solvers [Lee et al., 2015, Jiang et al., 2020]. Specifically, the dual problem of the Mahalanobis projection reduces to a one-dimensional root-finding problem (Exercise 4.22 of Boyd and Vandenberghe [2004]), solvable via bisection.
We propose FastProj in Algorithm˜4, which offers two implementation choices for different purposes. Choice 1 attains superior theoretical dependence on $d$ by exploiting fast matrix multiplication with exponent $\omega<2.3714$ [Alman et al., 2025]. Choice 2 offers stronger practical performance via tridiagonalization tricks [Parlett, 1998, Golub and Van Loan, 2013], which is more favorable in practice where $\omega=3$. The following lemma characterizes the error and runtime of FastProj.
Lemma 7.
Let $\mathbf{v}$ be the output of FastProj (Algorithm˜4), and $\mathbf{v}^{*}=\Pi^{A}_{\mathcal{B}(R)}[\mathbf{u}]$ be the exact Mahalanobis projection, then $\mathbf{v}\in\mathcal{B}(R)$ and $\left\|\mathbf{v}-\mathbf{v}^{*}\right\|_{2}\leq\zeta$. The number of bisection $n=O(\log(\frac{1}{\zeta}\frac{\penalty 10000\ \overline{\lambda}\penalty 10000\ }{\penalty 10000\ \underline{\lambda}\penalty 10000\ }))$. With choice 1, the runtime is $O(d^{\omega}n)$, and with choice 2, the runtime is $O(d^{3}+dn)$.
In the context of LightONS (Algorithm˜3), it suffices to set the error tolerance as $\zeta_{t}=O(1/t^{2})$ at the $t$-th round. The accumulated truncation error contributes only a negligible additive constant to the regret of LightONS. A detailed discussion is deferred to Lemma˜12 in Section˜C.4. Moreover, $\underline{\lambda}_{t}=\epsilon$ and $\overline{\lambda}_{t}=\epsilon+c_{g}^{2}G^{2}t$, we have $n_{t}=O(\log t)$ at the $t$-th round.
To prove Lemma˜7, we first present two supporting lemmas. Lemma˜8 reduces the Mahalanobis projection onto an Euclidean ball to a one-dimensional root-finding problem. Lemma˜9 bridges the truncation error of root-finding and the truncation error of Mahalanobis projection.
Lemma 8.
$\Pi^{A}_{\mathcal{B}(R)}[\mathbf{y}]=(A+\mu^{\ast}I)^{-1}A\mathbf{y}$, where $\mu^{\ast}$ is the only positive zero of the following function
$\rho(\mu)=\left\|\left(A+\mu I\right)^{-1}A\mathbf{y}\right\|_{2}^{2}-R^{2}=\sum_{i=1}^{d}\frac{v_{i}^{2}}{\left(1+\frac{\mu}{\lambda_{i}}\right)^{2}}-R^{2},$ (17)
and $\mu^{\ast}$ satisfies that
$\left(\frac{\left\|\mathbf{y}\right\|_{2}}{R}-1\right)\lambda_{d}<\mu^{\ast}<\left(\frac{\left\|\mathbf{y}\right\|_{2}}{R}-1\right)\lambda_{1},$ (18)
where $v_{i}=\mathbf{e}_{i}^{\mathsf{T}}Q^{\mathsf{T}}\mathbf{y}$, $\mathbf{e}_{i}$ is the $i$-th standard basis vector, and $A=Q\Lambda Q^{\mathsf{T}}$ is the eigendecomposition, $\Lambda=\operatorname{diag}\left(\lambda_{1},\dots,\lambda_{d}\right)$ with $\lambda_{1}\geq\dots\geq\lambda_{d}>0$.
Proof of Lemma˜8.
The Mahalanobis projection onto an Euclidean ball is formulated as
[…]
s.t. $\mathbf{x}^{\mathsf{T}}\mathbf{x}\leq R^{2}$
The Lagrangian of this quadratic program is $\mathcal{L}=(\mathbf{x}-\mathbf{y})^{\mathsf{T}}A(\mathbf{x}-\mathbf{y})+\mu(\mathbf{x}^{\mathsf{T}}\mathbf{x}-R^{2})$. According to the KKT conditions, $\nabla\mathcal{L}=2A(\mathbf{x}-\mathbf{y})+2\mu\mathbf{x}=\bm{0}$ with $\mu>0$. (Otherwise, $\mu=0$ implies $\mathbf{y}\in\mathcal{B}(R)$ and the projection is trivial.) Rearranging $\nabla\mathcal{L}=\bm{0}$ yields $\mathbf{x}=(A+\mu I)^{-1}A\mathbf{y}$, thus $\mathbf{x}^{\mathsf{T}}\mathbf{x}=R^{2}$ is equivalent to $\rho(\mu)=0$, which proves Eq.˜17.
To prove Eq.˜18, since $\rho$ monotonically decreases on $0,\infty)$, it suffices to bound $\rho(\mu)$. By the orthogonality of $Q$, we have $\sum_{i=1}^{d}v_{i}^{2}=\left\|Q\mathbf{y}\right\|_{2}^{2}=\left\|\mathbf{y}\right\|_{2}^{2}$. Therefore,
$\frac{\left\|\mathbf{y}\right\|_{2}^{2}}{(1+\frac{\mu}{\lambda_{d}})^{2}}-R^{2}\leq\rho(\mu)=\sum_{i=1}^{d}\frac{v_{i}^{2}}{(1+\frac{\mu}{\lambda_{i}})^{2}}-R^{2}\leq\frac{\left\|\mathbf{y}\right\|_{2}^{2}}{(1+\frac{\mu}{\lambda_{1}})^{2}}-R^{2}.$
Then it is straightforward that $\rho((\frac{\left\|\mathbf{y}\right\|_{2}}{R}-1)\lambda_{d})\geq 0$ and $\rho((\frac{\left\|\mathbf{y}\right\|_{2}}{R}-1)\lambda_{1})\leq 0$, yielding that $\mu^{\ast}\in((\frac{\left\|\mathbf{y}\right\|_{2}}{R}-1)\lambda_{d},(\frac{\left\|\mathbf{y}\right\|_{2}}{R}-1)\lambda_{1})$ and $\mu^{\ast}$ is unique on $0,\infty)$. ∎
Lemma 9.
Let $\mathbf{x}^{\ast}=(A+\mu^{\ast}I)^{-1}A\mathbf{y}$, where $\rho(\mu^{\ast})=0$ as in Lemma˜8. If $\widetilde{\mu}\geq 0$ and $\left|\widetilde{\mu}-\mu^{\ast}\right|<\frac{\lambda_{d}}{\left\|\mathbf{y}\right\|_{2}}\zeta$, then $\mathbf{x}\in\mathcal{B}(R)$ and $\left\|\mathbf{x}-\mathbf{x}^{\ast}\right\|_{2}\leq\zeta$, where $\mathbf{x}$ is constructed from $\widetilde{\mu}$ as
$\mathbf{x}=\Pi_{\mathcal{B}(R)}\left[\widetilde{\mathbf{x}}\right]=\frac{R}{\left\|\widetilde{\mathbf{x}}\right\|_{2}}\cdot\widetilde{\mathbf{x}},\quad\widetilde{\mathbf{x}}=\left(A+\widetilde{\mu}I\right)^{-1}A\mathbf{y}.$ (19)
Proof of Lemma˜9.
Eq.˜19 immediately implies $\mathbf{x}\in\mathcal{B}(R)$.
To prove $\left\|\mathbf{x}-\mathbf{x}^{\ast}\right\|_{2}\leq\zeta$, since $\left\|\mathbf{x}-\mathbf{x}^{\ast}\right\|_{2}\leq\left\|\widetilde{\mathbf{x}}-\mathbf{x}^{\ast}\right\|_{2}$ by Lemma˜10, it suffices to prove $\left\|\widetilde{\mathbf{x}}-\mathbf{x}^{\ast}\right\|_{2}\leq\zeta$. By the definition of $v_{i}$ in Lemma˜8, we have
$\left\|\widetilde{\mathbf{x}}-\mathbf{x}^{\ast}\right\|_{2}^{2}$ $=\left\|\left(A+\widetilde{\mu}I\right)^{-1}A\mathbf{y}-\left(A+\mu^{\ast}I\right)^{-1}A\mathbf{y}\right\|_{2}^{2}=\sum_{i=1}^{d}v_{i}^{2}\left(\frac{1}{1+\frac{\widetilde{\mu}}{\lambda_{i}}}-\frac{1}{1+\frac{\mu^{\ast}}{\lambda_{i}}}\right)^{2}$
$\leq\sum_{i=1}^{d}v_{i}^{2}\left(1-\frac{1}{1+\frac{\left|\widetilde{\mu}-\mu^{\ast}\right|}{\lambda_{i}}}\right)^{2}\leq\sum_{i=1}^{d}v_{i}^{2}\left(\frac{\widetilde{\mu}-\mu^{\ast}}{\lambda_{i}}\right)^{2}\leq\left\|\mathbf{y}\right\|_{2}^{2}\left(\frac{\widetilde{\mu}-\mu^{\ast}}{\lambda_{d}}\right)^{2}<\zeta^{2}.$
The first inequality uses that $\lambda_{i}>0$, $\widetilde{\mu}\geq 0$ and $\mu^{\ast}\geq 0$; The second inequality comes from the following inequality, $(1-\frac{1}{1+a})^{2}=\frac{a^{2}}{(a+1)^{2}}\leq a^{2}$ for any $a\geq 0$; The third inequality uses that $\sum_{i=1}^{d}v_{i}^{2}=\left\|\mathbf{y}\right\|_{2}^{2}$; The last inequality uses that $|\widetilde{\mu}-\mu^{\ast}|<\frac{\lambda_{d}}{\left\|\mathbf{y}\right\|_{2}}\zeta$. ∎
With the help of these lemmas, we can prove Lemma˜7.
Proof of Lemma˜7.
First, we justify the bisection. By Lemma˜8, the wanted zero satisfies that $\mu^{\ast}\in(a_{1},b_{1})=((\frac{\left\|\mathbf{y}\right\|_{2}}{R}-1)\lambda_{d},(\frac{\left\|\mathbf{y}\right\|_{2}}{R}-1)\lambda_{1})$, which implies the initial interval of bisection (Line 3 in Algorithm˜4). Furthermore, $\rho$ monotonically decreases on $0,\infty)$, which implies the selection of $a_{t+1}$ and $b_{t+1}$ (Lines 4–6 in Algorithm˜4).
Next, we show the convergence of the bisection. With the interval length halving each iteration, based on the value of $T$ (Line 3 in Algorithm˜4), we have
$|\mu_{T+1}-\mu^{\ast}|\leq\frac{b_{1}-a_{1}}{2^{T}}=\frac{(\frac{\left\|\mathbf{y}\right\|_{2}}{R}-1)(\lambda_{1}-\lambda_{d})}{2^{T}}\leq\frac{(\frac{\left\|\mathbf{y}\right\|_{2}}{R}-1)(\overline{\lambda}-\underline{\lambda})}{2^{T}}\leq\frac{\underline{\lambda}}{\left\|\mathbf{y}\right\|_{2}}\zeta\leq\frac{\lambda_{d}}{\left\|\mathbf{y}\right\|_{2}}\zeta.$
Then by Lemma˜9, Algorithm˜4 achieves an error $\left\|\mathbf{x}-\mathbf{x}^{\ast}\right\|_{2}\leq\zeta$.
Finally, we bound the runtime. We note that choice 1 and 2 are equivalent, since
$\left\|\left(A+\mu I\right)^{-1}A\mathbf{y}\right\|_{2}^{2}=\left\|P\left(C+\mu I\right)^{-1}P^{\mathsf{T}}PCP^{\mathsf{T}}\mathbf{y}\right\|_{2}^{2}=\left\|\left(C+\mu I\right)^{-1}\mathbf{w}\right\|_{2}^{2}.$
With choice 1, each iteration requires $d^{\omega}$ arithmetic operations, where the computational bottleneck lies in the matrix inversion $(A+\mu I)^{-1}$, resulting in a runtime of $O(d^{\omega}n)$. With choice 2, the tridiagonalization of $A$ requires $O(d^{3})$ arithmetic operations [Golub and Van Loan, 2013], and each iteration can evaluate $(C+\mu I)^{-1}\mathbf{w}$ via the Thomas algorithm with only $O(d)$ arithmetic operations [Golub and Van Loan, 2013], resulting in a runtime of $O(d^{3}+dn)$. ∎
Appendix C Proofs for Section˜3
In this section, we prove the theoretical guarantees of LightONS.
C.1 Proof of Lemma˜2
Proof.
Let $\Phi_{T}$ denote the sum of squared norms of updates, as in the left-hand side of Eq.˜14 in Lemma˜6. Note that $A_{t}=\epsilon I+\sum_{i=1}^{t}\nabla f_{t}(\mathbf{x}_{t})\nabla f_{t}(\mathbf{x}_{t})^{\mathsf{T}}$ in Algorithm˜1 and $\left\|\nabla f_{t}(\mathbf{x}_{t})\right\|_{2}\leq G$ in Assumption˜2. Then by Lemma˜6 we have
$\Phi_{T}\triangleq\sum_{t=1}^{T}\left\|\frac{1}{\gamma}A_{t}^{-1}\nabla f_{t}(\mathbf{x}_{t})\right\|_{2}^{2}\leq\frac{1}{\gamma^{2}}\frac{d}{\epsilon}.$
We prove Lemma˜2 by bounding $N$ with $\Phi_{T}$. Let $\{\tau_{i}\}_{i=1}^{N}\subseteq[T]$ denote all the rounds when LightONS.Core performs the Mahalanobis projection. With $\tau_{0}=1$, we have that for any $i\in[N]$,
$\sum_{t=\tau_{i-1}}^{\tau_{i}-1}\frac{1}{\gamma}A_{t}^{-1}\nabla f_{t}(\mathbf{x}_{t})=\widehat{\mathbf{x}}_{\tau_{i}}-\mathbf{x}_{\tau_{i-1}}.$
Since $\left\|\widehat{\mathbf{x}}_{\tau_{i}+1}\right\|_{2}>kD/2$ and $\left\|\mathbf{x}_{\tau_{i}+1}\right\|_{2}=\left\|\Pi^{A_{\tau_{i}}}_{\mathcal{X}}[\widehat{\mathbf{x}}_{\tau_{i}+1}]\right\|_{2}\leq D/2$, we have
$\left\|\sum_{t=\tau_{i-1}}^{\tau_{i}-1}\frac{1}{\gamma}A_{t}^{-1}\nabla f_{t}(\mathbf{x}_{t})\right\|_{2}=\left\|\widehat{\mathbf{x}}_{\tau_{i}}-\mathbf{x}_{\tau_{i-1}}\right\|_{2}>(k-1)\frac{D}{2}.$
By applying the inequality $n\sum_{i=1}^{n}\left\|\mathbf{v}_{i}\right\|_{2}^{2}\geq\left\|\sum_{i=1}^{n}\mathbf{v}_{i}\right\|_{2}^{2}$, we obtain
$(\tau_{i}-\tau_{i-1})\frac{1}{\gamma^{2}}\sum_{t=\tau_{i-1}}^{\tau_{i}-1}\left\|A_{t}^{-1}\nabla f_{t}(\mathbf{x}_{t})\right\|_{2}^{2}\geq\left\|\sum_{t=\tau_{i-1}}^{\tau_{i}-1}\frac{1}{\gamma}A_{t}^{-1}\nabla f_{t}(\mathbf{x}_{t})\right\|_{2}^{2}>(k-1)^{2}\frac{D^{2}}{4}.$
Rearranging the preceding inequality yields
$\frac{1}{\tau_{i}-\tau_{i-1}}<\frac{4}{(k-1)^{2}D^{2}\gamma^{2}}\sum_{t=\tau_{i-1}}^{\tau_{i}-1}\left\|A_{t}^{-1}\nabla f_{t}(\mathbf{x}_{t})\right\|_{2}^{2}.$
By applying the inequality $\frac{n}{\sum_{i=1}^{n}\frac{1}{a_{i}}}\leq\frac{\sum_{i=1}^{n}a_{i}}{n}$ and the fact that $\tau_{N}\leq T$, we obtain
$\frac{N^{2}}{T}\leq\frac{N^{2}}{\tau_{N}-1}\leq\sum_{i=1}^{N}\frac{1}{\tau_{i}-\tau_{i-1}}<\frac{4}{(k-1)^{2}D^{2}\gamma^{2}}\sum_{t=1}^{\tau_{N}}\left\|A_{t}^{-1}\nabla f_{t}(\mathbf{x}_{t})\right\|_{2}^{2}\leq\frac{4}{(k-1)^{2}D^{2}\gamma^{2}}\Phi_{T}.$
Finally, rearranging the preceding inequality yields the desired result $N\leq\frac{2}{D\gamma}\sqrt{\frac{d}{\epsilon}T}$. ∎
We remark that Lemma˜2 is tight in terms of $T$ up to logarithmic factors. Consider $d=1$ and $\nabla f_{t}(\mathbf{x}_{t})=1/\sqrt{T}$, then $A_{t}^{-1}\nabla f_{t}(\mathbf{x}_{t})=\Omega(1/\sqrt{T})$ and $\|\sum_{t=1}^{T}A_{t}^{-1}\nabla f_{t}(\mathbf{x}_{t})\|_{2}=\Omega(\sqrt{T})$. We notice that Theorem 2.3 of Simchowit [2020] corroborates this point.
C.2 Proof of Theorem˜1
Proof.
The proof of Theorem˜1 follows as a specialization of the proof of Theorem˜2 given in Section˜C.4. The only substantive difference is that, in the present setting, the analysis is carried out directly on the original loss functions $f_{t}$, rather than on surrogate losses $g_{t}$.
We therefore omit the repetitive details and highlight only the key intermediate arguments.
  • Choice of the expanded curvature parameter $\gamma$ in LightONS.Core. The selection of $\gamma=\frac{1}{2}\min\{\frac{2}{(k+1)DG},\alpha\}$ directly follows from Lemma˜1.
  • Regret analysis with selective projections. The regret analysis follows the same structure as Lemma˜11, which accounts for projections triggered only when the iterate exits the expanded domain $\widetilde{\mathcal{X}}_{k}$. Details are given in Section˜C.4.
  • Runtime analysis with selective projections. The runtime bound is derived in the same manner as in the proof of Theorem˜2 in Section˜C.4.
C.3 Proof of Lemma˜4
Proof.
By the selection of curvature parameters $\gamma_{0}$ (in Lemma˜1) and $\gamma^{\prime}$ (in Algorithm˜3), it holds that $\gamma^{\prime}\leq\gamma_{0}$. Thus by relaxing Eq.˜2 in Lemma˜1 we have
$f_{t}(\mathbf{x}_{t})-f_{t}(\mathbf{u})$ $\leq\nabla f_{t}(\mathbf{x}_{t})^{\mathsf{T}}(\mathbf{x}_{t}-\mathbf{u})-\frac{\gamma_{0}}{2}\left(\nabla f_{t}(\mathbf{x}_{t})^{\mathsf{T}}(\mathbf{x}_{t}-\mathbf{u})\right)^{2}$
$\leq\nabla f_{t}(\mathbf{x}_{t})^{\mathsf{T}}(\mathbf{x}_{t}-\mathbf{u})-\frac{\gamma^{\prime}}{2}\left(\nabla f_{t}(\mathbf{x}_{t})^{\mathsf{T}}(\mathbf{x}_{t}-\mathbf{u})\right)^{2}.$
To finish the proof, it suffices to show that, with $U(a)=a-\frac{\gamma^{\prime}}{2}a^{2}$,
$U\left(\nabla f_{t}(\mathbf{x}_{t})^{\mathsf{T}}(\mathbf{x}_{t}-\mathbf{u})\right)\leq U\left(\nabla g_{t}(\mathbf{y}_{t})^{\mathsf{T}}(\mathbf{y}_{t}-\mathbf{u})\right).$
The preceding inequality holds because $U$ monotonically increases on $(-\infty,\frac{1}{\gamma^{\prime}}]$ and that
$\nabla f_{t}(\mathbf{x}_{t})^{\mathsf{T}}(\mathbf{x}_{t}-\mathbf{u})\leq c_{f}\nabla g_{t}(\mathbf{y}_{t})^{\mathsf{T}}(\mathbf{y}_{t}-\mathbf{u})\leq c_{f}G\cdot\max\left\{\frac{k+1}{2}c_{g}D,2c_{g}D\right\}\leq\frac{1}{\gamma^{\prime}}.$
The first inequality follows from Condition˜1; The second inequality from the Cauchy-Schwarz inequality $\mathbf{u}^{\mathsf{T}}\mathbf{v}\leq\left\|\mathbf{u}\right\|_{2}\left\|\mathbf{v}\right\|_{2}$; The last inequality from the selection of $\gamma^{\prime}$. ∎
C.4 Proof of Theorem˜2
We first introduce a property of the Mahalanobis projection. This lemma is often referred to as the Pythagorean theorem in Banach space or the non-expansiveness of projections.
Lemma 10 (Lemma 8 of Hazan et al. [2007]).
If $A\in\mathbb{R}^{d\times d}$ is positive-definite and symmetric matrix, then for any compact convex set $\mathcal{X}\subseteq\mathbb{R}^{d}$, any point $\mathbf{y}\in\mathbb{R}^{d}$ and any point $\mathbf{u}\in\mathcal{X}$, $\left\|\Pi^{A}_{\mathcal{X}}[\mathbf{y}]-\mathbf{u}\right\|_{A}\leq\left\|\mathbf{y}-\mathbf{u}\right\|_{A}$.
Based on Lemma˜10, we can decompose the regret of LightONS as the following lemma shows.
Lemma 11.
Ignoring the truncation error of Mahalanobis projections, under the same assumptions as Theorem˜2, in Algorithm˜3, for any $t\in[T]$ and all $\mathbf{u}\in\mathcal{X}$, it holds that
$2\nabla g_{t}(\mathbf{y}_{t})^{\mathsf{T}}(\mathbf{y}_{t}-\mathbf{u})\leq\frac{1}{\gamma^{\prime}}\left\|\nabla g_{t}(\mathbf{y}_{t})\right\|_{A_{t}^{-1}}^{2}+\gamma^{\prime}\left\|\mathbf{y}_{t}-\mathbf{u}\right\|_{A_{t}}^{2}-\gamma^{\prime}\left\|\mathbf{y}_{t+1}-\mathbf{u}\right\|_{A_{t}}^{2}.$ (20)
Proof of Lemma˜11.
When $\left\|\widehat{\mathbf{y}}_{t+1}\right\|_{2}\leq kD/2$, no Mahalanobis projection is triggered, i.e., $\mathbf{y}_{t+1}=\widehat{\mathbf{y}}_{t+1}=\mathbf{y}_{t}-\frac{1}{\gamma^{\prime}}A_{t}^{-1}\nabla g_{t}(\mathbf{y}_{t})$, then for any $\mathbf{u}\in\mathbb{R}^{d}$, we have
$\left\|\mathbf{y}_{t+1}-\mathbf{u}\right\|_{A_{t}}^{2}=\left\|\mathbf{y}_{t}-\frac{1}{\gamma^{\prime}}A_{t}^{-1}\nabla g_{t}(\mathbf{y}_{t})-\mathbf{u}\right\|_{A_{t}}^{2}.$
Otherwise, when $\left\|\widehat{\mathbf{y}}_{t+1}\right\|_{2}>kD/2$ and the Mahalanobis projection is triggered, by Lemma˜10, for any $\mathbf{u}\in\mathcal{X}$, we have
$\left\|\mathbf{y}_{t+1}-\mathbf{u}\right\|_{A_{t}}^{2}=\left\|\Pi_{\mathcal{B}(D/2)}^{A_{t}}\left[\mathbf{y}_{t}-\frac{1}{\gamma^{\prime}}A_{t}^{-1}\nabla g_{t}(\mathbf{y}_{t})\right]-\mathbf{u}\right\|_{A_{t}}^{2}\leq\left\|\mathbf{y}_{t}-\frac{1}{\gamma^{\prime}}A_{t}^{-1}\nabla g_{t}(\mathbf{y}_{t})-\mathbf{u}\right\|_{A_{t}}^{2}.$
Combining both cases, we conclude that, for any $t\in[T]$ and any $\mathbf{u}\in\mathcal{X}$,
$\left\|\mathbf{y}_{t+1}-\mathbf{u}\right\|_{A_{t}}^{2}\leq\left\|\left(\mathbf{y}_{t}-\mathbf{u}\right)-\frac{1}{\gamma^{\prime}}A_{t}^{-1}\nabla g_{t}(\mathbf{y}_{t})\right\|_{A_{t}}^{2}.$
Rearranging the preceding inequality yields the desired result of Eq.˜20. ∎
We note that Lemma˜11 ignores the truncation error of the Mahalanobis projections. The matrix factorization underlying Mahalanobis projections is related to eigendecomposition, equivalent to finding roots of a $d$-degree polynomial, which is not exactly solvable by finitely many arithmetic operations when $d\geq 5$. The next lemma complements the analysis. Thanks to FastProj (Algorithm˜4 in Section˜B.3), the truncation error only incurs a negligible additive $O(1/t^{2})$ term in the regret decomposition of Lemma˜11, which can be safely ignored in the final regret bound of LightONS.
Lemma 12.
Let the Mahalanobis projection of Algorithm˜3 be implemented with Algorithm˜4 with $A=A_{t}$, $R=D/2$, $\mathbf{u}=\widehat{\mathbf{y}}_{t+1}$, $\zeta=\zeta_{t}$, $\underline{\lambda}=\epsilon$, $\overline{\lambda}=\epsilon+G^{2}t$, where $\zeta_{t}$ is defined as
$\zeta_{t}=\min\left\{\frac{\gamma^{\prime}}{2kD(c_{g}^{2}G^{2}t+\epsilon)t^{2}},\frac{1}{t}\sqrt{\frac{\gamma^{\prime}}{2(c_{g}^{2}G^{2}t+\epsilon)t}}\right\}.$ (21)
Then under the same assumptions as Theorem˜2, in Algorithm˜3, for any $t\in[T]$ and all $\mathbf{u}\in\mathcal{X}$,
$2\nabla g_{t}(\mathbf{y}_{t})^{\mathsf{T}}(\mathbf{y}_{t}-\mathbf{u})\leq\frac{1}{\gamma^{\prime}}\left\|\nabla g_{t}(\mathbf{y}_{t})\right\|_{A_{t}^{-1}}^{2}+\gamma^{\prime}\left\|\mathbf{y}_{t}-\mathbf{u}\right\|_{A_{t}}^{2}-\gamma^{\prime}\left\|\mathbf{y}_{t+1}-\mathbf{u}\right\|_{A_{t}}^{2}+\frac{1}{t^{2}}.$ (22)
Proof of Lemma˜12.
It suffices to prove the case when $\left\|\widehat{\mathbf{y}}_{t+1}\right\|_{2}>kD/2$. Let $\mathbf{y}^{\ast}_{t+1}=\Pi_{\mathcal{B}(D/2)}^{A_{t}}\left[\widehat{\mathbf{y}}_{t+1}\right]$ be the exact Mahalanobis projection without truncation error, and $\bm{\delta}_{t}=\mathbf{y}^{\ast}_{t+1}-\mathbf{y}_{t+1}$ be the truncation error. By Lemma˜9, $\left\|\bm{\delta}_{t}\right\|_{2}=\left\|\mathbf{y}_{t+1}-\mathbf{y}^{\ast}_{t+1}\right\|_{2}\leq\zeta_{t}$; By Lemma˜10, for any $\mathbf{u}\in\mathcal{X}$,
$\left\|\mathbf{y}_{t+1}-\mathbf{u}\right\|_{A_{t}}^{2}$ $=\left\|\mathbf{y}^{\ast}_{t+1}-\mathbf{u}\right\|_{A_{t}}^{2}+2\left(\mathbf{y}^{\ast}_{t+1}-\mathbf{u}\right)^{\mathsf{T}}A_{t}\bm{\delta}_{t}+\left\|\bm{\delta}_{t}\right\|_{A_{t}}^{2}$
$\leq\left\|\mathbf{y}^{\ast}_{t+1}-\mathbf{u}\right\|_{A_{t}}^{2}+2\cdot 2D\cdot(c_{g}^{2}G^{2}T+\epsilon)\cdot\left\|\bm{\delta}_{t}\right\|_{2}+(c_{g}^{2}G^{2}T+\epsilon)\cdot\left\|\bm{\delta}_{t}\right\|_{2}^{2}$
$\leq\left\|\mathbf{y}_{t}-\frac{1}{\gamma^{\prime}}A_{t}^{-1}\nabla g_{t}(\mathbf{y}_{t})-\mathbf{u}\right\|_{A_{t}}^{2}+\frac{2\gamma^{\prime}}{t^{2}}.$
The first inequality considers the operator norm of $A_{t}$ and uses that $\left\|\nabla g_{t}(\mathbf{y}_{t})\right\|_{2}\leq c_{g}G$ (Condition˜1); The second inequality comes from the selection of $\zeta_{t}$ in Eq.˜21. ∎
For conciseness, we ignore the truncation error of the Mahalanobis projections in the proof of Theorem˜2, since it only incurs an additive $O(\sum_{t=1}^{T}1/t^{2})=O(1)$ term in the final regret bound.
Proof of Theorem˜2.
Plugging Lemma˜11 into Lemma˜4 yields the following inequalities:
[…]
Telescoping the right-hand side establishes the desired regret of Eq.˜10a:
$\sum_{t=1}^{T}\left(f_{t}(\mathbf{x}_{t})-f_{t}(\mathbf{u})\right)$ $\leq\sum_{t=1}^{T}\left(\frac{1}{2\gamma^{\prime}}\left\|\nabla g_{t}(\mathbf{y}_{t})\right\|_{A_{t}^{-1}}^{2}+\frac{\gamma^{\prime}}{2}\left\|\mathbf{y}_{t}-\mathbf{u}\right\|_{A_{t-1}}^{2}-\frac{\gamma^{\prime}}{2}\left\|\mathbf{y}_{t+1}-\mathbf{u}\right\|_{A_{t}}^{2}\right)$
$=\left(\frac{1}{2\gamma^{\prime}}\sum_{t=1}^{T}\left\|\nabla g_{t}(\mathbf{y}_{t})\right\|_{A_{t}^{-1}}^{2}\right)+\frac{\gamma^{\prime}}{2}\left\|\mathbf{y}_{1}-\mathbf{u}\right\|_{A_{0}}^{2}-\frac{\gamma^{\prime}}{2}\left\|\mathbf{y}_{T+1}-\mathbf{u}\right\|_{A_{T}}^{2}$
[…]
Finally, the desired runtime of Eq.˜10b follows from the following two parts:
  • Runtime aside from FastProj. In each round, the domain conversion of Lemma˜3 takes $O(\texttt{EP}_{\mathcal{X}}+d)$ time, updating and inverting $A_{t}$ as Eq.˜16 takes $O(d^{2})$ time, and other operations, such as computing the surrogate gradient as Eq.˜8b, take only $O(d)$ time. The overall runtime aside from FastProj is $O((\texttt{EP}_{\mathcal{X}}+d^{2})T)$.
  • Runtime of FastProj. By Lemma˜7, the number of bisections $n_{t}=O(\log t)=O(\log T)$ and each bisection takes $O(d^{\omega})$ time with choice 1. Since LightONS.Core is a subroutine in LightONS, the number of calls to FastProj is at most $O((k-1)^{-1}d^{0.5}\sqrt{T/\epsilon})$ by Lemma˜2. The overall runtime of FastProj is $O((k-1)^{-1}d^{0.5}\sqrt{T/\epsilon}\cdot d^{\omega}\log T)$.
Appendix D Proofs for Section˜4
In this section, we provide the proofs and details for Section˜4.
D.1 Proof of Theorem˜3
Before proving Theorem˜3, we first give a more detailed presentation of the convergence rates induced by LightONS. With $T=\Theta\left(\frac{d}{\epsilon}\log\frac{d}{\epsilon}\log\frac{1}{\delta}\right)$, we have
$F(\bar{\mathbf{x}}_{T})-\min_{\mathbf{x}\in\mathcal{X}}F(\mathbf{x})\leq\frac{1}{T}\left(\textsc{Reg}_{T}+4\sqrt{\frac{\textsc{Reg}_{T}}{2\gamma_{0}}\log\frac{4\log T}{\delta}}+\frac{8}{\gamma_{0}}\log\frac{4\log T}{\delta}\right)=O(\epsilon),$ (23a)
$\textsc{Runtime}\leq O\left(\frac{d^{3}}{\epsilon}\log\frac{d}{\epsilon}\log\frac{1}{\delta}+\frac{d^{3.5}}{\sqrt{\epsilon}}\log\frac{d}{\epsilon}\log\frac{1}{\delta}\right)=\widetilde{O}\left(\frac{d^{3}}{\epsilon}\right),$ (23b)
where $\textsc{Reg}_{T}$ follows Eq.˜11a in Corollary˜1. With $\frac{1}{\delta}=T^{\prime}=\Theta\left(\frac{d}{\epsilon}\log\frac{d}{\epsilon}\right)$, we have
$\mathop{\mathbb{E}}\left[F(\bar{\mathbf{x}}_{T^{\prime}})-\min_{\mathbf{x}\in\mathcal{X}}F(\mathbf{x})\right]\leq O(\epsilon),$ (24a)
$\textsc{Runtime}\leq O\left(\frac{d^{3}}{\epsilon}\log\frac{d}{\epsilon}+\frac{d^{3.5}}{\sqrt{\epsilon}}\log\frac{d}{\epsilon}\right)=\widetilde{O}\left(\frac{d^{3}}{\epsilon}\right).$ (24b)
Proof.
First, we prove Eq.˜23a. Without loss of generality, we consider $T\geq 3$. Then the high-probability excess risk bound of Eq.˜23a directly follows from Corollary 2 of Mehta [2017] by substituting LightONS’s regret in Corollary˜1.
Then, we verify that the choice of $T=\Theta\left(\frac{d}{\epsilon}\log\frac{d}{\epsilon}\log\frac{1}{\delta}\right)$ yields a high-probability excess risk of $O(\epsilon)$. Let $\Gamma_{\delta}=\frac{4\log T}{\delta}=O\left(\frac{1}{\delta}\log\frac{d}{\epsilon}+\frac{1}{\delta}\log\log\frac{1}{\delta}\right)$, we have
$F(\bar{\mathbf{x}}_{T})-\min_{\mathbf{x}\in\mathcal{X}}F(\mathbf{x})\leq O\left(\frac{\textsc{Reg}_{T}+\sqrt{\textsc{Reg}_{T}\cdot\log\Gamma_{\delta}}+\log\Gamma_{\delta}}{T}\right)\leq O\left(\frac{\textsc{Reg}_{T}+\log\Gamma_{\delta}}{T}\right).$
The first inequality is essentially Corollary 2 of Mehta [2017], and the second inequality follows from the fact that $\sqrt{ab}\leq\frac{a+b}{2}=O(a+b)$ for any positive terms $a$ and $b$. Since $\textsc{Reg}_{T}=O(d\log T)=O\left(d\log\frac{d}{\epsilon}+d\log\log\frac{1}{\delta}\right)$, it suffices to show that
$O\left(\frac{\textsc{Reg}_{T}}{T}\right)$ $=O\left(\frac{d\left(\log\frac{d}{\epsilon}+\log\log\frac{1}{\delta}\right)}{\frac{d}{\epsilon}\log\frac{d}{\epsilon}\log\frac{1}{\delta}}\right)=O(\epsilon),\penalty 10000\ \penalty 10000\ \text{and}$
$O\left(\frac{\log\Gamma_{\delta}}{T}\right)$ $=O\left(\frac{\log\frac{1}{\delta}+\log\left(\log\frac{d}{\epsilon}+\log\log\frac{1}{\delta}\right)}{\frac{d}{\epsilon}\log\frac{d}{\epsilon}\log\frac{1}{\delta}}\right)=O\left(\frac{\epsilon}{d}\right)=O(\epsilon).$
Next, we prove Eq.˜24a from Eq.˜23a. By the definition of expectation, we have
$\mathop{\mathbb{E}}\left[F(\bar{\mathbf{x}}_{T^{\prime}})-\min_{\mathbf{x}\in\mathcal{X}}F(\mathbf{x})\right]$ $\leq\left(1-\frac{1}{T^{\prime}}\right)\cdot O\left(\frac{\textsc{Reg}_{T^{\prime}}+\sqrt{\textsc{Reg}_{T^{\prime}}\cdot\log\Gamma_{1/T^{\prime}}}+\Gamma_{1/T^{\prime}}}{T^{\prime}}\right)$
$\quad+\frac{1}{T^{\prime}}\left(\max_{\mathbf{y}\in\mathcal{X}}F(\mathbf{y})-\min_{\mathbf{x}\in\mathcal{X}}F(\mathbf{x})\right)$
$\leq O\left(\frac{\textsc{Reg}_{T^{\prime}}+\Gamma_{1/T^{\prime}}}{T^{\prime}}\right)+\frac{DG}{T^{\prime}}\leq O\left(\frac{\textsc{Reg}_{T^{\prime}}+\Gamma_{1/T^{\prime}}}{T^{\prime}}\right).$
The above inequalities use the Lipschitzness of $F$ and the boundedness of $\mathcal{X}$. Similarly, we verify that the choice of $T^{\prime}=\Theta\left(\frac{d}{\epsilon}\log\frac{d}{\epsilon}\right)$ yields an in-expectation excess risk of $O(\epsilon)$ with $\frac{1}{\delta}=T^{\prime}=\Theta\left(\frac{d}{\epsilon}\log\frac{d}{\epsilon}\right)$. It suffices to note that $\Gamma_{1/T^{\prime}}=4T^{\prime}\log T^{\prime}=O\left(\frac{d}{\epsilon}\log^{2}\frac{d}{\epsilon}\right)$ and that
$O\left(\frac{\log\Gamma_{1/T^{\prime}}}{T^{\prime}}\right)=O\left(\frac{\log\frac{d}{\epsilon}}{\frac{d}{\epsilon}\log\frac{d}{\epsilon}}\right)=O\left(\frac{\epsilon}{d}\right)=O(\epsilon).$
Finally, the total runtime in Eqs.˜23b and 24b follows from Theorem˜2. ∎
D.2 Proofs and Details for ERM-based SXO Methods
In this subsection, we provide two observations which support our conjecture in Section˜4.2 that ERM-based SXO methods fail to surpass the $\widetilde{O}(d^{3}/\epsilon)$ runtime barrier, even when these methods use more working memory. Moreover, we give an intuitive explanation of the runtime barrier $\widetilde{O}(d^{3}/\epsilon)$.
Fast matrix multiplication.
Linear regression with random design, a special case of SXO, reduces to solving a linear system with $d$ variables and $\widetilde{O}(d/\epsilon)$ equations. Although fast matrix multiplication accelerates this to $\widetilde{O}(d^{\omega}/\epsilon)$ runtime [Ibarra et al., 1982], this runtime does not plausibly extend beyond linear regression and reverts to $\widetilde{O}(d^{3}/\epsilon)$ in practice.
Cutting-plane methods.
The following proposition constructs an SXO instance where applying CPM correspondingly incurs $\widetilde{O}(d^{3}/\epsilon)$ runtime. We remark that, apart from the runtime, CPM further assumes well-roundedness of the domain, which may incur additional computational overhead.
Proposition 3.
Under Assumption˜4 and that $\mathcal{X}=\mathcal{B}(1)$, let the stochastic functions take the form $f(\mathbf{x};\xi)=\phi\left(\mathbf{w}(\xi)^{\mathsf{T}}\mathbf{x}\right)$, where $\phi:\mathbb{R}\to\mathbb{R}$ is a black-box function, and $\mathbf{w}:\Xi\to\mathbb{R}^{d}$ is a fixed mapping. By [Koren and Levy, 2015, Mehta, 2017], obtaining an $\epsilon$-optimal solution for SXO reduces to obtaining an $O(\epsilon)$-optimal minimizer to the offline $\alpha$-exp-concave objective
$\widehat{F}(\mathbf{x})=\frac{1}{T}\sum_{t=1}^{T}\phi\left(\mathbf{w}(\xi_{t})^{\mathsf{T}}\mathbf{x}\right),$
where $T=\widetilde{O}(d/\epsilon)$ is the necessary sample size. Cutting-Plane Methods (CPM) of Lee et al. [2015], Jiang et al. [2020] solves this offline problem to $O(\epsilon)$-accuracy in $\widetilde{O}(d^{3}/\epsilon)$ time.
Proof of Proposition˜3.
Best known CPM of Lee et al. [2015], Jiang et al. [2020] query the gradient of $\widehat{F}$ for $\widetilde{O}(d)$ times, and each gradient query costs $O(dT)$ time due to the finite-sum structure of $\widehat{F}$ and the black-box nature of $\phi$. Therefore, the total runtime is $\widetilde{O}(d^{2}T+d^{3})=\widetilde{O}(d^{3}/\epsilon)$. ∎
An intuitive explanation of the runtime $\widetilde{O}(d^{3}/\epsilon)$.
With the online-to-batch conversion, OGD’s regret $O(\sqrt{T})$ for OCO translates to a sample complexity of $O(1/\epsilon^{2})$ and a runtime of $\widetilde{O}(d/\epsilon^{2})$ for SXO. When the target excess risk $\epsilon$ is relatively large, i.e., $\epsilon=\widetilde{\Omega}(1/d^{2})$, the OGD-based SXO method may circumvent the $\widetilde{O}(d^{3}/\epsilon)$ runtime barrier. Thus the computational challenge of SXO mainly manifests when pursuing the (near) optimal sample complexity of $\widetilde{O}(d/\epsilon)$.
Lemma˜1 indicates that an exp-concave function can be regarded as a strongly convex function in the direction of its gradient. Comparing the sample complexity lower bound $\Omega(d/\epsilon)$ for SXO [Mahdavi et al., 2015] to the $\Omega(1/\epsilon)$ rate for stochastic strongly convex optimization [Hazan and Kale, 2011], the exp-concave offline objective $F$ can be viewed as the sum of $d$ strongly convex objectives on $d$ directions, one along each dimension corresponding to the eigenvectors of the Hessian-related matrix $A_{t}$ in ONS.
Our intuition is as follows, the difficulty of pursuing the (near) optimal sample complexity arises from exploiting the directional strong convexity, which necessitates estimating the covariance structure of the stochastic gradients. Cai et al. [2010] establish a minimax sample complexity for covariance matrix estimation under squared Frobenius norm, which coincides with the sample complexity lower bound for SXO. Specifically, for some family of covariance matrices $S\SS$, with $T$ samples i.i.d. sampled from a distribution with covariance $S\Sigma\in\SS$, the estimator $\widehat{\Sigma}$ satisfies
[…]
We refer readers to Theorem 1 of Cai et al. [2010] for a formal statement. Since computing the sample covariance matrix requires $O(d^{2}T)$ time, estimating the covariance structure of stochastic gradients plausibly incurs at least $O(d^{2}T)$ runtime, leading to the $\widetilde{O}(d^{3}/\epsilon)$ runtime barrier.
Appendix E Proofs for Section˜5
In this section, we provide algorithms and proofs for LightONS’s applications listed in Section˜5.
E.1 Proofs and Details for Gradient-Norm Adaptivity
In this subsection, we target OXO regret bounds that scale with the accumulated squared gradient norms $G_{T}$ instead of the time horizon $T$, i.e.,
$\textsc{Reg}_{T}(\mathbf{u})=O\left(d\log G_{T}\right),\quad\text{where}\penalty 10000\ G_{T}\triangleq\sum_{t=1}^{T}\left\|\nabla f_{t}(\mathbf{x}_{t})\right\|_{2}^{2}.$ (25)
In benign environments, $G_{T}$ can be $O(1)$, for example, when the loss is nearly fixed, decisions converge quickly, yielding regret far below worst-case bounds that scale with $T$. Yet, since $G_{T}\leq G^{2}T$, the gradient-norm adaptive bound safeguards the minimax-optimality in $T$. Prior work has leveraged Eq.˜25 to achieve such stronger types of adaptivity, including small-loss adaptivity for OXO with smoothness and comparator-norm adaptivity for unbounded OCO, as discussed below.
OXO with smoothness.
Orabona et al. [2012] demonstrate that ONS directly implies the desired regret bound in Eq.˜25 for OXO. Under smoothness assumptions, the gradient-norm adaptivity can further be improved to small-loss adaptivity [Srebro et al., 2010, Zhao et al., 2020], where the regret scales with the cumulative loss of the best comparator in hindsight, regardless of online decisions. This is formalized as follows:
$\textsc{Reg}_{T}=O\left(d\log L_{T}\right),\quad\text{where}\penalty 10000\ L_{T}\triangleq\min_{\mathbf{u}\in\mathcal{X}}\sum_{t=1}^{T}\left(f_{t}(\mathbf{u})-\min_{\mathbf{x}\in\mathcal{X}}f_{t}(\mathbf{x})\right),$ (26)
when the loss $f_{t}$ is $H$-smooth on $\mathcal{X}$ for any $t\in[T]$, i.e., $\left\|\nabla f(\mathbf{x})-\nabla f(\mathbf{y})\right\|_{2}\leq H\left\|\mathbf{x}-\mathbf{y}\right\|_{2}$ for any pair $(\mathbf{x},\mathbf{y})\in\mathcal{X}^{2}$.
Unbounded OCO.
Cutkosky and Orabona [2018] extend the gradient-norm adaptive OXO regret in Eq.˜25 to a comparator-norm adaptive OCO regret, i.e.,
$\textsc{Reg}_{T}(\mathbf{u})=\widetilde{O}\left(\left\|\mathbf{u}\right\|_{2}\sqrt{dG_{T}}\right),\quad\text{for any unbounded comparator}\penalty 10000\ \mathbf{u}\in\mathbb{R}^{d},$ (27)
without any prior knowledge of $\mathbf{u}$. Such guarantees for unbounded comparators are unachievable via classical OCO algorithms that explicitly depend on the domain diameter $D$, for example, OGD demands an explicit $D$ to obtain $O(DG\sqrt{T})$ regret [Zinkevich, 2003, Abernethy et al., 2008].
Nevertheless, both [Orabona et al., 2012, Cutkosky and Orabona, 2018] incur the worst-case $\widetilde{O}(d^{\omega}T)$ runtime bottleneck of ONS. Below, we show that LightONS can replace ONS in these algorithms to substantially improve efficiency without compromising statistical guarantees.
Improvements by LightONS.
LightONSrecovers ONS’s gradient-norm adaptive regret bound in Eq.˜25, thereby leading to small-loss adaptivity for OXO with smoothness, as well as comparator-norm adaptivity for unbounded OCO. We state the theoretical guarantees based on LightONS and omit those of the original ONS-based methods as they are identical except for the runtime.
Theorem 4 (LightONS’s improvement for OXO with smoothness).
Under Assumptions˜1, 2 and 3, and that $f_{t}$ is $H$-smooth for any $t\in[T]$, LightONS (Algorithm˜3) satisfies that,
$\textsc{Reg}_{T}$ $\leq\frac{d}{2\gamma_{0}}\log\left(\frac{8H}{d\epsilon}L_{T}+\frac{4H}{\gamma_{0}\epsilon}\log\frac{4H}{\mathrm{e}\gamma_{0}\epsilon}+\frac{\gamma_{0}D^{2}H}{d}+2\right)+\frac{\gamma_{0}\epsilon D^{2}}{8},$
Runtime $\leq O\left(\left(\texttt{EP}_{\mathcal{X}}+d^{2}\right)T+d^{\omega}\sqrt{T}\log T\right),$
where $L_{T}$ is defined in Eq.˜26, and $\gamma_{0}$ is defined in Corollary˜1.
Proof of Theorem˜4.
Based on the proof of Theorem 1 of Orabona et al. [2012], it suffices to bound the gradient norms of the surrogate loss by those of the original loss. We note that Appendix B.4.2 of Yang et al. [2024] also discusses the small-loss bounds of ONS with domain conversion, although with a different domain conversion. Since $\left\|\nabla g_{t}(\mathbf{y}_{t})\right\|_{2}\leq\left\|\nabla f_{t}(\mathbf{x}_{t})\right\|_{2}$ by Lemma˜3, we have
$G_{T,g}\triangleq\sum_{t=1}^{T}\left\|\nabla g_{t}(\mathbf{y}_{t})\right\|_{2}^{2}\leq\sum_{t=1}^{T}\left\|\nabla f_{t}(\mathbf{x}_{t})\right\|_{2}^{2}\triangleq G_{T,f}.$
Theorem˜2 implies the runtime with $\epsilon=d$ and the regret of
$\textsc{Reg}_{T}(\mathbf{u})\leq\frac{d}{2\gamma_{0}}\log\left(1+\frac{G_{T,g}}{d\epsilon}\right)+\frac{\gamma_{0}\epsilon D^{2}}{8}\leq\frac{d}{2\gamma_{0}}\log\left(1+\frac{G_{T,f}}{d\epsilon}\right)+\frac{\gamma_{0}\epsilon D^{2}}{8}.$
The preceding inequality uses Jensen’s inequality and $\operatorname{tr}(A_{T})=d\epsilon+G_{T,g}$, i.e.,
$\log\operatorname{det}(A_{T})-\log\operatorname{det}(A_{0})\leq d\log\frac{\operatorname{tr}(A_{T})}{d}-\log\operatorname{det}(A_{0})=d\log\left(1+\frac{G_{T,g}}{d\epsilon}\right).$ (28)
Then Corollary 5 of Orabona et al. [2012] converts the gradient-norm adaptive bound to the small-loss bound with respect to $L_{T}$ and completes the proof. ∎
Theorem 5 (LightONS’s improvement for unbounded OCO).
If the loss function $f_{t}:\mathbb{R}^{d}\to\mathbb{R}$ is convex and $1$-Lipschitz for any $t\in[T]$, then there exists an algorithm satisfying that, for any $\mathbf{u}\in\mathbb{R}^{d}$,
$\textsc{Reg}_{T}(\mathbf{u})$ $\leq\widetilde{O}\left(\sqrt{d\sum_{t=1}^{T}\left(\nabla f_{t}(\mathbf{x}_{t})^{\mathsf{T}}\mathbf{u}\right)^{2}}\right)\leq\widetilde{O}\left(\left\|\mathbf{u}\right\|_{2}\sqrt{dG_{T}}\right),$
Runtime $\leq O\left(d^{2}T+d^{\omega}\sqrt{T}\log T\right).$
Proof of Theorem˜5.
Theorem 8 of Cutkosky and Orabona [2018] hinges on their Algorithm 7, Lemmas 16 and 17, apart from the coin-betting framework. To prove Theorem˜5, we show how LightONS adapts their analysis with minimal changes.
  • Modifications to their Algorithm 7. Since the decisions of their ONS are intermediate decisions to maximize the “wealth” in the coin-betting framework instead of true decisions to minimize regret, we can ignore the improper-to-proper conversion and replace their ONS with LightONS.Core. Specifically, their ONS runs on $\mathcal{X}=\mathcal{B}(1/2)$ while LightONS.Core runs on $\mathcal{Y}=\mathcal{B}(3/4)$ with the deferral coefficient $k=3/2$.
  • Modifications to their Lemma 16. Their ONS’s domain $\mathcal{X}=\mathcal{B}(1/2)$ implies a curvature parameter $\gamma=\frac{2-\log 3}{2}$ while LightONS.Core’s domain $\mathcal{Y}=\mathcal{B}(3/4)$ implies $\gamma=\frac{6-\log 7}{18}$.
  • Modifications to their Lemma 17. Relaxing the radius from $1/2$ to $3/4$ enlarges constants in the regret. Nonetheless, the numerical constants in their Lemma 17 are loose enough to accommodate our changes, greatly simplifying our analysis. Following their proof, let $\mathbf{z}_{t}$ denote the gradient that LightONS.Core receives at time $t$, Theorem˜2 and Eq.˜28 imply
    $\textsc{Reg}_{T}\leq\frac{d}{2\gamma}\log\left(1+\frac{\sum_{t=1}^{T}\left\|\mathbf{z}_{t}\right\|_{2}^{2}}{d\epsilon}\right)+\frac{\gamma\epsilon D^{2}}{8}.$
    Plugging $D=1$, $\gamma=\frac{6-\log 7}{18}\in(0.225,0.226)$, $\epsilon=d$, and $\left\|\mathbf{z}_{t}\right\|_{2}^{2}\leq 16\left\|\nabla f_{t}(\mathbf{x}_{t})\right\|_{2}^{2}$ yields
    $\textsc{Reg}_{T}\leq d\left(\frac{5}{2}\log\left(1+\frac{16}{d^{2}}\sum_{t=1}^{T}\left\|\nabla f_{t}(\mathbf{x}_{t})\right\|_{2}^{2}\right)+\frac{1}{35}\right).$
    The preceding regret bound fully recovers their Lemma 17 when $d\geq 2$.
The runtime follows from Theorem˜2 with $\epsilon=d$. ∎
Comparison with OQNS.
We remark that OQNS can hardly achieve full gradient-norm adaptivity, as the log-barrier regularization introduces an unavoidable $O(\log T)$ bias term independent of $G_{T}$. From the perspective of OMD, the regret can be decomposed into a stability term and a bias term. Adopted from Hazan [2016], Orabona [2019], the regret decomposition of a typical OMD algorithm is as follows:
$\textsc{Reg}_{T}(\mathbf{u})\leq\underbrace{\frac{\eta}{2}\sum_{t=1}^{T}\left\|\nabla f_{t}(\mathbf{x}_{t})\right\|_{\ast}^{2}}_{\text{stability}}+\underbrace{\frac{\left\|\mathbf{u}-\mathbf{x}_{1}\right\|^{2}}{2\eta}}_{\text{bias}}.$ (29)
While ONS and LightONS, as instances of OMD, can flexibly balance stability and bias,  [^3] For ONS and LightONS, the learning rate $\eta$ is replaced with the time-varying Hessian-related matrix’s inverse $A_{t}^{-1}$. OQNS sacrifices such flexibility for computational efficiency. OQNS employs a highly stable log-barrier, which suppresses the stability term but inflates the bias term in a manner resistant to small gradient norms. Specifically, the log-barrier contributes a term $\log(D^{2}-\left\|\mathbf{u}\right\|_{2}^{2})$ in the regret for any comparator $\mathbf{u}\in\mathcal{B}(D)$. As $\left\|\mathbf{u}\right\|_{2}$ approaches $D$, this term diverges to infinity. This issue is mitigated with the fixed-share trick [Mhammedi and Gatmiry, 2023], an illustration of which is as follows:
$\sum_{t=1}^{T}\left(f_{t}(\mathbf{x}_{t})-f_{t}(\mathbf{u})\right)\leq\underbrace{\sum_{t=1}^{T}\left(f_{t}(\mathbf{x}_{t})-f_{t}(\mathbf{v})\right)}_{\text{fixed-share regret}}+\underbrace{\sum_{t=1}^{T}\left(f_{t}(\mathbf{v})-f_{t}(\mathbf{u})\right)}_{\text{fixed-share margin}},\penalty 10000\ \penalty 10000\ \text{where}\penalty 10000\ \mathbf{v}=\left(1-\frac{1}{T}\right)\mathbf{u}.$
The fixed-share margin term $f_{t}(\mathbf{v})-f_{t}(\mathbf{u})\leq O(GD/T)$ due to boundedness and Lipschitzness, summing to $O(1)$ over $T$ rounds. Since $1-\left\|\mathbf{v}\right\|_{2}^{2}\geq\Omega(D^{2}/T)$, the fixed-share regret term contributes $\log(D^{2}-\left\|\mathbf{v}\right\|_{2}^{2})=O(\log T)$ to the regret, which prevents full gradient-norm adaptivity.
E.2 Proofs and Details for Logistic Bandits
In this subsection, we focus on logistic bandits as a representative instance of generalized linear bandits (GLB) to concretely illustrate the applicability of LightONS. GLB constitutes an important class of parametric stochastic bandits, where the expected loss depends on an unknown parameter through a known link function. This formulation introduces non-linearities into linear bandits, enhancing their expressivity while retaining tractable solutions.
Problem setting.
Logistic bandits can be interpreted as interactions between a learner and an adversary, which unfolds as follows: At each round $t\in[T]$, the learner selects an arm $\mathbf{x}_{t}$ from an arm set $\mathcal{X}_{t}\subseteq\mathcal{B}(1)^{K}$ with $K$ arms, and suffers a stochastic loss $y_{t}\in\{0,1\}$, where $\mathop{\mathbb{P}}\left(y_{t}=1\,|\,\mathbf{x}_{t}\right)=\sigma(\mathbf{x}_{t}^{\mathsf{T}}\mathbf{w}^{\star})$. The true parameter $\mathbf{w}^{\star}\in\mathcal{W}=\mathcal{B}(D/2)$ is unknown, and $\sigma(z)=1/(1+\exp(-z))$ denotes the sigmoid function. The performance of the learner is measured by its pseudo-regret, quantifying the cumulative expected loss against the optimal arms in hindsight, which is defined as
$\textsc{Reg}_{T}=\sum_{t=1}^{T}\left(\sigma(\mathbf{x}_{t}^{\mathsf{T}}\mathbf{w}^{\star})-\sigma({\mathbf{x}^{\star}_{t}}^{\mathsf{T}}\mathbf{w}^{\star})\right),$
where $\mathbf{x}^{\star}_{t}=\operatorname*{arg\,max}_{\mathbf{x}\in\mathcal{X}_{t}}\mathbf{x}^{\mathsf{T}}\mathbf{w}^{\star}$.
A challenge in GLB lies in its dependence on $\kappa=\max_{\mathbf{x}\in\mathcal{X},\mathbf{w}\in\mathcal{W}}1/\sigma^{\prime}(\mathbf{x}^{\mathsf{T}}\mathbf{w})$, a problem-dependent constant that may grow expoenentially with $D$. Directly applying ONS to GLB yields a pseudo-regret of $\widetilde{O}(\kappa d\sqrt{T})$ [Jun et al., 2017], which becomes vacuous for $\kappa=\Omega(\sqrt{T})$.
Current results.
Recently, Zhang et al. [2025] propose the first one-pass GLB algorithm achieving $\kappa$-free-leading-term pseudo-regret with constant working memory and constant per-round time. Their pseudo-regret is $O(d\sqrt{T}\log T+\kappa(d\log T)^{2})$, which remains sublinear in $T$ for $\kappa=o(T)$.
Their key idea is to estimate the parameter $\mathbf{w}^{\star}$ with a “look-ahead” variant of ONS, whose analytical properties can remove the dependence on $\kappa$ in the pseudo-regret’s leading terms while retaining the constant-memory and constant-time efficiency of ONS. An illustration of their “look-ahead” ONS is provided as follows:
$\widetilde{H}_{t}=\epsilon I+\left(\sum_{i=1}^{t-1}\nabla^{2}\ell_{i}(\mathbf{w}_{i+1})\right)+\eta\nabla^{2}\ell_{t}(\mathbf{w}_{t}),\quad\widehat{\mathbf{w}}_{t+1}=\mathbf{w}_{t}-\frac{1}{\eta}\widetilde{H}_{t}^{-1}\nabla\ell_{t}(\mathbf{w}_{t}),\quad\mathbf{w}_{t+1}=\Pi^{\widetilde{H}_{t}}_{\mathcal{B}(D/2)}[\widehat{\mathbf{w}}_{t+1}],$ (30)
where $\ell_{t}(\mathbf{w})=-y_{t}\log\sigma(\mathbf{x}_{t}^{\mathsf{T}}\mathbf{w})-(1-y_{t})\log(1-\sigma(\mathbf{x}_{t}^{\mathsf{T}}\mathbf{w}))$ denotes the logistic loss for parameter estimation at round $t$.
Because of ONS, their algorithm incurs the worst-case $\widetilde{O}((d^{2}K+d^{\omega})T)$ runtime bottleneck, where $K$ is the number of arms. Below we show that plugging LightONS can improve the runtime while preserving the $\kappa$-free-leading-term pseudo-regret guarantee.
Improvements by LightONS.
We propose an analogous “look-ahead” variant of LightONS in Algorithm˜5, which replaces the ONS-based subroutine in Zhang et al. [2025]. Line 7 of Algorithm˜5 introduces the deferred-projection mechanism of LightONS.Core to replace the Mahalanobis projection in Eq.˜30.
Algorithm˜5 improves the runtime from $\widetilde{O}((d^{2}K+d^{\omega})T)$ to $\widetilde{O}(d^{2}KT+d^{\omega}\cdot\min\{\sqrt{\kappa dT},T\})$, while preserving the $\kappa$-free-leading-term pseudo-regret guarantee. When $\kappa=o(T)$, as required for sublinear pseudo-regret, the improved algorithm is asymptotically faster than the original algorithm. We state the theoretical guarantee of Algorithm˜5 in the theorem below.
Algorithm 5 “look-ahead” LightONS for [Zhang et al., 2025]
0: domain $\mathcal{W}=\mathcal{B}(D/2)$, regularization coefficient $\lambda$, inverse step size $\eta$.
1: Initialize $H_{0}=\epsilon I$, $\mathbf{w}=\bm{0}$.
2:for $t=1,\dots,T$ do
3:  Update the lower confidence bound function as in Algorithm 1 of Zhang et al. [2025].
4:  Select the arm $\mathbf{x}_{t}$ as the Algorithm 1 of Zhang et al. [2025] and observe the loss $y_{t}$.
5:  $\widetilde{H}_{t}=H_{t-1}+\eta\nabla^{2}\ell_{t}(\mathbf{w}_{t})$.
6:  $\widehat{\mathbf{w}}_{t+1}=\mathbf{w}_{t}-\frac{1}{\eta}\widetilde{H}_{t}^{-1}\nabla\ell_{t}(\mathbf{w}_{t})$.
7:  $\mathbf{w}_{t+1}=\begin{cases}\widehat{\mathbf{w}}_{t+1}&\text{if}\penalty 10000\ \left\|\widehat{\mathbf{w}}_{t+1}\right\|_{2}\leq kD/2\\ \Pi^{\widetilde{H}_{t}}_{\mathcal{B}(D/2)}[\widehat{\mathbf{w}}_{t+1}]&\text{otherwise}\end{cases}$.
8:  $H_{t}=H_{t-1}+\nabla^{2}\ell_{t}(\mathbf{w}_{t+1})$.
9:end for
Theorem 6 (LightONS’s improvement for logistic bandits).
If $\bigcup_{t=1}^{T}\mathcal{X}_{t}\subseteq\mathcal{B}(1)$, $\left\|\mathbf{w}^{\star}\right\|\leq D/2$ and $D$ is known to the learner, then there exists an algorithm for binary logistic bandits, whose pseudo-regret is $O(d\sqrt{T}\log T+\kappa(d\log T)^{2})$, and whose total runtime is $\widetilde{O}(d^{2}KT+d^{\omega}\cdot\min\{\sqrt{\kappa dT},T\})$.
We remark that that Algorithm˜5 does not suffer from improper learning concerns, because the algorithm’s decision is the arm $\mathbf{x}_{t}$, rather than the estimated parameter $\mathbf{w}_{t}$. Consequently, Algorithm˜5 is free from the improper-to-proper conversion and resembles LightONS.Core.
Before proving Theorem˜6, we first present two critical lemmas that facilitate migrating the original ONS-based subroutine to LightONS-based Algorithm˜5. The following lemma resembles our Lemma˜11 and Lemma 1 of Zhang et al. [2025], showing that Algorithm˜5 implies an OMD-like regret decomposition form similar to that of original ONS-based subroutine.
Lemma 13.
Algorithm˜5 satisfies that, for any $\mathbf{u}\in\mathcal{B}(D/2)$
$\left\|\mathbf{w}_{t+1}-\mathbf{u}\right\|_{H_{t-1}}^{2}\leq 2\eta\nabla\widetilde{\ell}_{t}(\mathbf{w}_{t+1})^{\mathsf{T}}(\mathbf{u}-\mathbf{w}_{t+1})+\left\|\mathbf{w}_{t}-\mathbf{u}\right\|_{H_{t-1}}^{2}-\left\|\mathbf{w}_{t}-\mathbf{w}_{t+1}\right\|_{H_{t-1}}^{2},$
where $\widetilde{\ell}_{t}$ is the second-order Taylor expansion of $\ell_{t}$ at $\mathbf{w}_{t}$, i.e.,
$\widetilde{\ell}_{t}(\mathbf{w})\triangleq\ell_{t}(\mathbf{w}_{t})+\nabla\ell_{t}(\mathbf{w}_{t})^{\mathsf{T}}(\mathbf{w}-\mathbf{w}_{t})+\frac{1}{2}\left\|\mathbf{w}-\mathbf{w}_{t}\right\|_{\nabla^{2}\ell_{t}(\mathbf{w}_{t})}^{2}.$
Proof of Lemma˜13.
We note that the descend-and-project update in Algorithm˜5 is equivalent to the OMD update, as Appendix D of Zhang et al. [2025] has shown. Specifically,
$\mathbf{w}_{t+1}$ $=\Pi^{\widetilde{H}_{t}}_{\mathcal{B}(D/2)}[\widehat{\mathbf{w}}_{t+1}]=\Pi^{\widetilde{H}_{t}}_{\mathcal{B}(D/2)}\left[\mathbf{w}_{t}-\frac{1}{\eta}\widetilde{H}_{t}^{-1}\nabla\ell_{t}(\mathbf{w}_{t})\right]$ (31)
$\iff\mathbf{w}_{t+1}$ $=\operatorname*{arg\,min}_{\mathbf{w}\in\mathcal{B}(D/2)}\widetilde{\ell}_{t}(\mathbf{w})+\frac{1}{2\eta}\left\|\mathbf{w}-\mathbf{w}_{t}\right\|_{H_{t-1}}^{2},$
and
$\mathbf{w}_{t+1}$ $=\widehat{\mathbf{w}}_{t+1}=\mathbf{w}_{t}-\frac{1}{\eta}\widetilde{H}_{t}^{-1}\nabla\ell_{t}(\mathbf{w}_{t})$ (32)
$\iff\mathbf{w}_{t+1}$ $=\operatorname*{arg\,min}_{\mathbf{w}\in\mathbb{R}^{d}}\widetilde{\ell}_{t}(\mathbf{w})+\frac{1}{2\eta}\left\|\mathbf{w}-\mathbf{w}_{t}\right\|_{H_{t-1}}^{2}.$
To recover Lemma 1 of Zhang et al. [2025], we examine whether the following inequality holds
$\nabla\widetilde{\ell}_{t}(\mathbf{w}_{t+1})^{\mathsf{T}}\left(\mathbf{w}_{t+1}-\mathbf{u}\right)\leq\frac{1}{2\eta}\left(\left\|\mathbf{w}_{t}-\mathbf{u}\right\|_{H_{t-1}}^{2}-\left\|\mathbf{w}_{t+1}-\mathbf{u}\right\|_{H_{t-1}}^{2}-\left\|\mathbf{w}_{t+1}-\mathbf{w}_{t}\right\|_{H_{t-1}}^{2}\right).$ (33)
When $\left\|\widehat{\mathbf{w}}_{t+1}\right\|_{2}\leq kD/2$ and the Mahalanobis projection is not triggered, by Eq.˜31 we have
$\nabla_{\mathbf{w}=\mathbf{w}_{t+1}}\left(\widetilde{\ell}_{t}(\mathbf{w})+\frac{1}{2\eta}\left\|\mathbf{w}-\mathbf{w}_{t}\right\|_{H_{t-1}}^{2}\right)^{\mathsf{T}}(\mathbf{w}_{t+1}-\mathbf{u})=\bm{0}^{\mathsf{T}}(\mathbf{w}_{t+1}-\mathbf{u})=0.$
Rearranging terms, we have
$\nabla\widetilde{\ell}_{t}(\mathbf{w}_{t+1})^{\mathsf{T}}\left(\mathbf{w}_{t+1}-\mathbf{u}\right)$ $=-\frac{1}{\eta}(\mathbf{w}_{t+1}-\mathbf{w}_{t})^{\mathsf{T}}H_{t-1}(\mathbf{w}_{t+1}-\mathbf{u})$
$=\frac{1}{2\eta}\left(\left\|\mathbf{w}_{t}-\mathbf{u}\right\|_{H_{t-1}}^{2}-\left\|\mathbf{w}_{t+1}-\mathbf{u}\right\|_{H_{t-1}}^{2}-\left\|\mathbf{w}_{t+1}-\mathbf{w}_{t}\right\|_{H_{t-1}}^{2}\right),$
which means that Eq.˜33 holds with equality.
When $\left\|\widehat{\mathbf{w}}_{t+1}\right\|_{2}>kD/2$ and the Mahalanobis projection is triggered, by Eq.˜32, we have
$\nabla_{\mathbf{w}=\mathbf{w}_{t+1}}\left(\widetilde{\ell}_{t}(\mathbf{w})+\frac{1}{2\eta}\left\|\mathbf{w}-\mathbf{w}_{t}\right\|_{H_{t-1}}^{2}\right)^{\mathsf{T}}(\mathbf{w}_{t+1}-\mathbf{u})\leq 0.$
Rearranging terms, we have
$\nabla\widetilde{\ell}_{t}(\mathbf{w}_{t+1})^{\mathsf{T}}\left(\mathbf{w}_{t+1}-\mathbf{u}\right)$ $\leq-\frac{1}{\eta}(\mathbf{w}_{t+1}-\mathbf{w}_{t})^{\mathsf{T}}H_{t-1}(\mathbf{w}_{t+1}-\mathbf{u})$
$=\frac{1}{2\eta}\left(\left\|\mathbf{w}_{t}-\mathbf{u}\right\|_{H_{t-1}}^{2}-\left\|\mathbf{w}_{t+1}-\mathbf{u}\right\|_{H_{t-1}}^{2}-\left\|\mathbf{w}_{t+1}-\mathbf{w}_{t}\right\|_{H_{t-1}}^{2}\right),$
which means that Eq.˜33 holds.
Therefore, combining both cases, we have that Eq.˜33 always holds. ∎
The next lemma gives the total runtime of Algorithm˜5.
Lemma 14.
Algorithm˜5 has a runtime of $\widetilde{O}(d^{2}T+d^{\omega}\cdot\min\{\sqrt{\kappa dT},T\})$.
Proof of Lemma˜14.
Compared with the runtime of LightONS Theorem˜2, the only difference the dependence is that $\kappa$ appears in the number of Mahalanobis projections. It suffices to show how the logistic loss function affects the analysis of Lemma˜2. Specifically, we need to bound the quantity
$\Phi^{\prime}_{T}\triangleq\sum_{t=1}^{T}\left\|\frac{1}{\eta}\widetilde{H}_{t}^{-1}\nabla\ell_{t}(\mathbf{w}_{t})\right\|_{2}^{2}.$
We recall that $\ell_{t}(\mathbf{w})=-y_{t}\log\sigma(\mathbf{x}_{t}^{\mathsf{T}}\mathbf{w})-(1-y_{t})\log(1-\sigma(\mathbf{x}_{t}^{\mathsf{T}}\mathbf{w}))$ is the logistic loss function, with $\nabla\ell_{t}(\mathbf{w})=\left(\sigma(\mathbf{x}_{t}^{\mathsf{T}}\mathbf{w})-y_{t}\right)\mathbf{x}_{t}$ and $\nabla^{2}\ell_{t}(\mathbf{w})=\sigma^{\prime}(\mathbf{x}_{t}^{\mathsf{T}}\mathbf{w})\mathbf{x}_{t}\mathbf{x}_{t}^{\mathsf{T}}$, where $\sigma(z)=1/(1+\exp(-z))$ is the sigmoid function.
By the update rule of Algorithm˜5 and that $\eta>1$ in [Zhang et al., 2025], we have
$\widetilde{H}_{t}=\epsilon I+\sum_{i=1}^{t-1}\sigma^{\prime}(\mathbf{x}_{i}^{\mathsf{T}}\mathbf{w}_{i+1})\mathbf{x}_{i}\mathbf{x}_{i}^{\mathsf{T}}+\eta\sigma^{\prime}(\mathbf{x}_{t}^{\mathsf{T}}\mathbf{w}_{t})\mathbf{x}_{t}\mathbf{x}_{t}^{\mathsf{T}}\succ\lambda I+\sum_{i=1}^{t}\frac{1}{\kappa}\mathbf{x}_{i}\mathbf{x}_{i}^{\mathsf{T}},$
where $\kappa=\max_{\mathbf{x}\in\mathcal{X},\mathbf{w}\in\mathcal{W}}1/\sigma^{\prime}(\mathbf{x}^{\mathsf{T}}\mathbf{w})$. Since $\left|\sigma(\mathbf{x}_{t}^{\mathsf{T}}\mathbf{w}_{t})-y_{t}\right|\leq 1$, we have
$\left\|\widetilde{H}_{t}^{-1}\nabla\ell_{t}(\mathbf{w}_{t})\right\|_{2}^{2}$ $=\left\|\widetilde{H}_{t}^{-1}(\sigma(\mathbf{x}_{t}^{\mathsf{T}}\mathbf{w}_{t})-y_{t})\mathbf{x}_{t}\right\|_{2}^{2}\leq\left\|\widetilde{H}_{t}^{-1}\mathbf{x}_{t}\right\|_{2}^{2}$
$<\left\|\left(\lambda I+\sum_{i=1}^{t}\frac{1}{\kappa}\mathbf{x}_{i}\mathbf{x}_{i}^{\mathsf{T}}\right)^{-1}\mathbf{x}_{t}\right\|_{2}^{2}=\kappa^{2}\left\|\left(\kappa\lambda I+\sum_{i=1}^{t}\mathbf{x}_{i}\mathbf{x}_{i}^{\mathsf{T}}\right)^{-1}\mathbf{x}_{t}\right\|_{2}^{2}.$
Finally, summing up the preceding inequality with Lemma˜6, we have
$\Phi^{\prime}_{T}<\frac{\kappa^{2}}{\eta^{2}}\sum_{t=1}^{T}\left\|\left(\kappa\lambda I+\sum_{i=1}^{t}\mathbf{x}_{i}\mathbf{x}_{i}^{\mathsf{T}}\right)^{-1}\mathbf{x}_{t}\right\|_{2}^{2}\leq\frac{\kappa d}{\eta^{2}\lambda}.$
Therefore, reusing the proof of Lemma˜2, we obtain that the runtime of Algorithm˜5 is $\widetilde{O}(d^{2}T+d^{\omega}\cdot\min\{\sqrt{\kappa dT},T\})$. We note that $H_{t}$ and $\widetilde{H}_{t}$ also admit rank-one updates. ∎
Based on the preceding lemmas, we are ready to prove Theorem˜6.
Proof of Theorem˜6.
The upper-confidence-bound-based pseudo-regret analysis in [Zhang et al., 2025] primarily relies on their Theorem 1, which constitutes their Lemmas 4–6. To prove Theorem˜6, we examine how replacing the ONS-based subroutine with LightONS-based Algorithm˜5 affects these lemmas, except the moderate expansion of the diameter from $D$ to $D^{\prime}=kD=2D$.
  • Modifications to their Lemma 4. Their Lemma 4 is supported by their Lemma 1 and local relaxation of generalized linear models. Lemma˜13 shows that their Lemma 1 in still holds when using Algorithm˜5 as a replacement. The local relaxation is independent of the specific update rules of $\mathbf{w}_{t}$ and remains valid.
  • Modifications to their Lemma 5 and 6. Their Lemma 5 and 6 depend solely on the structure of the covariance matrix, i.e., $H_{t}=\lambda I+\sum_{i=1}^{t}\sigma^{\prime}(\mathbf{x}_{i}^{\mathsf{T}}\mathbf{w}_{i+1})\mathbf{x}_{i}\mathbf{x}_{i}^{\mathsf{T}}$. Therefore, Algorithm˜5 does not affect their Lemma 5 and 6.
Finally, the runtime of Algorithm˜5 follows from Lemma˜14. With the overhead $O(d^{2}KT)$ for selecting arms as in [Zhang et al., 2025], the total runtime is $\widetilde{O}(d^{2}KT+d^{\omega}\cdot\min\{\sqrt{\kappa dT},T\})$. ∎
We remark that extending Algorithm˜5 from the binary logistic bandits to the generalized linear bandits directly follows from [Zhang et al., 2025].
Comparison with OQNS.
OQNS can hardly be integrated into the method of [Zhang et al., 2025], as it is tailored to the OXO protocol and does not admit a regret decomposition analogous to Lemma˜13, which is essential for eliminating the $\kappa$ of leading terms in pseudo-regret. Moreover, OQNS lacks the flexibility to accommodate customized local norms beyond those used in standard ONS. Specifically, the local norm $\widetilde{H}_{t}$ in [Zhang et al., 2025], i.e., Eq.˜30, is not a simple accumulation of gradient outer products, unlike the OQNS’s local norm in Eq.˜4. In contrast, LightONS retains the structural flexibility of ONS, enabling integration into GLB with minimal modifications.
E.3 Proofs and Details for Memory-Efficient OXO
In this subsection, we seek to reduce the working memory required to achieve the optimal $O(d\log T)$ regret in OXO, albieit conditioned on geometric properties of the problem instance.
Current results.
Luo et al. [2016] propose Sketched Online Newton Step (SON), which achieves linear-in-$d$ runtime and working memory when the sketched dimension $d^{\prime}\ll d$. The value of $d^{\prime}$ typically depends on the problem’s intrinsic dimensionality, such as the number of non-zero eigenvalues in the Hessian-related matrix. The theoretical guarantees of SON are summarized in the following proposition.
Proposition 4 (Theorem 3 of Luo et al. [2016]).
Under Assumption˜2 and additional assumptions described in Eqs.˜36 and 37, SON (Algorithms 1 and 6 of Luo et al. [2016]) satisfies that,  [^4] The $\log T$ term in SON’s runtime arises from the eigendecomposition underlying sketching. Although Theorem 3 of Luo et al. [2016] ignores it, we include it for comparison with LightONS and OQNS.
$\textsc{Reg}_{T}(\mathbf{u})$ $\leq\frac{d^{\prime}}{\gamma}\log\left(1+\frac{G^{2}}{2d^{\prime}\epsilon}T\right)+\frac{\gamma\epsilon D^{2}}{8}+\frac{\Delta_{1:T}}{2\gamma},$ (34a)
Runtime $\leq O\left(d^{\prime}dT\log T\right),$ (34b)
and the working memory is $O(d^{\prime}d)$. The cumulative sketching error $\Delta_{1:T}$ can be bounded as
$\Delta_{1:T}\leq\min_{j\in[d^{\prime}]}\frac{2d^{\prime}}{(d^{\prime}-j+1)\epsilon}\sum_{i=j}^{d}\lambda_{i}\left(\sum_{t=1}^{T}\nabla f_{t}(\mathbf{x}_{t})\nabla f_{t}(\mathbf{x}_{t})^{\mathsf{T}}\right).$ (35)
However, SON demands additional assumptions. Its domain $\mathcal{X}_{t}$ must be an intersection of two parallel half-spaces, onto which Mahalanobis projections admit closed-form with $O(d^{2})$ time. For general bounded convex domains in OXO, SON loses its computational advantage and reverts to the high computational cost of ONS. The domain restriction further imposes a stronger assumption on loss functions. The curvature parameter $\gamma_{0}=\frac{1}{2}\min\{\frac{1}{DG},\alpha\}$ in Lemma˜1 collapses to zero for the unbounded intersection-of-parallel-half-spaces domain, reducing exp-concavity to convexity. Consequently, SON assumes an explicit quadratic property mirroring Lemma˜1, rather than the standard Assumption˜3. We list these additional assumptions as follows:
  • Additional assumption on domains. For any $t\in[T]$, the domain at the $t$-th round is the intersection of two parallel half-spaces, i.e.,
    $\mathcal{X}_{t}=\{\mathbf{x}\,|\,|\mathbf{w}_{t}^{\mathsf{T}}\mathbf{x}|<D/2\},\quad\text{where}\penalty 10000\ \mathbf{w}_{t}\in\mathbb{R}^{d}\penalty 10000\ \text{is known and}\penalty 10000\ \left\|\mathbf{w}_{t}\right\|=1.$ (36)
  • Additional assumption on loss functions. For any $t\in[T]$, the curvature parameter $\gamma$, the loss function $f_{t}$, the trajectory $\{\mathbf{x}_{t}\}_{t=1}^{T}$ and the comparator $\mathbf{u}\in\bigcap_{t=1}^{T}\mathcal{X}_{t}$ satisfies that
    $f_{t}(\mathbf{x})-f_{t}(\mathbf{u})\leq\nabla f_{t}(\mathbf{x})^{\mathsf{T}}(\mathbf{x}-\mathbf{u})-\frac{\gamma}{2}\left(\nabla f_{t}(\mathbf{x})^{\mathsf{T}}(\mathbf{x}-\mathbf{u})\right)^{2}.$ (37)
Improvements by LightONS.
We propose LightONS.Sketch in Algorithm˜6, replacing LightONS’s accesses to the Hessian-related matrix $A_{t}$ with the sketching primitives as in SON. LightONS.Sketch combines the projection efficiency of LightONS with the memory efficiency of SON. This hybrid method extends linear-in-$d$ runtime and working memory to the standard OXO setting (Assumptions˜1, 2 and 3), while retaining the SON’s regret $O(d^{\prime}\log T)$. Theoretical guarantees of the hybrid method are summarized in the following theorem.
Theorem 7 (LightONS’s improvement for memory-efficient OXO).
Under Assumptions˜1, 2 and 3,
with $\gamma_{0}=\frac{1}{2}\min\{\frac{1}{DG},\alpha\}$, LightONS.Sketch (Algorithm˜6) satisfies that, for any $\mathbf{u}\in\mathcal{X}$,
$\textsc{Reg}_{T}(\mathbf{u})$ $\leq\frac{d^{\prime}}{\gamma_{0}}\log\left(1+\frac{G^{2}}{2d^{\prime}\epsilon}T\right)+\frac{\gamma_{0}\epsilon D^{2}}{8}+\frac{\Delta_{1:T}}{2\gamma_{0}},$ (38a)
Runtime $\leq O\left(\left(\texttt{EP}_{\mathcal{X}}+d^{\prime}d\log T\right)T+d^{\omega}\sqrt{(d+\Delta_{1:T})T/\epsilon}\log T\right),$ (38b)
and the working memory is $O(d^{\prime}d)$. With the surrogate gradient $\nabla g_{t}(\mathbf{y}_{t})$ as in Lemma˜3, the cumulative sketching error $\Delta_{1:T}$ can be bounded as
$\Delta_{1:T}\leq\min_{j\in[d^{\prime}]}\frac{2d^{\prime}}{(d^{\prime}-j+1)\epsilon}\sum_{i=j}^{d}\lambda_{i}\left(\sum_{t=1}^{T}\nabla g_{t}(\mathbf{y}_{t})\nabla g_{t}(\mathbf{y}_{t})^{\mathsf{T}}\right).$ (39)
Algorithm 6 LightONS.Sketch
0: preconditioner coefficient $\epsilon$, deferral coefficient $k$, dimension to reduce to $d^{\prime}$.
1: Initialize $\gamma^{\prime}=\frac{1}{2}\min\{\frac{1}{c_{f}c_{g}DG},\frac{4}{c_{f}c_{g}(k+1)DG},\alpha\}$, $S_{0}=O_{2d^{\prime}\times d}$, $R_{0}=\frac{1}{\epsilon}I_{2d^{\prime}\times 2d^{\prime}}$, $\mathbf{x}_{1}=\mathbf{y}_{1}=\bm{0}$.
2:for $t=1,\dots,T$ do
3:  Observe $\nabla f_{t}(\mathbf{x}_{t})$; and construct $\nabla g_{t}(\mathbf{y}_{t})$ satisfying Condition˜1.
[…]
7:  $\mathbf{x}_{t+1}=\Pi_{\mathcal{X}}[\mathbf{y}_{t+1}]$.
8:end for
Algorithm 7 Fast Frequent Directions of Ghashami et al. [2016]
0: new gradient $\nabla g_{t}(\mathbf{y}_{t})$, frequent directions $S_{t-1}$, low-dimension inverse $R_{t-1}$.
0: updated frequent directions $S_{t}$, updated low-dimension inverse $R_{t}$.
[…]
2:if $S_{t}$ still has all-zero rows then
3:  $R_{t}=\left(\epsilon I+S_{t}S_{t}^{\mathsf{T}}\right)^{-1}$.
4:else
5:  $U_{t}\Sigma_{t}V_{t}^{\mathsf{T}}=S_{t}$, truncated SVD with top $d^{\prime}$ singular values.
[…]
7:  $R_{t}=\operatorname{diag}\left(\frac{1}{\epsilon+(\Sigma_{t})_{1,1}^{2}-(\Sigma_{t})_{d^{\prime},d^{\prime}}^{2}},\dots,\frac{1}{\epsilon+(\Sigma_{t})_{d^{\prime},d^{\prime}}^{2}-(\Sigma_{t})_{d^{\prime},d^{\prime}}^{2}},\frac{1}{\epsilon},\dots,\frac{1}{\epsilon}\right)$.
8:end if
The key difference between LightONS.Sketch and LightONS is the storage strategy. Instead of storing full matrices $A_{t}\in\mathbb{R}^{d\times d}$ and $V_{t}\in\mathbb{R}^{d\times d}$, LightONS.Sketch maintains compact sketches $S_{t}\in\mathbb{R}^{2d^{\prime}\times d}$ and $R_{t}\in\mathbb{R}^{2d^{\prime}\times 2d^{\prime}}$. Each row of $S_{t}$ stores a principal gradient component, while $R_{t}$ plays a role analogous to $V_{t}$ in Eq.˜16. The full matrix and its inverse are reconstructed as
$\widetilde{A}_{t}=\epsilon I+S_{t}^{\mathsf{T}}S_{t},\quad\widetilde{A}_{t}^{-1}=\frac{1}{\epsilon}\left(I-S_{t}^{\mathsf{T}}R_{t}S_{t}\right),\quad R_{t}=\left(\epsilon I+S_{t}S_{t}^{\mathsf{T}}\right)^{-1}.$
This relationship follows from the Sherman-Morrison-Woodbury formula:
$(A+BCD)^{-1}=A^{-1}-A^{-1}B(C^{-1}+DA^{-1}B)^{-1}DA^{-1}.$
Following Luo et al. [2016], LightONS.Sketch uses Fast Frequent Directions of Ghashami et al. [2016] in Algorithm˜7 to update $S_{t}$ and $R_{t}$.
Before proving Theorem˜7, we first introduce three lemmas to characterize the error introduced by sketching. We introduce the notation $\Delta_{t}$ to denote the sketching error at the $t$-th round.
$\Delta_{t}\triangleq\begin{cases}\frac{2d^{\prime}}{\epsilon}\sigma_{d^{\prime}}(S_{t})^{2}&\text{if SVD is triggered at $t$-th round}\\ 0&\text{otherwise}\end{cases},$
where $\sigma_{i}(S_{t})$ is the $i$-th greatest singular value of $S_{t}$. Specifically, Lemma˜15 bound the error for regret analysis with Lemma˜5, while Lemma˜16 bound the error for projection-count analysis with Lemma˜6. Lemma˜17 bounds the total error with the spectrum of the Hessian-related matrix.
Lemma 15.
At the $t$-th round of Algorithm˜6, in Algorithm˜7, $\Delta_{t}$ satisfies:
$\left\|\nabla g_{t}(\mathbf{y}_{t})\right\|_{\widetilde{A}_{t}^{-1}}^{2}\leq\left\langle\widetilde{A}_{t}^{-1},\widetilde{A}_{t}-\widetilde{A}_{t-1}\right\rangle_{\operatorname{F}}+\Delta_{t}.$
Proof of Lemma˜15.
Let $\left\langle\cdot,\cdot\right\rangle_{\operatorname{F}}$ denote the inner product induced by the Frobenius matrix norm, let $(\cdot)_{:,i}$ denote the $i$-th column of the matrix, and let $\bar{U}_{t}\bar{\Sigma}_{t}\bar{V}_{t}^{\mathsf{T}}=S_{t}$ be the full SVD with all $2d^{\prime}$ singular values. When SVD is not triggered, then $\widetilde{A}_{t}=\widetilde{A}_{t-1}+\nabla g_{t}(\mathbf{y}_{t})\nabla g_{t}(\mathbf{y}_{t})^{\mathsf{T}}$ and $\Delta_{t}=0$ trivially holds. When SVD is triggered,
$\quad\left\langle\widetilde{A}_{t}^{-1},\widetilde{A}_{t-1}+\nabla g_{t}(\mathbf{y}_{t})\nabla g_{t}(\mathbf{y}_{t})^{\mathsf{T}}-\widetilde{A}_{t}\right\rangle_{\operatorname{F}}$
$=\sum_{i=1}^{2d^{\prime}}\min\left\{(\bar{\Sigma}_{t})_{i,i}^{2},(\bar{\Sigma}_{t})_{d^{\prime},d^{\prime}}^{2}\right\}\left\|(\bar{V}_{t})_{:,i}\right\|_{\widetilde{A}_{t}^{-1}}^{2}\leq\frac{2d^{\prime}}{\epsilon}(\bar{\Sigma}_{t})_{d^{\prime},d^{\prime}}^{2}\triangleq\Delta_{t}.$
The inequality uses the fact that $\widetilde{A}_{t}\succeq\epsilon I$ and $\left\|(\bar{V}_{t})_{:,i}\right\|_{\widetilde{A}_{t}^{-1}}^{2}\leq\left\|(\bar{V}_{t})_{:,i}\right\|_{\left(\epsilon I\right)^{-1}}^{2}=\frac{1}{\epsilon}$. Substituting $(\bar{\Sigma}_{t})_{i,i}=\sigma_{i}(S_{t})$ completes the proof. ∎
Lemma 16.
At the $t$-th round of Algorithm˜6, in Algorithm˜7, $\Delta_{t}$ also satisfies:
$\left\|\nabla g_{t}(\mathbf{y}_{t})\right\|_{\widetilde{A}_{t}^{-2}}^{2}=\left\langle\widetilde{A}_{t}^{-2},\widetilde{A}_{t}-\widetilde{A}_{t-1}\right\rangle_{\operatorname{F}}+\frac{\Delta_{t}}{\epsilon}.$
We omit proof of Lemma˜16, as it directly reuses the proof of Lemma˜15.
Lemma 17 (Theorem 1.1 and Section 3 of Ghashami et al. [2016]).
In Algorithm˜6, $\Delta_{1:T}$ satisfies that, for any $j\in[d^{\prime}]$,
$\Delta_{1:T}\triangleq\sum_{t=1}^{T}\Delta_{t}\leq\frac{2d^{\prime}}{(d^{\prime}-j+1)\epsilon}\sum_{i=j}^{d}\lambda_{i}\left(\sum_{t=1}^{T}\nabla g_{t}(\mathbf{y}_{t})\nabla g_{t}(\mathbf{y}_{t})^{\mathsf{T}}\right).$
Now we can prove Theorem˜6 by bounding difference between LightONS and LightONS.Sketch with the help of the preceding lemmas characterizing the sketching error.
Proof of Theorem˜7.
We note that Lemma˜11 holds for Algorithm˜6, as the proof of Lemma˜11 only requires $A_{t}$ to be positive-definite. Thus by Lemmas˜4 and 11 we obtain
[…]
where the last inequality uses the fact that $\widetilde{A}_{t}\preceq\widetilde{A}_{t-1}+\nabla g_{t}(\mathbf{y}_{t})\nabla g_{t}(\mathbf{y}_{t})^{\mathsf{T}}$, which is ensured by Algorithm˜7. Then plugging Lemma˜15 into the preceding inequality yields
$\quad f_{t}(\mathbf{x}_{t})-f_{t}(\mathbf{u})$
$\leq\frac{1}{2\gamma^{\prime}}\left(\left\langle\widetilde{A}_{t}^{-1},\widetilde{A}_{t}-\widetilde{A}_{t-1}\right\rangle_{\operatorname{F}}+\Delta_{t}\right)+\frac{\gamma^{\prime}}{2}\left\|\mathbf{y}_{t}-\mathbf{u}\right\|_{\widetilde{A}_{t-1}}^{2}-\frac{\gamma^{\prime}}{2}\left\|\mathbf{y}_{t+1}-\mathbf{u}\right\|_{\widetilde{A}_{t}}^{2}$
$\leq\frac{1}{2\gamma^{\prime}}\left(\log\operatorname{det}(\widetilde{A}_{t})-\log\operatorname{det}(\widetilde{A}_{t-1})+\Delta_{t}\right)+\frac{\gamma^{\prime}}{2}\left\|\mathbf{y}_{t}-\mathbf{u}\right\|_{\widetilde{A}_{t-1}}^{2}-\frac{\gamma^{\prime}}{2}\left\|\mathbf{y}_{t+1}-\mathbf{u}\right\|_{\widetilde{A}_{t}}^{2},$
where the equality uses the fact that $\left\langle X^{-1},X-Y\right\rangle_{\operatorname{F}}\leq\log\operatorname{det}(X)-\log\operatorname{det}(Y)$, which comes from the proof of Lemmas˜5 and 6. Telescoping the preceding inequality yields
$\sum_{t=1}^{T}\left(f_{t}(\mathbf{x}_{t})-f_{t}(\mathbf{u})\right)\leq\frac{1}{2\gamma^{\prime}}\log\frac{\operatorname{det}(\widetilde{A}_{T})}{\operatorname{det}(\widetilde{A}_{0})}+\frac{1}{2\gamma^{\prime}}\Delta_{1:T}+\frac{\gamma^{\prime}}{2}\left\|\mathbf{y}_{1}-\mathbf{u}\right\|_{\widetilde{A}_{0}}^{2},$
where the logarithmic term is further bounded with Jensen’s inequality, which differs slightly from Lemma˜5 due to the number of non-zero eigenvalues:
$\log\frac{\operatorname{det}(\widetilde{A}_{T})}{\operatorname{det}(\widetilde{A}_{0})}=\sum_{i=1}^{2d^{\prime}}\log\left(1+\frac{\sigma_{i}^{2}(S_{T})}{\epsilon}\right)\leq 2d^{\prime}\log\left(1+\frac{\left\|S_{T}\right\|_{\operatorname{F}}^{2}}{2d^{\prime}\epsilon}\right)\leq 2d^{\prime}\log\left(1+\frac{G^{2}}{2d^{\prime}\epsilon}T\right).$
Finally, the runtime follows from the following two parts:
  • Runtime aside from Algorithm˜7. Following the same analysis of Lemmas˜2 and 14, overall runtime aside from Algorithm˜7 is $O((k-1)^{-1}\sqrt{(d+\Delta_{1:T})T/\epsilon}\cdot d^{\omega}\log T)$ due to FastProj. It suffices to verify the following inequality which follows from Lemma˜16:
    $\Phi^{\prime\prime}_{T}\triangleq\sum_{t=1}^{T}\left\|\frac{1}{\gamma^{\prime}}\widetilde{A}_{t}^{-1}\nabla g_{t}(\mathbf{y}_{t})\right\|_{2}^{2}\leq\frac{1}{\gamma^{\prime 2}}\frac{d+\Delta_{1:T}}{\epsilon}.$
  • Runtime of Algorithm˜7. Each time when SVD is triggered, the last $d^{\prime}+1$ rows of $S_{t}$ become all-zero, thus the SVD is triggered at most $\lceil T/(d^{\prime}+1)\rceil$ times. The runtime of SVD is $O(d^{\prime 2}d\log T)$ [Golub and Van Loan, 2013] to achieve the desired accuracy that does not affect the regret bound.  [^5] The runtime of SVD can be improved to $O\big(d^{\prime 2}(d+\log T)\big)$ and the factor $O(\log T)$ can be independent of the minimal singular value gap of $S_{t}$ and only depends on the scale of singular values of $S_{t}$ [Parlett, 1998]. The runtime of updating $R_{t}$ is $O(d^{\prime}d)$, as this can be implemented with twice rank-one updates:
    $(A+\mathbf{u}\mathbf{v}^{\mathsf{T}})^{-1}=A^{-1}-\frac{1}{1+\mathbf{v}^{\mathsf{T}}A^{-1}\mathbf{u}}A^{-1}\mathbf{u}\mathbf{v}^{\mathsf{T}}A^{-1}.$
    Specifically,
    $\quad R_{t}^{-1}=\epsilon I+S_{t}S_{t}^{\mathsf{T}}=\epsilon I+\left(S_{t-1}+\mathbf{e}_{i_{t}}\nabla g_{t}(\mathbf{y}_{t})^{\mathsf{T}}\right)\left(S_{t-1}+\mathbf{e}_{i_{t}}\nabla g_{t}(\mathbf{y}_{t})^{\mathsf{T}}\right)^{\mathsf{T}}$
    $=\underbrace{\epsilon I+S_{t-1}S_{t-1}^{\mathsf{T}}}_{R_{t-1}^{-1}}+\mathbf{e}_{i_{t}}\underbrace{\left(S_{t-1}\nabla g_{t}(\mathbf{y}_{t})\right)^{\mathsf{T}}}_{\mathbf{a}_{t}^{\mathsf{T}}}+\underbrace{\left(S_{t-1}\nabla g_{t}(\mathbf{y}_{t})+\mathbf{e}_{i_{t}}\nabla g_{t}(\mathbf{y}_{t})\nabla g_{t}(\mathbf{y}_{t})^{\mathsf{T}}\right)}_{\mathbf{b}_{t}}\mathbf{e}_{i_{t}}^{\mathsf{T}},$
    where $\mathbf{e}_{i_{t}}$, $\mathbf{a}_{t}$ and $\mathbf{b}_{t}$ are $2d^{\prime}$-dimensional vectors that can be computed in $O(d^{\prime}d)$ time. Therefore, the overall runtime of Algorithm˜7 is $O(d^{\prime 2}d(\log T)\lceil T/(d^{\prime}+1)\rceil+d^{\prime}dT)=O(d^{\prime}dT\log T)$.
Comparison with OQNS.
Mhammedi and Gatmiry [2023] mention the possibility of combining OQNS with sketching but provide neither algorithms nor analysis. Incorporating sketching further complicates OQNS’s already intricate analysis, as the sketching error interacts with the log-barrier, Hessian approximation and decisions’ update in Eq.˜4. In contrast, within the OMD framework, sketching errors naturally appear as an additive term in the regret bound via OMD’s stability term, illustrated in Eq.˜29, since OMD’s regret decomposition is somewhat orthogonal to sketching.